### v2

Abstract
Two office-driven politicians compete in an election by proposing policies. There are two possible states of the world: climate change is either mild, with no lasting effect on welfare if addressed properly, or severe, leading to reduced welfare even with appropriate measures. Voters receive signals about the state but may interpret them in a non-Bayesian way, holding motivated beliefs. An equilibrium always exists where voters ignore signals suggesting severe consequences, causing politicians to propose policies for mild climate change—even when they know otherwise. If severe climate change leads to only moderate welfare losses, an efficient equilibrium also exists. In this equilibrium, voters trust politicians to choose the optimal policies, implying voters choose to trust their signals, which in turn encourages optimal policy choices by politicians. The model highlights the role of political rhetoric and trust in government, and a first glance at the data reveals patterns consistent with the models predictions.
JEL Codes: D72, D91, H12
Keywords: political competition, climate change, motivated reasoning, trust in government, political rhetoric, misinformation
1 Introduction
“[…] information that increases perceptions of the reality of climate
change may feel so frightening that it leads to denial and thus
a reduction in concern and support for action.”
Clayton, Manning, Krygsman, and Speiser (2017)
“Yes, There Has Been Progress on Climate.
No, It’s Not Nearly Enough.”
The New York Times, 25 Oct 2021
Climate change is one of the most pressing issues of our time, threatening the livelihoods of millions of people around the globe. Information about climate change and global warming has been available for more than a century, dating back at least to Svante Arrhenius’ (1896) famous paper “On the Influence of Carbonic Acid in the Air upon the Temperature of the Ground.” However, despite an ever-growing scientific consensus that increasing atmospheric CO2 concentrations will severely impact our planet, little to no action was taken to stop this process for most of the last few decades. Many reasons have been put forward to explain the inaction of political decision-makers, ranging from lobbying interests to widespread misinformation campaigns like those of ExxonMobil. In this paper, I discuss a different, simpler channel: the electoral incentives of politicians when voters may choose to hold motivated beliefs.
The looming dire consequences of climate change may, if taken seriously, create stress and anxiety. In an effort to avoid these negative emotions, people may choose to hold motivated beliefs, ignoring information that suggests climate change is severe, while overreacting to information suggesting there is nothing to worry about. Psychologists and economists alike have long been aware of such information processing biases, as exemplified by Kunda’s (1990) seminal paper or the recent survey by Amelio and Zimmermann (2023), and multiple studies have provided evidence for their empirical relevance (for example, Lewandowsky, Gignac, and Oberauer, 2013, Thaler, 2021, 2024, and Lois, Tsakas, Yuen, and Riedl, 2024). To study the potential consequences of such behavior for the implementation of policies targeted at climate change, I construct a game-theoretic model in which two politicians compete in an election by proposing policies. There are two possible states of the world: severe climate change and moderate (or no) climate change. The state determines which policy is optimal and establishes a baseline welfare level. If climate change is severe, the baseline welfare is lower than in the mild state, regardless of the chosen policy. Voters and politicians receive signals about the state of the world, and society is best served when the policy that matches the state is implemented. As discussed by authors like Bénabou and Tirole (2016) or Spiegler (2016), voters have anticipatory utility, meaning they feel anxiety or stress when the future looks grim. To cope with these negative emotions, they may choose to hold motivated beliefs, interpreting information about the future in a non-Bayesian way to increase anticipatory utility. How voters react to information is an important determinant of how politicians compete to win an election.
The model highlights the importance of the portrayed severity of climate change for the efficiency of policy-making. In particular, if severe climate change has only moderate consequences for baseline welfare levels, then there always exists an efficient equilibrium in which politicians choose the optimal policies for each state. However, this equilibrium only exists if voters have sufficient trust in government, i.e., if they belief that elected politicians will enact the right policies with a sufficiently high probability. Otherwise, with low trust in government, there always exists an equilibrium in which politicians ignore their information and choose policies for mild or no climate change.
The analysis suggests two important determinants for the efficiency of policy-making. First, trust in government is important, because it determines in which way voters interpret information when severe climate change only has moderate effects on baseline welfare. With sufficient trust in government, there always exists an efficient equilibrium in which politicians strictly choose the optimal policies. To the contrary, if trust in government is low, then politicians will ignore their information and policy choice will be inefficient. Second, the analysis suggests that political rhetoric could be an important determinant of the quality of policy choice as well. In particular, if the consequences of severe climate change are portrayed as catastrophical, then the efficient equilibrium ceases to exist. Voters will ignore any information that climate change is severe, inducing office motivated politicians to ignore their information as well, and in the unique equilibrium politicians always choose policies for mild or no climate change. However, if the consequences of severe climate change are portrayed as moderate or mild, then the efficient equilibrium survives.
Literature:
The paper contributes to several areas of literature. Firstly, it adds to the literature that examines the incentives of political candidates to select efficient policy platforms when the true state of the world is unknown. Important contributions to this literature include Heidhues and Lagerlof (2003), Laslier and van der Straeten (2004), and Kartik, Squintani, and Tinn (2015). Unlike Heidhues and Lagerlof (2003) and Kartik, Squintani, and Tinn (2015), I assume that voters also receive information, which creates the main tension of whether to ignore it or not. In Laslier and van der Straeten (2004), voters receive a single public signal as well, whereas in the current paper, each voter receives an independent signal. The two most closely related papers in this literature are Gratton (2014) and Crutzen, Sisak, and Swank (2024). Gratton (2014) assumes that candidates observe the state perfectly while voters receive imperfect but informative signals about the state, and the focus of the analysis is to identify conditions that lead equilibria such that candidates choose efficient policies. Crutzen, Sisak, and Swank (2024) study a game like Gratton (2014), but there are two groups of voters, elites and commoners, who stochastically differ in their policy preferences and who always differ in quality of the information they receive. The analysis focusses on deriving conditions such that candidates cater to the elites or engage in populism. Unlike in those papers, the focus in the current paper is on misinterpreting information, and how this influences candidates’ incentives to provide efficient policies.
A cornerstone of the model is that voters have anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare” (Bénabou and Tirole, 2016). As in Akerlof and Dickens (1982), Caplin and Leahy (2001), Bénabou and Tirole (2002), Spiegler (2016), Eliaz and Spiegler (2020), and other papers, this means that voters have preferences not only about states and policies but also about their own beliefs. A consequence of this is that voters may engage in motivated reasoning; see, for example, Kunda (1990), Bénabou and Tirole (2016), Zimmermann (2020), Oprea and Yuksel (2021), or Thaler (2024). The most closely related papers in this area are Levy (2014) and Le Yaouanq (2023), which also examine electoral contexts. Levy (2014) presents a model featuring a policymaker trying to signal quality/congruency to voters. Voters have imperfect memory and can suppress certain news by purposefully conflating positive and negative signals. In contrast, there is no signaling in the current paper. Furthermore, voters have perfect memory but may choose to interpret some signals as more or less informative than they really are. Finally, it is not just voters’ beliefs about the state that matter, but also their expectations regarding policy outcomes. This difference introduces a novel self-fulfilling aspect that becomes critical for equilibrium. Le Yaouanq (2023) suggests that voters may interpret signals in a contrary manner at a cost. In the current paper, I assume that voters can modulate the perceived informativeness of a signal, but they cannot alter its direction. Moreover, unlike in Le Yaouanq (2023), a key component of my model is that the state influences baseline welfare levels, which has important consequences for equilibrium.
The paper also contributes to a recently revived literature on environmental policymaking. Delfgaauw and Swank (2024) show that there can be locked-in effects that prevent the adoption of environmentally friendly policies. Blumenthal (2024) shows that efficient environmental policies may not be adopted if voters’ preferences are expected to change over time. Van der Straten, Perotti, and Van der Ploeg (2024) study the relationship between climate change adoption policies and inequality in society. Besley and Persson (2023) examine the conditions for green transitions to occur, depending on voters’ ideological preferences and the extent to which the production sector remains reliant on brown energy sources. Gullberg (2008), Shapiro (2016), and Balles, Matter, and Stutzer (2024) show that special interest groups may prevent climate change policies from being adopted. Denter (2024) shows that the wrong kind of climate activism may decrease support for climate policies. In contrast to these papers, the current paper provides an explanation based purely on the electoral incentives of office-motivated politicians in the presence of a common information processing bias among voters.
The paper is organized as follows. In Section 2, I describe the baseline model, which is then solved in Section 3. Section 4 extends the baseline model along several dimensions. In Section 5, I derive testable predictions of the model and have a first quick look at the data. Section 6 concludes. All proofs are relegated to the appendix.
2 The Model
In this section I introduce the main building blocks of the model. There is an unknown state of the world, $\omega\in\{0,1\}$, indicating whether climate change is mild, $\omega=0$, or severe, $\omega=1$. The prior probability that $\omega=1$ is $q\in(0,1)$. There are two different kinds of actors. On the one hand, there are two kinds of purely office motivated politicians, indexed by $i=1,2$. On the other hand, there is a continuum of voters of mass 1. Candidates vie for voters’ support by proposing policies, $p_{i}\in\{0,1\}$, to which they commit. Denote the vector of policies by $\mathbf{p}\equiv(p_{1},p_{2})$.
Voters policy preferences depend on the state of the world $\omega$. Given a winning policy $p$ and given a state realization $\omega$, voters’ realized policy utility is
$u(p,\omega)=\left\{\begin{array}[]{cl}-p&\text{if }\omega=0,\\ -\Delta-\beta\left|p-1\right|&\text{if }\omega=1.\end{array}\right.$ (1)
Hence, for each state realization $\omega$, the optimal policy is $p=\omega$. But the state $\omega$ determines not only optimal policy, but the baseline welfare level in society. In particular, if $\omega=1$, i.e., when climate change is severe, then baseline welfare decreases by $\Delta>0$, independent of the chosen policy, while it remains constant if climate change is mild, $\omega=0$. The parameter $\beta>0$ measures the importance of choosing the right policy when climate change is severe: the greater is $\beta$, the more important is taking the right action. Note that the worst outcome is choosing policy incorrectly in state 1, yielding $-\Delta-\beta$, while the best outcome is choosing correctly in state 0, yielding a utility of zero. I make the following assumption to capture the idea that at the outset, voters believe the optimal policy choice is $p=0$: [^1] This assumption is not without loss of generality for the model’s conclusions. Allowing for $q\geq\frac{1}{1+\beta}$ would imply that the set of possible equilibria is larger. I chose to assume $q<\frac{1}{1+\beta}$, because it seems realistic that voters need to be convinced to support anti-climate change policies. Moreover, as a second benefit, Assumption 1 keeps the paper’s exposition more concise.
Assumption 1.
Throughout I assume that $q<\frac{1}{1+\beta}.$
At the beginning of the game, both types of players receive a signal that is informative about $\omega$. In particular, as in Gratton (2014), before announcing policy platforms every politician $i$ receives a signal $s_{i}^{P}$ that perfectly reveals $\omega$ Hence, after receiving $s_{i}^{P}$, each politician perfectly knows wether climate change is severe or not.
Voters receive two kinds of signals. On the one hand, the policy platforms $\mathbf{p}$ may function as signals about $\omega$, depending on the strategies chosen by the candidates. On the other hand, they receive information from news or from direct experience about $\omega$. I model this by assuming that each voter receives a signal $s\in\mathbb{R}$. When the state is $\omega$, the c.d.f. of $s$ is $\Phi\left(\frac{s+(2\omega-1)\mu}{\sigma}\right)$, where $\Phi$ is the c.d.f. of the standard Gaussian distribution and $\mu>0$ and $\sigma>0$ are parameters. Therefore, the typical monotone likelihood ratio property (MLRP) is satisfied, meaning that any $s>0$ is evidence for $\omega=1$ and increases the belief compared to the prior, while any $s<0$ is evidence for $\omega=0$ and decreases the belief. Following Callander (2011) and others, I interpret $\mu$ as a measure of the complexity of the issue climate change. If $\mu$ is small, the issue is very complex, and thus voters tend to hold imprecise beliefs about $\omega$. To the contrary, if $\mu$ is large, climate change is not a very complex issue and beliefs tend to be precise.
We assume that the complexity of the issue climate change is such that, absent motivated beliefs and any signaling about the state $\omega$ through platform choices $\mathbf{p}$, the election aggregates information whenever $p_{1}\neq p_{2}$. In other words: if both policies are offered, a majority chooses to vote for the welfare-maximizing policy. The following assumption guarantees that this is true:
Assumption 2.
Throughout I assume that
$\mu\geq\mu^{\text{Bayes}}\equiv\frac{\sigma\sqrt{\ln\left(\frac{1-q}{\beta q}\right)}}{\sqrt{2}}.$
I assume that voters have anticipatory utility. In particular, a voter has expectations about her future utility, which is derived from a belief bout the true state and a second belief about the enacted policy. This causes anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare(Bénabou and Tirole, 2016). To increases anticipatory utility (to be defined precisely below), a voters may use motivated reasoning, which means that she may update beliefs using a distorted complexity parameter $\tilde{\mu}$. Clearly, processing information using such a distorted complexity parameter comes at a cost, because higher anticipatory utility may imply lower utility because of imperfect decision making. As Bénabou and Tirole (2016) write, “one can react to bad news objectively, which leads to better decisions but having to live with grim prospects for some time […], or adopt a more “defensive” cognitive response that makes life easier until the day of reckoning, when mistakes will have to be paid for.” I do not model these costs in the baseline model to be able to identify the pure effect of motivated reasoning, but in Section 4.4 two models of distortions are discussed. If a voter is indifferent between having $\tilde{\mu}=\mu$ and some other $\tilde{\mu}\neq\mu$, I assume without loss of generality that she chooses $\tilde{\mu}^{*}=\mu$. Choosing $\tilde{\mu}>\mu$ implies the voter interprets the signal as more informative than it really is, and hence beliefs will change excessively, while choosing $\tilde{\mu}<\mu$ implies a more conservative stance and that beliefs move less than they should.
Based on $s$ and $\mathbf{p}$, and after choosing $\tilde{\mu}$, each voter forms beliefs $\pi(s,\mathbf{p},\tilde{\mu})$ and $\kappa$. $\pi(s,\mathbf{p},\tilde{\mu})$ is the posterior belief about $\omega$, whereas $\kappa$ is the probability assigned to the event that policy 1 will be implemented after the election. Because candidates commit to policy platforms, clearly $\kappa=1$ if $p_{1}=p_{2}=1$, and $\kappa=0$ if $p_{1}=p_{2}=0$.
Equipped with these beliefs, we can now calculate a voter’s anticipatory utility:
$AU(s,\mathbf{p},\tilde{\mu})=-\kappa\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot\Delta+(1-\pi(s,\mathbf{p},\tilde{\mu}))\right]-(1-\kappa)\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot(\Delta+\beta)\right]$ (2)
Thus, anticipatory utility equals the utility the voter expects to receive once policies are determined.
Note that a voter is never pivotal, because there is a continuum of voters. I thus assume that voters vote sincerely as defined by Austen-Smith and Banks (1996). In words, each voter forms beliefs based on $s$, $\mathbf{p}$, and $\tilde{\mu}$ and then votes to maximize (1).
We are now in a position to define equilibrium of our game. (i) each voter maximizes $AU(s,\mathbf{p},\tilde{\mu})$ given a belief about the probability that policy 1 will be enacted $\kappa$; (ii) each voter votes sincerely given $s$ and $\tilde{\mu}$; (iii) $\kappa$ is correct given $s$, $\tilde{\mu}$, and given the strategies of all other voters; (iv) politicians choose policy platforms that maximize the probability to be elected, given their expectations about voters’ behavior. As typical in signaling games, there are multiple equilibria. I focus attention on symmetric equilibria, that is, equilibria in which candidates, holding the same information and trying to achieve the same, choose identical strategies. Further, if on the equilibrium path both candidates choose the same pure strategy, then we must have $p_{1}=p_{2}$. If, off the equilibrium path, voters observe $p_{1}\neq p_{2}$, I assume that they do not learn anything from this about $\omega$. That is, if off the equilibrium path $p_{1}\neq p_{2}$, then observing $\mathbf{p}$ does not change voters’ belief about $\omega$ [^2] This could be justified, for example, because candidates have types as well. An ideological candidate always chooses $p_{i}=i-1$, $i\in\{0,1\}$. A strategic candidate chooses the policy that is expected to maximize the probability to win the election. A model in which each candidate $i$ is ideological with probability $\gamma\in(0,1)$ and non-ideological with probability $1-\gamma$ would yield the same results regarding the platform choices of office-motivated candidates as the assumption of non-informative deviations.
3 Equilibrium
In this section, I solve the baseline game. I begin by analyzing equilibrium play in the voting subgame. In Section 3.2, I analyze candidates’ incentives and equilibrium behavior, and in Section 3.3 I study the equilibrium of the whole game.
3.1 Voter Behavior
First I turn to the optimal motivated beliefs of voters. The focus is on situations in which $p_{1}\neq p_{2}$, because otherwise voters choose a winner, but not a winning policy. If $p_{1}\neq p_{2}$ off the equilibrium path, $\mathbf{p}$ is not informative about $\omega$. However, it may be so on the equilibrium path. In a symmetric equilibrium, $p_{1}\neq p_{2}$ is only possible in mixed strategies. Because both candidates hold the same information, assume that, when the state is $\omega$, each chooses $p=1$ when with probability $\rho_{\omega}$ and $p=0$ with probability $1-\rho_{\omega}$. Therefore, if $\rho_{1}=1$ and $\rho_{0}=0$, then both choose the policy that matches the state with probability 1. If $\rho_{0}=\rho_{1}=0$, then candidates never choose $p=1$, irrespective of $\omega$.
If $p_{1}\neq p_{2}$ off-equilibrium, $\mathbf{p}$ is not informative by assumption, and $p_{1}\neq p_{2}$ in a symmetric pure strategy equilibrium is not possible. Hence, the belief about $\omega$ of a voter who receives signal $s$ and who chooses $\tilde{\mu}$, is
$\pi(s,\mathbf{p},\tilde{\mu})=\dfrac{q\rho_{1}(1-\rho_{1})}{q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}}.$ (3)
To see how a voter optimally chooses her motivated belief $\tilde{\mu}$, take the derivative of $AU(s,\mathbf{p},\tilde{\mu})$ with respect to $\tilde{\mu}$:
$\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}=-\frac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[\beta+\Delta-\kappa(1+\beta)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}$ (4)
If $\kappa=\tilde{\kappa}\equiv\frac{\beta+\Delta}{\beta+1}$, then this is equal to zero, independent of $s$. If $\Delta>1$, then $\tilde{\kappa}>1$, and because $\kappa$ is a belief, it must be true that $\kappa<\tilde{\kappa}$, implying that the sign of (4) is independent of $\kappa$. In particular, if this is the case, then $AU(s,\mathbf{p},\tilde{\mu})$ decreases in $\tilde{\mu}$ if $s>0$, and it increases in $\tilde{\mu}$ if $s<0$. Otherwise, that is, if $\Delta\leq 1$, then whether or not $AU(s,\mathbf{p},\tilde{\mu})$ increases in $\tilde{\mu}$ depends on both $s$ and the belief $\kappa$. Note that for each voter $\kappa$ can be interpreted as a measure of trust in government. $\kappa$ is the distribution of beliefs about the policies that will be enacted. If a voter believes that the government will always enact the policy that matches the true state, then $\kappa=\pi(s,\mathbf{p},\tilde{\mu})$. If, however, there is little trust in government regarding policy choices, then $\kappa\neq\pi(s,\mathbf{p},\tilde{\mu})$, and if trust is very low, we have $\kappa\in\{0,1\}$.
We can now derive the optimally distorted complexity of the issue climate change:
Lemma 1.
Let $\tilde{\kappa}\equiv\frac{\beta+\Delta}{\beta+1}$.
  1. (a)
    If $s=0$ or $\kappa=\tilde{\kappa}$, then generally $\tilde{\mu}^{*}=\mu$.
  2. (b)
    If $\Delta>1$, then a voter’s optimal distortion of $\mu$ is (i) $\tilde{\mu}^{*}=\infty$ if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
  3. (c)
    If $\Delta=1$, then a voter’s optimal distortion of $\mu$, if $\kappa<1$, is (i) $\tilde{\mu}^{*}=\infty$, if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
  4. (d)
    If $\Delta<1$, then a voter’s optimal distortion of $\mu$ is (i) $\tilde{\mu}^{*}=\infty$, if either $s<0$ and $\kappa<\tilde{\kappa}$, or $s>0$ and $\kappa>\tilde{\kappa}$, and (ii) $\tilde{\mu}^{*}=0$, if either $s<0$ and $\kappa>\tilde{\kappa}$, or $s>0$ and $\kappa<\tilde{\kappa}$.
The lemma provides us with an important intermediate result, the optimal distortions of $\mu$ for different voters. We see that beliefs about policy only matter if severe climate change causes only moderate baseline welfare losses. Otherwise, if severe climate change leads to catastrophical welfare losses, then any signal indicating that $\omega=1$ will be interpreted as pure noise and hence completely ignored. To the contrary, any signal indicating $\omega=0$ will be accepted as a perfect indication that climate change is indeed mild, independent of the signal’s strength. Finally, a voter with a signal $s=0$ or a voter with $\kappa=\tilde{\kappa}$ has no incentive to distort her signal’s precision and hence updates her belief about $\omega$ like a perfect Bayesian.
What does this imply for the beliefs about $\omega$ the voters hold? Clearly, if $\tilde{\mu}=0$, then beliefs do not change at all, and $\pi(s,\mathbf{p},0)=q$. The same is true for a voter receiving a signal $s=0$, which is not informative about $\omega$. Moreover, if $\tilde{\mu}=\infty$, then any signal that contains only the slightest bit of information will completely move beliefs to the extremes, and hence $\pi(s,\mathbf{p},\infty)\in\{0,1\}$. Only if $s\neq 0$ and $\kappa=\tilde{\kappa}$ will the belief be a non-constant continuous function of $s$ and it equals $\pi(s,\mathbf{p},\mu)$:
Corollary 1.
Let $\hat{\pi}\equiv\frac{q\rho_{1}(1-\rho_{1})}{q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})}$.
  1. (a)
    If $\Delta>1$, then
    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}\hat{\pi}&\text{if}s\geq 0\\ 0&\text{if }s<0\end{array}\right.$
  2. (b)
    If $\Delta\leq 1$,
    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}0&\text{if }\left(s<0\wedge\kappa<\tilde{\kappa}\right)\\ \hat{\pi}&\text{if }\left(s>0\wedge\kappa<\tilde{\kappa}\right)\vee\left(s<0\wedge\kappa>\tilde{\kappa}\right)\\ \pi(s,\mathbf{p},\mu)&\text{if }\kappa=\tilde{\kappa}\\ 1&\text{if }\left(s>0\wedge\kappa>\tilde{\kappa}\right)\end{array}\right.$
We now know the beliefs of all voters as functions of their signals and of their policy belief $\kappa$. Corollary 1 shows that if $\Delta>1$, and hence severe climate change has catastrophic consequences, then the equilibrium has a simple structure, and $\kappa$ actually plays no roll. Voters have two different beliefs, $\pi(s,\mathbf{p},\tilde{\mu})=0$ and $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$. What does this imply for voters’ decisions at the ballot? Recall that voters vote sincerely for the alternative that they believe maximizes (1). The expected utility from policy 1 is $u(p=1)=-\pi(s,\mathbf{p},\tilde{\mu})\Delta-(1-\pi(s,\mathbf{p},\tilde{\mu}))$, while from policy 0 she gets $u(p=0)=-\pi(s,\mathbf{p},\tilde{\mu})(\Delta+\beta)$. Hence, the voter cast her ballot for policy 1 iff
$u(p=1)>u(p=0)\Leftrightarrow\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}\equiv\frac{1}{1+\beta}.$ (5)
If the reverse is true, then policy 0 is strictly preferred, and if $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$, then a voter is indifferent. Clearly, $01>\tilde{\pi}>$ 0. Note that by Assumption 1 we have $\tilde{\pi}>q$.
If $\Delta>1$, then voters hold two different beliefs, either $\pi(s,\mathbf{p},\tilde{\mu})=0$ or $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$. A voter with the former belief always votes for policy 0, while the decision at the ballot of the other voter depends on $q$, $\beta$, as well as on the politicians’ strategies $\rho_{0}$ and $\rho_{1}$. If $\hat{\pi}>\tilde{\pi}$, then these voters vote for policy 1, and as a consequence all voters vote informatively, i.e., they vote according to their signals. This means that in each state $\omega$ the policy that matches this state is chosen. If, however, $\hat{\pi}\leq\tilde{\pi}$, then a majority of voters always supports policy 0, implying it wins independent of the state.
Things are slightly different when $\Delta\leq 1$. On the one hand, the equilibrium just described continues to exist also when severe climate change decreases baseline welfare only moderately. In particular, this is the case if $\hat{\pi}\leq\tilde{\pi}$ and $\kappa<\tilde{\kappa}$ for sufficiently many $s$. To the contrary, if $\hat{\pi}\geq\tilde{\pi}$ and $\kappa>\tilde{\kappa}$ for sufficiently many $s$, then policy 1 wins independent of the state. Hence, when severe climate change is not catastrophic, then there might exist an equilibrium in which politicians always campaign on the policy tailored for severe climate change, independent of the true state. It follows that we cannot yet exclude that, when $\Delta\leq 1$, a second inefficient equilibrium may exist.
What about the efficient equilibrium, in which candidates choose the optimal policies given $\omega$? Assume each voters wants to believe her signal and thus has $\kappa<\tilde{\kappa}$ when $s<0$ and $\kappa>\tilde{\kappa}$ when $s>0$. Then every voter votes informatively, and therefore the policy matching the state wins for all $\hat{\pi}\in(0,1)$. Hence, if even severe climate change decreases baseline welfare only moderately, then there always may exist an equilibrium in which voters vote informatively, and hence the correct policy is chosen with probability 1.
We summarize the results of this section in our next proposition:
Proposition 1.
Assume $\mathbf{p}\in\left\{\left(0,1\right),\left(1,0\right)\right\}$.
  1. 1.
    If $\Delta>1$, then there is a unique subgame equilibrium.
    1. (a)
      If $\hat{\pi}\leq\tilde{\pi}$, then $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ for all $s$ and a majority of voters always votes for policy 1.
    2. (b)
      If $\hat{\pi}>\tilde{\pi}$, then $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$ and $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=1$ if $s\geq 0$, and a majority always votes for the policy $p_{i}$ that matches the true state $\omega$.
  2. 2.
    If $\Delta\leq 1$, then there exist multiple subgame equilibria.
    1. (a)
      For any $\hat{\pi}\in(0,1)$, there exists a subgame equilibrium with $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$ and $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=1$ if $s\geq 0$, and a majority of voters always votes for the policy $p_{i}$ that matches the true state $\omega$.
    2. (b)
      For $\hat{\pi}\geq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=1$ and policy 1 always wins.
    3. (c)
      For $\hat{\pi}\leq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=0$ and policy 0 always wins.
    4. (d)
      There is no subgame equilibrium in which the policy $p_{i}$, that does not match the state $\omega$, always wins.
On the positive side, there always exists a subgame equilibrium in which information aggregates, if severe climate change decreases baseline welfare not too much, i.e., if $\Delta$ is small. However, this equilibrium depends on the expectations about the efficiency of political process. If voter believe that the policy suggested by their signal is chosen, then the election leads to efficient results. However, if voters are pessimistic, and believe policy 0 will be chosen no matter what is the true state $\omega$, then this becomes a self-fulfilling prophecy if $\hat{\pi}$ is low. When $\Delta$ becomes larger, and hence severe climate change leads to more dire welfare losses, then the efficient equilibrium disappears when $\hat{\pi}$ is small.
3.2 Candidates
We now study the platform choices of the candidates as a function of their signals. Of course, when they decide which policies to offer they think ahead which policies are likely to lead to electoral success, and therefore voters’ reactions, in particular the one’s formalized in Proposition 1, matter for candidates’ incentives.
From the perspective of an office motivated candidate, the optimal platform is the one that maximizes the chance to win the election. In a symmetric equilibrium, if both candidates choose identical platforms, voters may learn from the platforms’ congruence something about the true state $\omega$, but they cannot choose policies anymore, since all candidates offer the same. Because voters only care about policies, both candidates win with an equal probability of 50 percent if $p_{1}=p_{2}$. If some candidate deviates and chooses off-equilibrium a platform different from his opponent, voters learn nothing about $\omega$ from this deviation. However, if equilibrium play permits $p_{1}\neq p_{2}$, then platforms may indeed be partially informative and change voters’ beliefs, as discussed before.
First consider potential mixed-strategy equilibria. Focussing on symmetric equilibria, the probability that each candidate choose policy 1 in state $\omega$ is $\rho_{\omega}$ as before. Note that if $\rho_{1}\neq\rho_{2}$, voters learn from platform choices even if $p_{1}\neq p_{2}$. Choosing such a mixed strategy can only be an equilibrium if it leads to a chance of winning of 50%, because candidates need to be indifferent. The probability to win is 50%, if either both receive exactly half the votes, or if the probability that a majority chooses either candidate is 50%. But note that candidates know the state $\omega$, and thus they can infer the exact distribution of signals voters receive. Therefore, in a mixed-strategy equilibrium we need to have that both receive exactly half the votes. This implies policy 1 is implemented with probability of 50%, independent of the state $\omega$. Hence, we need to have $\kappa=\frac{1}{2}$ for all $s$. If $\tilde{\kappa}<\frac{1}{2}$, any voter with $s>0$ will hold belief $\pi(s,\mathbf{p},\tilde{\mu})=1$ and thus they all vote for 1. But this means that in state $\omega=1$, the candidate offering $p=1$ wins with certainty, and thus this cannot be equilibrium. If $\tilde{\kappa}>\frac{1}{2}$, any voter with $s<0$ will hold belief $\pi(s,\mathbf{p},\tilde{\mu})=0$ and thus they all vote for policy 0. But this means that in state $\omega=0$ the candidate offering $p=0$ wins with certainty, and thus this cannot be equilibrium, either. Finally, if $\tilde{\kappa}=\frac{1}{2}$, any voter will choose $\tilde{\mu}^{*}(s)=\mu$. But then information aggregates by Assumption 2, and thus this cannot be an equilibrium belief, either. Hence, we can conclude that no mixed strategy equilibrium can exist.
What about pure strategy equilibria? In a symmetric equilibrium, we must have $p_{1}=p_{2}$, because candidates hold the same information. Suppose voters observe off-equilibrium that $p_{1}\neq p_{2}$ and both candidates expect voters to vote informatively. Hence, they vote for policy 1 if $s>0$, for policy 0 if $s<0$, and they randomize when $s=0$. Then, in each state $\omega$, the candidate offering the policy that matches the state wins the election. Hence, when choosing which policy platform to offer, candidate $i$ knows that $p_{i}=\omega$ wins the election if $p_{-i}\neq\omega$ and otherwise both win with a probability of 50%. To the contrary, choosing $p_{i}\neq\omega$ wins the election with 50% if also $p_{-i}\neq\omega$ and otherwise loses the election for sure. Hence, $p_{i}=s_{i}$ is the unique optimal strategy. It follows that if sufficiently many voters vote informative, candidates will choose informative policy platforms. To the contrary, if candidates expect a sufficiently large number of voters to not vote informatively when $p_{1}\neq p_{2}$, then there exists a policy $\hat{p}\in\{0,1\}$ that wins the election with certainty, independent of the state $\omega$. Candidates then know they win the election by choosing $p_{i}=\hat{p}$ if $p_{-i}\neq\hat{p}$, and they win with a probability of 50% if $p_{-i}=\hat{p}$. Choosing $p_{i}\neq\hat{p}$ will lose the election for sure if $p_{-i}=\hat{p}$, and hence both candidates are best served by offering $\hat{p}$.
This leads to the next formal result:
Proposition 2.
There is a unique symmetric equilibrium in the platform choice stage, which is in pure strategies. Candidates choose $p_{i}^{*}=\omega$, if they expect a majority of voters to vote for the policy that matches the state, when $p_{1}\neq p_{2}$. Otherwise, there exists a policy $\hat{p}$ that is expected to win independent of the realization of $\omega$ if $p_{1}\neq p_{2}$, and both candidates choose $p_{i}^{*}=\hat{p}$.
3.3 Equilibrium Policy Platforms
We can now determine equilibrium platform choices as functions of the parameters of the game. If severe climate change leads to catastrophic baseline welfare level losses, $\Delta>1$, then voters ignore any signal indicating that $\omega=1$. Our analysis so far reveals that then a majority always votes for policy 0, whenever $p_{1}\neq p_{2}$. Proposition 2 tells us that in this case candidates never campaign on policy $p=1$, independent of the information they hold. Hence, if severe climate change is catastrophic, then candidates will ignore any information and always choose the optimal policy for mild climate change, $\mathbf{p}^{*}=(0,0)$.
When $\Delta\leq 1$, the set of equilibria becomes larger. The reason is that now voters’ beliefs about the enacted policy become self-fulfilling. If voters trust the government in the sense that they believe the policy that is optimal given $\omega$ wins, then voters always vote informatively, implying that $\mathbf{p}^{*}=(\omega,\omega)$ can indeed be equilibrium. However, there always coexists another equilibrium, in which candidates again ignore their information. If voters are convinced that $p=0$ is chosen, then any voter with $s<0$ will choose $\tilde{\mu}=\infty$, while a voter with $s>0$ chooses $\tilde{\mu}=0$. Therefore, as in the case of $\Delta>1$, a majority always votes for policy 0, and candidates choose $\mathbf{p}^{*}=(0,0)$.
The next proposition is the paper’s main result and formalizes the above intuitions:
Proposition 3.
If $\Delta>1$, then there exists a unique pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ and candidates choose $\mathbf{p}^{*}=(0,0)$. If instead $\Delta\leq 1$, then there are multiple equilibria:
  1. 1.
    If $\Delta\leq 1$ and voters have trust in the government, then there exists a pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=\pi(s,\mathbf{p},\tilde{\mu}^{*})$ and candidates choose $\mathbf{p}^{*}=(\omega,\omega)$.
  2. 2.
    If $\Delta\leq 1$ and voters have no trust in the government, then there exists a pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
4 Extensions
In this section I informally discuss further extensions that seem relevant. For the sake of brevity and clarity, throughout this section I assume that $\Delta>1$. In Section 4.1 I discuss how conclusions change if candidates are not only office motivated, but care also about welfare. In Section 4.2 I discuss the implications of differences in candidate valence. In Section 4.3, I discuss the implication of misinformation campaigns. Finally, in Section 4.4 I study the implication of costly deviations from Bayesian rationality.
4.1 Policy Motivated Politicians
Suppose candidates not only care about being elected but also value welfare, as captured by (1). Knowing which state has materialized, candidates are aware of which policy maximizes welfare. Thus, one might hypothesize that adding a policy motivation would improve the situation. However, this is only partially correct. On the one hand, if policy motivation is sufficiently important compared to simply being elected to office, there is indeed an equilibrium in which candidates always choose $\mathbf{p}^{*}=(\omega,\omega)$. To see this, assume the spoils of office have value $V$ and that each candidate receives $V+u(p,\omega)$ if elected and $u(p,\omega)$ otherwise.
Now, suppose candidate 1 expects candidate 2 to choose $p_{2}=\omega=1$. Choosing $p_{1}=0$ would win the election but result in a welfare loss due to the wrong policy being chosen. Alternatively, choosing $p_{1}=1$ reduces the chance of winning the election to 50%, but it ensures that the optimal policy is selected with certainty. Candidate 1 has a strict incentive to also propose $p_{1}=1$ if
$\frac{V}{2}-\Delta\geq V-\Delta-\beta\Leftrightarrow\beta\geq\frac{V}{2}.$
If, however, $\beta<\frac{V}{2}$, there remains a unique pure strategy equilibrium with $\mathbf{p}^{*}=(0,0)$.
Thus, if policy motivation is sufficiently strong, the previously identified problems may not occur. However, there is always another equilibrium in pure strategies where candidates continue to ignore their information despite policy motivation. Assume candidate 1 expects candidate 2 to choose $p_{2}=0$. In this case, the unique best response is to also choose $p_{1}=0$, because policy 0 now wins with certainty. Although candidate 1 cannot influence the policy, she can increase her chances of winning. Therefore, despite policy motivations, and even in the absence of any office motivation, there always exists an equilibrium with $\mathbf{p}^{*}=(0,0)$.
4.2 Valence Differences
Next, consider valence differences as in Ansolabehere and Snyder (2000), Groseclose (2001), Aragones and Palfrey (2002), and more recently Denter (2021). Specifically, assume that voters evaluate candidates not only by the policies they offer but by $v_{i}+u(p_{i},\omega)$, where $v_{i}\in\mathbb{R}$ is the valence utility the voter gets from candidate $i$ in the office and $u(p,\omega)$ is policy utility and defined as before in (1). Let $v_{1}>v_{2}$, implying that candidate 1 has a valence advantage. How does introducing valence alter the model’s conclusions? On the one hand, it has no effect on $\tilde{\mu}^{*}(s)$, which remains as described in Lemma 1, because valence is independent of the state $\omega$. However, valence does affect a voter’s choice at the ballot box for a given belief $\pi(s,\mathbf{p},\tilde{\mu})$. If both candidates choose the same policy, candidate 1 always wins because $v_{1}>v_{2}$. Equilibrium in the extended game resembles equilibrium in Ansolabehere and Snyder (2000). If the valence advantage is not too large, candidate 1 will always choose $p_{1}=0$, while candidate 2 may choose any $p_{2}\in\{0,1\}$ without having a chance to win. If the valence advantage is large enough, candidate 1 may choose any $p_{1}\in\{0,1\}$ and still win with certainty. Thus, it is possible that a candidate with a large valence advantage might use this advantage to propose an optimal policy. In fact, if we combine a large valence advantage with policy motivation as described in Section 4.1, then it is a dominant strategy for candidate 1 to choose $p_{1}=\omega$, and this policy indeed wins the election.
4.3 Disinformation Campaigns
The analysis so far has revealed a robust mechanism through which the electoral incentives of candidates are sufficient to prevent the implementation of adequate policies against climate change. However, climate change is also one of the topics where organized misinformation is most prevalent (see, for example, Lewandowsky, 2021). Perhaps the most famous example of such campaigns is that of ExxonMobil, which, despite having very precise knowledge about the causes and consequences of climate change (see Supran, Rahmstorf, and Oreskes, 2023), chose to downplay the phenomenon in their public communications and editorial-style advertisements in The New York Times (see Supran and Oreskes, 2017).
I follow Papanastasiou (2020), Sisak and Denter (2024) and Denter and Ginzburg (2024) to incorporate misinformation into the model. In particular, I assume that each voter receives a fake signal with a probability of $\alpha$, while with the complementary probability of $1-\alpha$ the signal comes from a proper source and is informative. An informative signal is drawn from the distribution $F_{\omega}$ with corresponding density $f_{\omega}$. I assume that $f_{1}(s)/f_{0}(s)$ is an strictly increasing function of $s$, which is the standard MLRP assumption. Note that in the model described in Section 2 with normally distributed signals this is automatically satisfied. Moreover, without loss of generality let $f_{1}(0)=f_{0}(0)$, implying that $s=0$ is an uninformative signal as before. Unlike proper signals, a fake signal is not informative. In particular, I assume that the distribution of fake signals is $G(s)$ which is independent of $\omega$. For simplicity, further assume that $F_{\theta}$ and $G$ have full support on $\mathbb{R}$.
Because the literature has shown that in the presence of fake news motivated reasoning is often regarding the reliability of an information source, voters cannot change the precision of a signal anymore, but have motivated beliefs about the probability that a given signal is fake or proper. In particular, they can choose $\tilde{\alpha}\in[0,1]$ instead of $\tilde{\mu}\in[0,\infty]$.
For a given signal $s$ and a given choice $\tilde{\alpha}$, the belief about the state $\omega$ is
$\pi(s,\mathbf{p},\tilde{\alpha})=\frac{q\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right]}{q\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right]+(1-q)\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right]}.$
Choosing $\hat{\alpha}=0$ means ignoring the chance that the signal could be fake, and hence the belief equals the perfect Bayesian belief when no fake signals exist. To the contrary, setting $\hat{\alpha}=1$ means ignoring the signal completely because it originates from a fake source.
Lemma 2.
The optimal motivated belief about $\alpha$ is
$\hat{\alpha}^{*}(s)=\left\{\begin{array}[]{cl}1&\text{if }s>0\\ \alpha&\text{if }s=0\\ 0&\text{if }s<0\end{array}\right.$
Hence, any bad news about climate change will be interpreted as fake, while any good news will be interpreted as proper and hence credible. This is in line with recent empirical evidence (Thaler, 2024).
What are the consequences for political decision making? The political equilibrium is the same as in Section 3. The reason is that any signal indicating climate change is severe will be ignored, and hence a majority of the electorate will prefer policy 00 in each state.
Proposition 4.
Suppose $\Delta>1$. There exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\alpha}^{*}(s))=0$, where $\tilde{\alpha}^{*}(s)$ follows from Lemma 2, and in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
4.4 Costly Distortions
In this section, I extend the previous analysis by accounting for the costs that a voter may incur when distorting the issue’s complexity. The goal is twofold: first, introducing costs allows us to analyze how beliefs may be distorted, which is valuable in itself. Second, it enables us to assess the robustness of the earlier results. I begin by considering a fixed cost of distortion, following the approach of other recent papers, such as Levy (2014) and Le Yaouanq (2023). I then explore a continuous cost function, where the magnitude of the distortion affects the costs, in Section 4.4.2. I will concentrate on conditions under which equilibria with inactive candidates continue to exist.
4.4.1 Fixed Costs
First consider that distorting the issue’s complexity implies a fixed cost of $\gamma>0$. That is, each voter now maximizes
$W(s,\mathbf{p},\tilde{\mu})=AU(s,\mathbf{p},\tilde{\mu})-\gamma\cdot\mathbbm{1}_{\tilde{\mu}\neq\mu}.$
As before, $W(s,\mathbf{p},\tilde{\mu})$ strictly decreases in $s$, implying the greatest incentive to distort $\mu$ when $s>0$ has a voter receiving $s\rightarrow\infty$, and when $s<0$ a voter receiving $s\approx 0$. Of course, if $\gamma$ is too large, no voter has an incentive to choose $\tilde{\mu}\neq\mu$.
Lemma 3.
If $\gamma<\gamma^{+}\equiv(1-q)(\beta(1-\kappa)+\Delta-\kappa)$, then there exists $s^{+}>0$ such that
$\tilde{\mu}^{*}=\left\{\begin{array}[]{cl}0&\text{if }s\geq s^{+}\\ \mu&\text{if }s\in[0,s^{+})\end{array}\right.$
However, if $\gamma\geq\gamma^{+}$, then $\tilde{\mu}^{*}=\mu$ for all $s\geq 0$. Moreover, if $\gamma<\gamma^{-}\equiv q(\beta+\Delta-(\beta+1)\kappa)$, then there exists $s^{-}<0$ such that
$\tilde{\mu}^{*}=\left\{\begin{array}[]{cl}\infty&\text{if }s\in[s^{-},0)\\ \mu&\text{if }s<s^{-}\end{array}\right.$
However, if $\gamma\geq\gamma^{-}$, then $\tilde{\mu}^{*}=\mu$ for all $s<0$.
Having established voters’ optimal deviations from Bayesian rationality, we next seek to determine under which conditions inactivity against climate change, despite better information, remains an equilibrium. First, note that if $\gamma\geq\max\left\{\gamma^{-},\gamma^{+}\right\}$, then no voter distorts their beliefs, and all voters interpret information as perfect Bayesians. It follows from Assumption 2 that in this case, information aggregates correctly in an election with $p_{1}\neq p_{2}$, and therefore the conditions for efficient policy platforms in equilibrium are satisfied. If $\gamma<\min\left\{\gamma^{-},\gamma^{+}\right\}$, there will be both voters who received good news ($s<0$) and those who received bad news ($s>0$) who choose $\tilde{\mu}^{*}\neq\mu$, while others will interpret signals correctly.
Proposition 5.
Suppose $\Delta>1$. There exists $\hat{\gamma}>0$ such that iff $\gamma<\hat{\gamma}$, then there exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, where $\tilde{\mu}^{*}$ follows from Lemma 3, and in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
The proposition shows that our earlier analysis is robust to the introduction of a fixed cost that needs to be paid if a voter chooses $\tilde{\mu}\neq\mu$. The result is of course not very surprising given Proposition 3. However, one thing to note is that the incentives to distort $\mu$ of voters with $s<0$ do not matter for the result. In fact, those voters vote more often informatively with motivated beliefs than without, which is beneficial for choosing the optimal policy.
4.4.2 Quadratic Costs
Next consider the case of a smooth cost of distortion function. We demand a number of properties our cost function should satisfy. It should be convex increasing when we move away from $\mu$ and zero when there is no distortion. Moreover, it should not be symmetric in the sense that $\tilde{\mu}=\mu+\epsilon$ and $\tilde{\mu}=\mu-\epsilon$ should not lead to the same cost. To see why assume $\mu=1$. A symmetric cost function would imply that $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=0}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=2}$. But note that $\tilde{\mu}=0$ implies that the issue is perceived as infinitely incomplex, and any signal $s\neq 0$ will lead to updating to either 0 or 1. At the same time, $\tilde{\mu}=2$ represents only a gradual increase in complexity. If both have the same cost, then downward distortions are generally “cheaper” than upward distortions. We therefore demand that the cost of equal relative distortions upwards and downwards are the same. This implies the cost function is compressed for downwards distortions. More formally, $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\lambda\cdot\mu}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\lambda}$ for all $\lambda>0$ and all $\mu>0$. In particular, we assume the following cost function:
$C(\tilde{\mu},\mu)=\left\{\begin{array}[]{ll}\frac{c}{2}\left(\tilde{\mu}-\mu\right)^{2}&\text{if }\tilde{\mu}\geq\mu\\ \frac{c}{2}\left(\frac{\mu}{\tilde{\mu}}\left(\tilde{\mu}-\mu\right)\right)^{2}&\text{if }\tilde{\mu}<\mu\end{array}\right.$ (6)
With this cost function we have $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\lambda\cdot\mu}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\lambda}=\frac{c}{2}\mu^{2}\left(\lambda-1\right)^{2}$ for all $\lambda>1$ as desired. In the left panel of Figure 1, the cost function is plotted for $\mu=1$ and $c=2$. One property of it is that $\lim_{\tilde{\mu}\rightarrow\infty}C(\tilde{\mu},\mu)=\lim_{\tilde{\mu}\rightarrow 0}C(\tilde{\mu},\mu)=\infty$, and therefore we have $\tilde{\mu}\in(0,\infty)$ for any $s$.
When deciding how to optimally distort $\mu$, each voter aims to maximize
Figure 1: Left panel: Cost function when $\mu=1$ and $c=2$. Right panel: $\tilde{\mu}$ as a function of $s$ for $\mu=1$ and different values of $c$. Higher $c$ moves $\tilde{\mu}$ closer to $\mu$.
$W(s,\mathbf{p},\tilde{\mu})=AU(s,\mathbf{p},\tilde{\mu})-C(\tilde{\mu},\mu).$
As before, focus on $p_{1}\neq p_{2}$. When the signal is uninformative, $s=0$, any distortion only causes cost, but does not change $AU(s,\mathbf{p},\tilde{\mu})$. Hence, in this case, $\tilde{\mu}^{*}=\mu$. But what happens when signals are very informative? Suppose $s\rightarrow\infty$. Then any distortion except $\tilde{\mu}=0$ will move the belief about $\omega$ to $\pi=1$. But the cost of choosing $\tilde{\mu}=0$ is infinite, and hence this cannot be optimal. Therefore, it is optimal to not distort at all. The same is true for $s\rightarrow-\infty$. For any other signal realization, $s\in\left(-\infty,0\right)\cup\left(0,\infty\right)$, the direction of the optimal distortion is the same as in Lemma 1. Moreover, $W(s,\mathbf{p},\tilde{\mu})$ is a continuous and smooth function of both $s$ and $\tilde{\mu}$, and hence also $\tilde{\mu}$ is a continuous and smooth function of $s$. To get a better intuition of how $\tilde{\mu}^{*}$ changes with $s$, I plotted it for $\mu=1$ and different values of $c$ in the right panel of Figure 1. The next lemma states the above intuitions formally:
Lemma 4.
Suppose $\Delta>1$. If $s\in\{-\infty,0,\infty\}$, then $\tilde{\mu}(s)=\mu$. Otherwise, and if $s>0$, then $\tilde{\mu}(s)<\mu$, and there exists $\bar{s}>0$ such that $\tilde{\mu}(s)$ decreases in $s$ if $s\in[0,s^{+})$, and it increases in $s$ if $s>s^{+}$. If $s<0$, then $\tilde{\mu}(s)>\mu$, and there exists $s^{-}<0$ such that $\tilde{\mu}(s)$ decreases in $s$ if $s\in(s^{-},0]$, and it increases in $s$ if $s<s^{-}$. Moreover, for every $s$, the absolute distortion $\left|\mu-\tilde{\mu}\right|$ decreases in $c$
Now let us discuss the implications of the lemma for voters’ decisions at the ballot, and by extension, for candidates’ incentives to offer the optimal policy. First, just as in the case of fixed costs, if $s<0$, voters tend to amplify the signals, leading more voters to vote informatively compared to the absence of motivated beliefs. This improves the ability to select the right policy through an election. Therefore, to understand the conditions under which candidates ignore information and choose policy 0 regardless of $\omega$, we need to focus on voters receiving $s>0$. These voters interpret signals as less informative than they truly are, resulting in fewer voters casting informative votes compared to the absence of motivated beliefs. Because $q<\tilde{\pi}$, if either the distortion of $\mu$ is significant or the signal is too weak, such a voter may cast a ballot for policy 0 despite $s>0$. Intuitively, as $c$ increases, voters choose to distort $\mu$ less and less. Consequently, $\lim_{c\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$. In this case, by Assumption 2, a majority will always cast a ballot for $p=\omega$, giving candidates an incentive to heed their information. However, if $c$ is small, we revert to the result from Proposition 3. By continuity, there exists a $\hat{c}>0$ such that if $c\in[0,\hat{c}]$, motivated beliefs will lead to distortions and voting behavior that prevent candidates from following their information.
Proposition 6.
Suppose $\Delta>1$. There exists $\hat{c}>0$ such that iff $c<\hat{c}$, then there exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, where $\tilde{\mu}^{*}$ follows from Lemma 4, in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
4.4.3 Comparing Fixed and Continuous Costs
The analysis in this section has given us more nuanced results for the optimal distortion of beliefs. Moreover, in both extensions discussed so far, the inactive equilibrium prevails as the unique equilibrium if the cost of distorting informativeness of signals is low enough. However, the two models discussed differ significantly in how voters choose to distort information and hence beliefs, see Figure 2, allowing to derive testable predictions.
Figure 2: Equilibrium belief distortion (solid, red curve) with fixed costs (left panel) and with a strictly convex cost function (right panel). The blue and dashed curve is the Bayesian belief.
With fixed costs, only the worst signals are distorted, both in the case of good news ($s<0$) and bad news ($s>0$). When news are bad, the worst signals are the ones suggesting $\omega=1$ the strongest. This leads to low anticipatory utility and thus the signal is interpreted as not informative. However, if bad news is weak, then the belief does not change a lot anyway, and hence choosing $\tilde{\mu}\neq\mu$ at a cost is not attractive. When news is good, the worst signals are the ones suggesting $\omega=0$ the least. Weak good news leads to little updating of beliefs, and overreacting to information is better for anticipatory utility, even if it is somewhat costly. However, if good news is strong enough, then overreacting to the signal does not change beliefs that much anymore, and hence distorting beliefs is not in a voters interest.
If, instead, we consider a continuous and convex cost function, the results are to some degree reversed. In fact, when $s$ is very large, implying a voter receives strong bad news, then interpreting the signal as not informative is very costly, and hence a voters will be better served simply interpreting the signal in an (almost) Bayesian fashion. The same is true when $s$ is close to zero. However, bad news of intermediate strength will be downplayed significantly. When a voter receives good news we find similar results. If $s$ is very low, the voter’s belief about $\omega$ is close to zero even if $\tilde{\mu}=\mu$, and hence distorting $\mu$ is not very attractive. The same is true if good news is very weak. However, if good news is of intermediate strength, then distorting $\mu$ upwards is attractive.
5 Discussion and Some Evidence
The formal analysis highlights at least two important determinants of the efficiency of policy-making in the face of climate change, rhetoric and trust in government.
The Importance of Rhetoric:
When $\Delta>1$, there is no equilibrium in which office-motivated candidates campaign with the policies needed for severe climate change, even if they know for sure that $\omega=1$. Of course, in real life, we do not discuss the issue of climate change in terms of baseline welfare losses, but rather using verbal descriptions. In recent years, an increasing number of people have stopped using the relatively neutral terms “global warming” or “climate change” and have instead adopted more extreme and loaded language such as “climate crisis” or “climate emergency.” Indeed, a comparison of Google search trends in the U.S. from January 2014 to July 2024 shows that, starting around 2019, the relative frequency with which these terms are used has been increasing, see Figure 3. One reason for this change is that many activist groups started to use a more aggressive rhetoric in an effort to emphasize the importance of swift climate policy changes. For example, the activist group Extinction Rebellion writes on their homepage: [^3] See https://rebellion.global/ (accessed on 24/09/2024).
(a) Climate Change
(b) Climate Emergency
(c) Climate Crisis
Figure 3: Five months moving averages of Google search trends in the U.S. for “Climate Change”, “Climate Emergency”, and “Climate Crisis” from January 2014 to July 2024.
This is an Emergency. Life on Earth is in crisis. Our climate is changing faster than scientists predicted and the stakes are high. Biodiversity loss. Crop failure. Social and ecological collapse. Mass extinction. We are running out of time, and our governments have failed to act.
The model suggests that such a shift in language may have undesired consequences, because it may lead voters to ignore information suggesting that climate change is a problem, whereas information stating the opposite is trusted more than it should. And in fact, the fraction of U.S. Americans who stated in PEW polls that they “personally care a great deal about the issue of climate change” increased between 2016 and 2018 from 36% to 44%, but then dropped to 37% in 2023. At the same time, the fraction of people who stated to either not care too much or to not care at all about the issue first decreased from 26% in 2016 to 22% in 2018, but then increased again to 27% in 2023.444See https://www.pewresearch.org/science/2023/10/25/how-americans-view-future-harms-from-climate-change-in-their-community-and-around-the-u-s/. These changes can be explained by the model, and they create stronger incentives for politicians to ignore their information about the true state of the world when designing policies.
That framing and language may have an important effect on wether people learn or not is well established in psychology. For example, Smith and Mackie (2007) mention the importance of not creating too much anxiety in receivers when the intent is to change their attitudes:
To be successful, fear appeals must arouse just the right amount of anxiety by showing the negative consequences […] that will follow if behavior doesn’t change. The provision of an explicit avenue of action […] eliminates the anxiety by showing how the negative consequences can be avoided.
Smith and Mackie (2007), page 257
Excessively negative rhetoric may create too much anxiety and may hence prevent attitude changes and learning, see also McDonald, Chai, and Newell (2015). The risk to trigger such excessive emotional responses is particularly large when people’s general livelihood seems threatened, which is true in the case of climate change. Writing explicitly about the issue climate change, Clayton, Manning, Krygsman, and Speiser (2017) state that
[t]he ability to process information and make decisions without being disabled by extreme emotional responses is threatened by climate change. Some emotional response is normal, and even negative emotions are a necessary part of a fulfilling life. In the extreme case, however, they can interfere with our ability to think rationally, plan our behavior, and consider alternative actions.
Clayton, Manning, Krygsman, and Speiser (2017), page 16
Other papers that highlight the importance of framing for the efficacy to learn about climate change include Spence and Pidgeon (2010), Jones and Peterson (2017), and Ngo, Poortvliet, and Klerkx (2022). Hence, language or rhetoric may indeed be important factors in explaining voters’ attitudes towards climate change, and therefore may as well be determinants of political inactivity in the face of climate change.
The Importance of Trust in Government:
On the other hand, Trust in the Government to choose appropriate policies for the problems at hand matters. In the model this is captured by $\kappa(s,\mathbf{p},\tilde{\mu})$. If voters belief the correct policy will be chosen, then $\kappa(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\tilde{\mu})$, and in this case an efficient equilibrium exists if $\Delta<1$. However, if voters are pessimistic about politician’s ability or incentives to implement the right policies—$\kappa(s,\mathbf{p},\tilde{\mu})$ is then independent of the signal $s$—then this becomes self-confirming, leading politicians to ignore their information, and the optimal policy may not be chosen. Hence, the model suggests that there is a positive relationship between efforts against climate change and trust in government. Note that in a model with no motivated beliefs, such a prediction could not emerge, because the belief about climate change is independent of trust in government, and hence so is the implemented policy. It is the interaction of trust in government $\kappa$ and motivated beliefs $\tilde{\mu}$ that causes this result.
Of course, trust in government is not the only determinant of progress against climate change. Different countries face different costs of adopting new policies. In the model, this is captured by $\beta$. When $\beta$ is large, inaction may be very costly, while if $\beta$ is small, inaction has little consequences. Empirically, $\beta$ could be interpreted as economic cost of (not) dealing with climate change in an appropriate way. Often, anti-climate change policies aim at reducing CO2 emissions. This may be quite costly for countries rich in oil and gas, while it is less costly for other countries. Hence, countries rich in fossil energy sources are less likely to make progress against climate change than countries with little or no such resources.
Statistic N Mean St. Dev. Min Max
CCPI 19 53.66 13.59 26.47 79.61
TGI 19 51.33 13.07 25.63 77.54
Oil 19 9.95 38.82 0.00 170.00
Risk 19 2.20 1.15 1.00 5.40
Table 1: Summary statistics of included variables.
We can test the hypotheses using cross-sectional data. As the dependant variable, I use the Climate Change Performance Index (CCPI) published jointly by Germanwatch, the NewClimate Institute, and the Climate Action Network International. [^5] See https://ccpi.org/. The CCPI takes values between 0 and 100, where higher values indicate that a country performs better in enacting appropriate policies against climate change. The considered policies include reductions in greenhouse gas emissions (40%), expansion of renewable energy sources (20%), general energy use (20%), and other climate policies (20%). As controls for trust in government I use the OECD’s Trust in Government Indicator (TGI). [^6] See https://www.oecd.org/en/data/indicators/trust-in-government.html The TGI also takes values between 0 and 100 and equals the share of people reporting to have confidence in the national government. As additional control variables, I use countries’ confirmed oil reserves in billion barrels (Oil), which can be interpreted as a measure of the cost of climate change policies, and the Global Climate Risk Index (Risk), published by Germanwatch, which can be interpreted as a measure of the benefits of climate change policies. [^7] See https://www.germanwatch.org/en/cri In Table 1 I summarize the included variables.
Figure 4: Scatterplot of Climate Change Performance Index (CCPI) in 2023 and the OECD Trust in Government Indicator (TGI) from 2022 with OLS trendline.
Unfortunately, the number of countries for which all of these variables are available is small ($N=19$), and hence power is limited. Nevertheless, let us now take a first look at the data. Figure 4 shows a scatter plot of CCPI and TGI for the different countries including a trend line. The plot indicates a positive relation between TGI and CCPI, as predicted by the model. The Scandinavian countries have the highest trust in government and also show some of the highest CCPI scores. The US, Poland, or Colombia have relatively low trust in government, and they also have low CCPI scores. Canada has a low CCPI score despite having a medium-high TGI score. [^8] Note that the theoretical analysis is valid only for democracies. However, because all countries in the sample are classified either as Democracy or Full Democracy in the Polity IV data series, and because the sample size is very small, I do not control for polity scores in the regressions. However, Canada has significant confirmed Oil reserves.
To shed further light on the empirical relation between trust in government and climate change policies, I regressed CCPI on TGI and a subset of the other control variables. In particular, I estimate the following regression equation:
$CCPI_{i}=\alpha+\beta\,TGI_{i}+\gamma\,\mathbf{X}_{i}+\epsilon_{i}$
The estimate of $\beta$ captures the relationship we are interested in. Table 2 below shows the regression results. Despite the low power, $\hat{\beta}$ is positive and statistically significant in all specifications. Moreover, a one standard deviation increase of TGI leads to an increase of CCPI between 0.463 and 0.495 standard deviations, depending on the chosen specification. Therefore, the relation appears to be of relevance. Checking for potential heteroscedasticity, the $p$-values of the Breusch-Pagan test are larger than 0.4 for all specifications, suggesting that heteroscedasticity is no problem. Inspecting the Q-Q plots of the different specifications (see Figure 5 in Appendix B), further supports this conclusion. Hence, there indeed seems to be a significant and positive correlation between CCPI and TGI, as suggested by the model.
Dependent variable:
CCPI
(1) (2) (3) (4)
TGI $0.484^{**}$ $0.481^{**}$ $0.498^{*}$ $0.515^{**}$
$(0.223)$ $(0.193)$ $(0.250)$ $(0.216)$
Oil $-0.168^{**}$ $-0.170^{**}$
$(0.065)$ $(0.067)$
Risk2022 $0.397$ $0.982$
$(2.832)$ $(2.458)$
Constant $28.825^{**}$ $30.629^{***}$ $27.249$ $26.754^{*}$
$(11.803)$ $(10.251)$ $(16.566)$ $(14.316)$
Observations 19 19 19 19
R2 0.217 0.446 0.218 0.452
Adjusted R2 0.170 0.377 0.120 0.343
F Statistic 4.699∗∗ (df = 1; 17) 6.451∗∗∗ (df = 2; 16) 2.224 (df = 2; 16) 4.128∗∗ (df = 3; 15)
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
Table 2: Regression Results.
6 Conclusion
This paper illustrates the negative impact that voter beliefs and expectations can have on political decision-making, particularly in the context of climate change policy. The model demonstrates that when voters are inclined to dismiss the severity of climate change, this skews the equilibrium toward suboptimal policy choices by incentivizing office-motivated politicians to act in ways that contradict their own information. The analysis reveals that—in the presence of motivated reasoning—trust in the functioning of government plays a critical role for the efficiency of policy-making. A first look at the data confirms this result. Moreover, the paper suggests that political rhetoric plays an important role as well: if politicians, scientists, journalists, and activists aim to contribute to better climate change policy, they must navigate a delicate balance between conveying information truthfully and avoiding the risk of pushing voters into denial.
Appendix A Mathematical Appendix
A.1 Proof of Lemma 1
The proof follows from (4):
$\begin{array}[]{rcl}\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}&>&0\\ \Leftrightarrow-\dfrac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}&>&0\\ \Leftrightarrow s\left[\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)\right]&<&0\end{array}$
Here we have $s$ and the expression in brackets that jointly determine the derivative’s sign. Consider next the term in parentheses:
$\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)>0\Leftrightarrow\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}\equiv\frac{\Delta+\beta}{1+\beta}$
If $\Delta>1$, then $\tilde{\kappa}>1$, and hence we cannot have $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$. It follows that in this case $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$. Hence, $\tilde{\mu}^{*}=\infty$ if $s>0$ and $\tilde{\mu}^{*}=0$ if $s<0$. If $s=0$, any $\tilde{\mu}$ yields the same $AU$ and hence $\tilde{\mu}^{*}=\sigma$.
If $\Delta<1$, $\tilde{\kappa}<1$ as well. Hence, if $\kappa(s,\mathbf{p},\tilde{\mu})<\tilde{\kappa}$, then again $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$. It follows that $\tilde{\mu}^{*}=\infty$ if $s>0$, $\tilde{\mu}^{*}=0$ if $s<0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$. To the contrary, if $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$. Hence, $\tilde{\mu}^{*}=0$ if $s>0$, $\tilde{\mu}^{*}=\infty$ if $s<0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$. If $\kappa(s,\mathbf{p},\tilde{\mu})=\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=0$, and any $\tilde{\mu}$ yields the same anticipatory utility. Thus, $\tilde{\mu}^{*}=\sigma$.
Finally, if $\Delta=1$, then we cannot have $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$. Thus, $AU(s,\mathbf{p},\tilde{\mu})$ is either flat in $\tilde{\mu}$ if $\kappa(s,\mathbf{p},\tilde{\mu})=1$, implying any $\tilde{\mu}$ yields the same anticipatory utility, and we have $\tilde{\mu}^{*}=\sigma$, or the situation resembles the case with $\Delta>1$ discussed above.∎
A.2 Proof of Corollary 1
Follows from the discussion in the text and Lemma 1.∎
A.3 Proof of Proposition 1
We prove the different cases successively.
$\Delta>1$:
We know from Corollary 1 that all voters with $s>0$ have belief $\pi(s)=\hat{\pi}$, while those with $s<0$ have belief $\pi(s)=0$. Hence, the latter always vote for policy 0. Wether a voter with $s>0$ votes for policy 1 or policy 0 depends on the comparison of $\tilde{\pi}$ and $\hat{\pi}$. In particular, we know from (5) that a voter votes for policy 1 if $\pi(s)>\tilde{\pi}$, for policy 0 if $\pi(s)<\tilde{\pi}$, and they choose each with a probability of 50% if $\pi(s)=\tilde{\pi}$. Hence, if $\hat{\pi}<\tilde{\pi}$, all voters vote for policy 0. If $\hat{\pi}=\tilde{\pi}$, all voters receiving $s<0$ plus half of the voter receiving $s>0$ vote for policy 0, while the other half of $s>0$ voters vote for policy 1. As a consequence, policy 0 wins all the time if $\hat{\pi}\leq\tilde{\pi}$. This implies that we must have $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ for all $s$.
However, if $\hat{\pi}>\tilde{\pi}$, all voters vote informatively, and thus the optimal policy always wins. It follows that in state 0 a majority votes for policy 0, while in state 1 a majority votes for policy 1. Hence, voters know that the policy that matches the state wins all the time. The belief about the state is $\pi(s)=\hat{\pi}$ if $s\geq 0$, implying $\left.\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s\geq 0}=\hat{\pi}$, and $\pi(s)=0$ if $s<0$, and hence $\left.\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s<0}=0$.
Could there be other pure strategy equilibria? $\tilde{\mu}^{*}$ is uniquely determined, and hence also $\pi(s,\mathbf{p},\tilde{\mu}^{*})$ is uniquely determined. But this implies that equilibrium vote shares are unique, leaving no room for beliefs $\kappa(s,\mathbf{p},\tilde{\mu})$ that are consistent with these vote shares and that differ from the ones established before. Hence, if $\Delta>1$, no other equilibrium in the voting subgame can exist.
$\Delta\leq 1$:
First assume that $\kappa(s,\mathbf{p},\tilde{\mu})$ is increasing in $s$ in the sense that $\text{Sign}\left[\kappa(s,\mathbf{p},\tilde{\mu})-\tilde{\kappa}\right]=\text{Sign}\left[s\right]$. Then any voter with $s<0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0. Moreover, any voter with $s>0$ chooses also $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1. This implies that indeed the policy that matches the state always wins and we need to have $\left.\kappa(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s<0}=0$ as well as $\left.\kappa(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s>0}=1$, proving part 2 (a) of the proposition.
Next, assume voters generally believe that policy 0 will be implemented, independent of the true state $\omega$. Then $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=0$ for all $s$. Then any voter with $s<0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0. Moreover, any voter with $s>0$ chooses $\tilde{\mu}^{*}=0$. If and only if $\hat{\pi}\leq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 0, and hence when $\hat{\pi}\leq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 0 is always chosen. This proves part 2 (b).
Now assume voters generally believe that policy 1 will be implemented, independent of the true state $\omega$. Then $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=1$ for all $s$, and any voter with $s>0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1. Moreover, any voter with $s<0$ chooses $\tilde{\mu}^{*}=0$. If and only if $\hat{\pi}\geq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 1, and hence when $\hat{\pi}\geq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 1 is always chosen. This proves part 2 (c).
Consider a situation in which beliefs are just as in the equilibrium discussed above for the case of $\Delta>1$. That is, $\kappa(s,\mathbf{p},\tilde{\mu})=0$ for all $s$. Then, as before, $\tilde{\mu}^{*}(s)=\infty$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ if $s<0$, and $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=q$ else. It follows that if $q\leq\frac{1}{1+\beta}$, a majority of voters always votes for policy 0. Hence, this is indeed an equilibrium.
Consider next a situation in which beliefs are just opposite to what we have seen above, namely $\kappa(s,\mathbf{p},\tilde{\mu})=1$ for all $s$. Then, $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ if $s>0$, and $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=q$ else. It follows that if $q\geq\frac{1}{1+\beta}$, a majority of voters always votes for policy 1. Hence, this is also an equilibrium.
Finally, consider part 2 (d). An equilibrium in which the optimal policy never wins would imply policy 1 wins if $\omega=0$ and policy 0 wins if $\omega=1$. Hence, we would need to have $\kappa(s,\mathbf{p},\tilde{\mu})$ weakly decreasing in $s$. Consider a voter with signal $s>0$. It must be true that most of these voters vote for policy 0. Hence, it cannot be true that $\kappa\geq\tilde{\kappa}$. If $\kappa(s,\mathbf{p},\tilde{\mu})<\tilde{\kappa}$, these voters all hold belief $\hat{\pi}$. Next consider voters with $s<0$. It must be true that most of these voters vote for policy 1, implying $\kappa(s,\mathbf{p},\tilde{\mu})\leq\tilde{\kappa}$ is not possible. Thus, because $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$, these voters all also hold belief $\hat{\pi}$. If $\hat{\pi}>\tilde{\pi}$, policy 1 always wins, contradicting that the optimal policy never wins. If $\hat{\pi}<\tilde{\pi}$, policy 0 always wins, also contradicting that the optimal policy never wins. Finally, if $\hat{\pi}=\tilde{\pi}$, in each state, each policy wins with 50%, and thus also the optimal policy wins with a chance of 50%. This implies that the actual state is irrelevant for the probability of each policy winning, and thus we must have $\kappa(s,\mathbf{p},\tilde{\mu})=\frac{1}{2}$ for all $s$. But then it is not possible that $\kappa(s,\mathbf{p},\tilde{\mu})$ differs for voters with $s>0$ and voters with $s<0$. Hence, such an equilibrium of the voting subgame cannot exist.
This complete the proof of the proposition.∎
A.4 Proof of Proposition 2
Follows from the discussion in the text.∎
A.5 Proof of Proposition 3
We first consider the case of severe climate change being catastrophic, $\Delta>1$. The other case of severe climate change having only mild baseline welfare consequences will be considered thereafter. Note that $\hat{\pi}=q\leq\tilde{\pi}$ in a symmetric pure strategy equilibrium. Suppose that off the equilibrium path voters observe $\tilde{\mathbf{p}}\in\{\{0,1\},\{1,0\}\}$.
$\Delta>1$:
We know from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=\hat{\pi}$ if $s\geq 0$. Hence, $\hat{\pi}\leq\tilde{\pi}$, and thus no voter ever votes for policy 1, whereas policy 0 always wins at the ballot. It follows that $\kappa^{*}(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$, with $\tilde{\mu}^{*}(s)$ as described in Lemma 1. Thus, the unique equilibrium is candidates choosing $\mathbf{p}^{*}=(0,0)$. This is supported by $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$ and $\tilde{\mu}^{*}(s)$ as described in Lemma 1.
$\Delta\leq 1$:
Assume first $\kappa(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ for all $s$. Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=\hat{\pi}=q$ if $s>0$. Because $q\leq\frac{1}{1+\beta}$, a majority of voters always votes for policy 0. It follows from Proposition 2 that no candidate has an incentive to campaign on $p=1$. Consequently, $\mathbf{p}^{*}=(0,0)$, $\kappa(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, and $\tilde{\mu}^{*}$ follows from Lemma 1.
Suppose next that $\Delta<1$ and that $\kappa>\tilde{\kappa}$ if $s>0$ and $\kappa<\tilde{\kappa}$ if $s<0$. Note that this means that we must have $\Delta<1$. Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=1$ if $s>0$. Hence, all voters all the time vote informatively, and thus the policy that matches the state wins, implying $\kappa=\pi(s,\tilde{\mathbf{p}},\tilde{\mu})$. It follows from Proposition 2 that, if candidates expect voters to behave this way, they have an incentive to choose the socially optimal policies. Hence, $\mathbf{p}^{*}=(\omega,\omega)$, $\kappa(s,\mathbf{p}^{*},\tilde{\mu}^{*})=\pi(s,\mathbf{p}^{*},\tilde{\mu}^{*})$, and $\tilde{\mu}^{*}$ follows from Lemma 1.
Now suppose that $\Delta=1$. Moreover, suppose an efficient equilibrium exists also in this case, with $\kappa(s,\mathbf{p},\tilde{\mu})=1=\tilde{\kappa}$ if $s>0$, and $\kappa=0<\tilde{\kappa}$ if $s<0$. Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$, and thus any voter with $s<0$ votes for policy 0. For voters with $s>0$, $\tilde{\mu}^{*}=\mu$, and thus these voters hold perfect Bayesian beliefs. This implies, by Assumption 2, that voters vote informatively. However, then the right policy is chosen in any state, and thus we need to have $\kappa(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\tilde{\mu})$. But $\pi(s,\mathbf{p},\tilde{\mu})<1$, and therefore voters with $s>0$ would choose $\tilde{\mu}=0$ instead of $\tilde{\mu}=\mu$. Hence, when $\Delta=1$, no efficient equilibrium exists.
It follows from Proposition 1 part 2 (d) that no equilibrium can exist in which always the wrong policy is chosen. Hence, we now only need to show that no equilibrium, in which policy 1 always wins, can exist. Suppose to the contrary that voters believe that policy 1 wins with certainty, $\kappa(s,\mathbf{p},\tilde{\mu})=1\geq\tilde{\kappa}$ for all $s$. Then it follows from Corollary 1 that, if $\Delta<1$, $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=q\leq\tilde{\pi}$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=1$ if $s>0$. Therefore, all voters with $s>0$ always vote for policy 1. If sufficiently many voters with $s<0$ also vote for policy 1, policy 1 wins all the time. This is the case iff $q\geq\frac{1}{1+\beta}$. However, this contradicts Assumption 1. A similar argument shows that $\mathbf{p}=(1,1)$ cannot be equilibrium, either, if $\Delta=1$. Hence, it follows that no equilibrium, in which policy 1 is always chosen, exists. This proves the proposition. ∎
A.6 Proof of Lemma 2
Take the belief about $\omega$ of a voter who receives $s$ and chooses $\tilde{\alpha}$:
$\pi(s,\tilde{\alpha})=\frac{q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)}{q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)+(1-q)\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right)}.$
Anticipatory utility is defined as before in (2). Hence, it’s derivative with respect to $\tilde{\alpha}$ is
$\dfrac{\partial AU}{\partial\tilde{\alpha}}=\frac{(1-q)qg(s)\left[\beta+\Delta-(\beta+1)\kappa\right](f_{1}(s)-f_{0}(s))}{\left[q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)+(1-q)\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right)\right]^{2}}$
The derivative is positive if $f_{1}(s)-f_{0}(s)>0\Leftrightarrow s>0$, negative if $f_{1}(s)-f_{0}(s)<0\Leftrightarrow s<0$, and zero else. This proves the result. ∎
A.7 Proof of Proposition 4
It follows from Lemma 2 that independent of the true state $\omega$, a majority of voters prefers to vote for policy 0 when offered the choice. The proof then follows from Proposition 2. ∎
A.8 Proof of Lemma 3
We focus on pure strategy equilibria. It then follows from Lemma 1 that when $s>0$, a voter would either like to choose $\tilde{\mu}=\infty$, or $\tilde{\mu}=\mu$. Moreover, when $s<0$, a voter would either like to choose $\tilde{\mu}=0$, or $\tilde{\mu}=\mu$. Next have a look how a voter’s objective function changes with $s$:
$\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial s}=-\frac{2\mu(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\mu s}{\tilde{\mu}^{2}}}}{\left(q\tilde{\mu}\left(e^{\frac{2\mu s}{\tilde{\mu}^{2}}}-1\right)+\tilde{\mu}\right)^{2}}<0,$ (7)
and hence $W(s,\mathbf{p},\tilde{\mu})$ strictly decreases in $s$ for all $\tilde{\mu}$. Note that when $\tilde{\mu}=0$, then the realization of $s$ is irrelevant as the belief is $\left.\pi(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}=q$. Hence, the voter with the greatest incentive to distort $\mu$ has $s\rightarrow\infty$. For this voter, the payoff of not distorting and instead choosing $\tilde{\mu}=\mu$ is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu\wedge s\rightarrow\infty}=-\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))-\Delta.$
If, to the contrary, any such voter chooses $\tilde{\mu}=0$, the payoff is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}=-\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta)-\gamma.$
If $\gamma\geq\gamma^{+}\equiv(1-q)(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta-\kappa(s,\mathbf{p},\tilde{\mu}))$, then no voter with $s>0$ has an incentive to distort the issue’s complexity. Otherwise, that is, if $\gamma<\gamma^{+}$, then there exists a unique signal realization $s=s^{+}>0$ such that a voter receiving this signal is indifferent between choosing $\tilde{\mu}=0$ and $\tilde{\mu}=\mu$. The payoff of a voter choosing to not distort a signal $s$ is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=-\frac{\kappa(s,\mathbf{p},\tilde{\mu})(1-q)+q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta)e^{\frac{2\mu s}{\sigma^{2}}}}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}.$
To find the indifferent voter, we need to solve
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}\Leftrightarrow s=s^{+}\equiv\frac{\sigma^{2}\ln\left(1+\frac{\gamma}{q(\beta-\gamma+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta+\Delta))}\right)}{2\mu}.$
Next consider a voter with $s<0$. It follows from Lemma 1 that such a voter would either like to choose $\tilde{\mu}=\infty$ or $\tilde{\mu}^{*}=\mu$. Moreover, by (7), when $\tilde{\mu}=\mu$, utility is the lowest when $s=0$. Hence, if a voter receiving $s=0$ has no incentive to distort $\mu$, then no voter with $s<0$ has an incentive to do so. The payoff of not distorting is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu\wedge s=0}=-\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta).$
If, however, a voter with $s<0$ chooses to distort beliefs, he receives
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\infty\wedge s<0}=-\kappa(s,\mathbf{p},\tilde{\mu})-\gamma.$
Hence, if $\gamma\geq\gamma^{-}\equiv q(\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu}))$, then no voter with $s<0$ chooses $\tilde{\mu}\neq\mu$. However, if $\gamma<\gamma^{-}$, then there exists a voter with $s=s^{-}<0$ who is indifferent between distorting and not distorting. To find the indifferent voter, we need to solve
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\infty\wedge s<0}\Leftrightarrow s=s^{-}\equiv\frac{\sigma^{2}\ln\left(\frac{\gamma(q-1)}{q(\beta(\kappa(s,\mathbf{p},\tilde{\mu})-1)+\gamma-\Delta+\kappa(s,\mathbf{p},\tilde{\mu}))}\right)}{2\mu}.$
Any voter with $s<s^{-}$ chooses $\tilde{\mu}=\mu$, while any voter with $s\in(s^{-},0)$ chooses $\tilde{\mu}=0$.∎
A.9 Proof of Proposition 5
In an inactivity equilibrium, we have $\kappa(s,\mathbf{p},\tilde{\mu})=0$. Then, $\gamma_{0}^{+}\equiv\left.\gamma^{+}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=(1-q)(\beta+\Delta)$ and $\gamma_{0}^{-}\equiv\left.\gamma^{-}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=q(\beta+\Delta)$. Note that when $q>\frac{1}{2}$, then $\gamma_{0}^{-}>\gamma_{0}^{+}$. Hence, if severe climate change is more likely a priori, then there are more distortions of good news than of bad news, and vice versa. Moreover, $\left.s^{+}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=\frac{\sigma^{2}\ln\left(1+\frac{\gamma}{q((\Delta+\beta)(1-q)-\gamma)}\right)}{2\mu}$ and $\left.s^{-}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=\frac{\sigma^{2}\ln\left(\frac{\gamma(1-q)}{q(\Delta+\beta-\gamma)}\right)}{2\mu}.$ Hence, if $\gamma<\gamma_{0}^{+}$, then beliefs after $s>0$ are $q$ if $s\geq s^{+}_{0}$ and $\pi(s,\mathbf{p},\tilde{\mu})\in(q,1)$ else. Moreover, if $s<0$ and $\gamma<\gamma_{0}^{-}$, then beliefs are 00 if $s\in[s^{-}_{0},0)$ and $\pi(s)\in(0,q)$ else.
Note that all voters with $s<0$ hold beliefs weakly below Bayesian beliefs. Because of motivated beliefs, they vote to a greater extent informatively than they should, which is positive for information aggregation. Because $q\leq\tilde{\pi}$, they all vote informatively. If $\omega=0$, therefore a majority votes for policy 0. Hence, what matters for candidate incentives is the share of $s>0$ voters voting informatively. Because $q<\tilde{\pi}$, those with a belief of $q$ vote for policy 0. That is, all voters with $s>s_{0}^{+}$. Those with $s\in(0,s^{+})$ may vote either way.
The belief of a voter who chooses $\tilde{\mu}=\mu$ and receives signal $s=s^{+}_{0}$ is $\pi(s_{0}^{+},\mathbf{p},0)=\frac{\gamma+\Delta q+q}{\Delta+1}$. This is smaller than $\tilde{\pi}$ if
$\frac{\gamma+\Delta q+q}{\Delta+1}<\frac{1}{1+\beta}\Leftrightarrow\gamma<\tilde{\gamma}\equiv\frac{(\Delta+1)(1-q(1+\beta))}{\beta+1}>0.$
The belief of a voter who chooses $\tilde{\mu}=\mu$ and receives signal $s>0$ is
$\pi(s,\mathbf{p},\tilde{\mu})=\frac{q}{q+(1-q)e^{-\frac{2s\mu}{\sigma^{2}}}}.$
The signal at which such a voter starts to vote for policy 1 is the signal such that $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$. In particular, this signal realization is $s^{*}$, where $s^{*}\equiv\frac{\sigma^{2}\ln\left(\frac{1-q}{\beta q}\right)}{2\mu}$. If $\Phi\left(\frac{s-\mu}{\sigma}\right)$ is the CDF of the normal distribution with mean $\mu$ and standard deviation $\sigma$, the share of voters voting for policy 1 if $\omega=1$ is
$\mathcal{V}_{0}=\Phi\left(\frac{s^{+}_{0}-\mu}{\sigma}\right)-\Phi\left(\frac{s^{*}-\mu}{\sigma}\right).$
Recall that $s_{0}^{+}$ is increasing in $\gamma$, and hence $\mathcal{V}_{0}$ is monotonically increasing in $\gamma$. When $\gamma=0$, we know from our earlier analysis that voters all vote for policy 0 despite $s>0$. If $\gamma\rightarrow\infty$, no voters distorts $\mu$ and, by Assumption 2, a majority votes for policy 1 iff $\omega=1$. By continuity, there is a unique $\hat{\gamma}$ such that a majority votes for policy 0 if and only if $\gamma<\hat{\gamma}$. ∎
A.10 Proof of Lemma 4
To see incentives, first consider how $W(s,\mathbf{p},\tilde{\mu})$ changes with $\tilde{\mu}$, when $\tilde{\mu}\geq\mu$. It is easy to show that $W(s,\mathbf{p},\tilde{\mu})$ is strictly concave in $\tilde{\mu}$ and thus an interior equilibrium exists for all $s$. The FOC for an interior optimum is
$\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c(\tilde{\mu}-\mu).$
if $\tilde{\mu}>\mu$ and
$\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu).$
if $\tilde{\mu}<\mu$. In both derivatives, the first term’s sign is the opposite of the sign of $s$ (when $\Delta>1$). Hence, if $s>0$, then $\tilde{\mu}^{*}(s)<\mu$. Similarly, if $s<0$, we must have $\tilde{\mu}^{*}(s)>\mu$. If $s=0$, the first term is zero, and hence we need to have $\tilde{\mu}^{*}(s)=\mu$. Moreover, $\lim_{s\rightarrow\infty}\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu)$, and hence $\lim_{s\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$. Similarly, $\lim_{s\rightarrow-\infty}\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-c(\tilde{\mu}-\mu)$, and hence also $\lim_{s\rightarrow-\infty}\tilde{\mu}^{*}(s)=\mu$.
Now consider $s<0$, for which we have $\tilde{\mu}>\mu$. Rearranging the FOC yields
$-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}+\mu=\tilde{\mu}$
That is, we can express $\tilde{\mu}$ as a deviation from $\mu$, and
$\Gamma=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}$
defines the deviation. $\Gamma$ increases in $s$ iff
$\frac{\partial\Gamma}{\partial s}=\frac{-2(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left((1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)\right)}{c\sigma^{4}\left(q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+1\right)^{3}}>0,$
which is the case iff $\Omega=(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)<0$. If $s=0$, $\left.\Omega\right|_{s=0}=\sigma^{2}>0$, and hence $\tilde{\mu}^{*}(s)$ is decreasing. Moreover, $\lim_{s\rightarrow-\infty}\Omega=-\infty$. Thus, if $\Omega$ is monotone in $s$, then there is a unique $s^{-}<0$ such that $\tilde{\mu}^{*}(s)$ increases in $s$ if $s<s^{-}$ and it decreases in $s$ if $s\in(s^{-},0]$. We have
$\frac{\partial\Omega}{\partial s}=2\tilde{\mu}\left(1-q-q\frac{2\tilde{\mu}se^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}}\right)>0.$
This proves the existence of a unique $s^{-}$. Moreover, because $\left|\Gamma\right|$ decreases in $c$, the absolute deviation from $\mu$ decreases in the cost $c$.
Next consider $s>0$, implying $\tilde{\mu}<\mu$. Rearranging the FOC yields
$-\frac{2\tilde{\mu}^{3}(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}+\mu=\tilde{\mu}$
That is, we can again express $\tilde{\mu}$ again as a deviation from $\mu$, and
$\Gamma^{\prime}=-\frac{2\tilde{\mu}^{3}(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}$
defines the deviation. $\Gamma^{\prime}$ increases in $s$ iff
$\frac{\partial\Gamma^{\prime}}{\partial s}=\frac{2\tilde{\mu}^{3}(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)\right)}{c\sigma^{4}\left(\mu+\mu q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)\right)^{3}}>0$
This is the case iff $\Omega^{\prime}=qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)>0$. If $s=0$, $\left.\Omega\right|_{s=0}=-\sigma^{2}<0$, and hence $\tilde{\mu}^{*}(s)$ is decreasing. Moreover, $\lim_{s\rightarrow\infty}\Omega^{\prime}=\infty$. Finally,
$\frac{\partial\Omega^{\prime}}{\partial s}=\frac{2\tilde{\mu}qse^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}}+q-1$
This is negative when $s=0$. Hence, if we can show that $\Omega^{\prime}$ is convex, we prove the existence of a unique $s^{+}$. Take the second derivative with respect to $s$:
$\frac{\partial^{2}\Omega^{\prime}}{\partial s^{2}}=\frac{2\tilde{\mu}qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(\sigma^{2}+2\tilde{\mu}s\right)}{\sigma^{4}}>0,$
and thus $\Omega^{\prime}$ is indeed convex in $s$. Thus, there exists a unique $s^{+}>0$ such that $\tilde{\mu}^{*}(s)$ increases in $s$ if $s>s^{+}$ and it decreases in $s$ when $s\in[0,s^{+})$. Moreover, because $\left|\Gamma^{\prime}\right|$ decreases in $c$, the absolute deviation from $\mu$ decreases in the cost $c$. This proves the lemma. ∎
A.11 Proof of Proposition 6
The proof basically follows along the lines of the discussion in the text. We know from Lemma 4 that $\left|\tilde{\mu}(s)-\mu\right|$ decreases in $c$. Moreover, $\lim_{c\rightarrow\infty}\tilde{\mu}(s)=\mu$, because then both $\Gamma=0$ and $\Gamma^{\prime}=0$ (see the proof of Lemma 4). Moreover, as $c\searrow 0$, we are back in the scenario of Proposition 3. Finally note that both $\tilde{\mu}(s)$ and $\pi(s,\mathbf{p},\tilde{\mu})$ are continuous functions of both $s$ and $c$, and thus so is the share of voters voting for policy 1 in state $\omega=1$ if $p_{1}\neq p_{2}$. Hence, there must exist $\hat{c}$ such that the share of voters $\mathcal{V}_{0}$ voting for policy 1 in state 1 is just $\frac{1}{2}$. Moreover, $\text{Sign}\left[c-\hat{c}\right]=\text{Sign}\left[\mathcal{V}_{0}-\frac{1}{2}\right]$. This proves the proposition. ∎
Appendix B Additional Tables and Figures
Figure 5: Q-Q plots for the different regressions reported in Table 2.

### v5

@@
 Abstract
-Two office-driven politicians compete in an election by proposing policies.
-There are two possible states of the world: climate change is either mild, with no lasting effect on welfare if addressed properly, or severe, leading to reduced welfare even with appropriate measures.
-Voters receive signals about the state but may interpret them in a non-Bayesian way, holding motivated beliefs.
-An equilibrium always exists where voters ignore signals suggesting severe consequences, causing politicians to propose policies for mild climate change—even when they know otherwise.
-If severe climate change leads to only moderate welfare losses, an efficient equilibrium also exists.
-In this equilibrium, voters trust politicians to choose the optimal policies, implying voters choose to trust their signals, which in turn encourages optimal policy choices by politicians.
-The model highlights the role of political rhetoric and trust in government, and a first glance at the data reveals patterns consistent with the models predictions.
+We study how motivated reasoning affects the provision of climate policy in an electoral competition framework.
+Voters experience anticipatory disutility when future outcomes appear grim and may therefore distort beliefs in response to adverse information.
+We develop a game-theoretic model in which voters and politicians receive signals about the severity of climate change.
+When the anticipated welfare losses from severe climate change are sufficiently large, voters optimally ignore unfavorable information, inducing politicians to campaign on policies appropriate for mild climate change only.
+When welfare losses are moderate, the model admits a second, efficient equilibrium in which voters trust politicians to implement welfare-maximizing policies and vote informatively, thereby creating incentives for politicians to propose adequate climate policy.
+The model shows how motivated belief formation and voters’ expectations about policy responsiveness jointly determine equilibrium selection between effective climate policy and persistent political inaction.
 JEL Codes: D72, D91, H12
-Keywords: political competition, climate change, motivated reasoning, trust in government, political rhetoric, misinformation
+Keywords: political competition, climate change, motivated reasoning, trust in government, political rhetoric
 1 Introduction
@@
 a reduction in concern and support for action.”
-Clayton, Manning, Krygsman, and Speiser (2017)
-“Yes, There Has Been Progress on Climate.
-No, It’s Not Nearly Enough.”
-The New York Times, 25 Oct 2021
+Clayton et al., (2017)
 Climate change is one of the most pressing issues of our time, threatening the livelihoods of millions of people around the globe.
 Information about climate change and global warming has been available for more than a century, dating back at least to Svante Arrhenius’ (1896) famous paper “On the Influence of Carbonic Acid in the Air upon the Temperature of the Ground.” However, despite an ever-growing scientific consensus that increasing atmospheric CO2 concentrations will severely impact our planet, little to no action was taken to stop this process for most of the last few decades.
-Many reasons have been put forward to explain the inaction of political decision-makers, ranging from lobbying interests to widespread misinformation campaigns like those of ExxonMobil.
-In this paper, I discuss a different, simpler channel: the electoral incentives of politicians when voters may choose to hold motivated beliefs.
+Many reasons have been put forward to explain the inaction of political decision-makers, ranging from lobbying interests to widespread misinformation campaigns like those of ExxonMobil (e.g., Oreskes and Conway,, 2011, or Supran et al.,, 2023).
+In this paper, we discuss a different, simpler channel: the electoral incentives of politicians when voters may hold motivated beliefs.
 The looming dire consequences of climate change may, if taken seriously, create stress and anxiety.
 In an effort to avoid these negative emotions, people may choose to hold motivated beliefs, ignoring information that suggests climate change is severe, while overreacting to information suggesting there is nothing to worry about.
-Psychologists and economists alike have long been aware of such information processing biases, as exemplified by Kunda’s (1990) seminal paper or the recent survey by Amelio and Zimmermann (2023), and multiple studies have provided evidence for their empirical relevance (for example, Lewandowsky, Gignac, and Oberauer, 2013, Thaler, 2021, 2024, and Lois, Tsakas, Yuen, and Riedl, 2024).
-To study the potential consequences of such behavior for the implementation of policies targeted at climate change, I construct a game-theoretic model in which two politicians compete in an election by proposing policies.
-There are two possible states of the world: severe climate change and moderate (or no) climate change.
-The state determines which policy is optimal and establishes a baseline welfare level.
-If climate change is severe, the baseline welfare is lower than in the mild state, regardless of the chosen policy.
-Voters and politicians receive signals about the state of the world, and society is best served when the policy that matches the state is implemented.
-As discussed by authors like Bénabou and Tirole (2016) or Spiegler (2016), voters have anticipatory utility, meaning they feel anxiety or stress when the future looks grim.
-To cope with these negative emotions, they may choose to hold motivated beliefs, interpreting information about the future in a non-Bayesian way to increase anticipatory utility.
-How voters react to information is an important determinant of how politicians compete to win an election.
-The model highlights the importance of the portrayed severity of climate change for the efficiency of policy-making.
-In particular, if severe climate change has only moderate consequences for baseline welfare levels, then there always exists an efficient equilibrium in which politicians choose the optimal policies for each state.
-However, this equilibrium only exists if voters have sufficient trust in government, i.e., if they belief that elected politicians will enact the right policies with a sufficiently high probability.
-Otherwise, with low trust in government, there always exists an equilibrium in which politicians ignore their information and choose policies for mild or no climate change.
-The analysis suggests two important determinants for the efficiency of policy-making.
-First, trust in government is important, because it determines in which way voters interpret information when severe climate change only has moderate effects on baseline welfare.
-With sufficient trust in government, there always exists an efficient equilibrium in which politicians strictly choose the optimal policies.
-To the contrary, if trust in government is low, then politicians will ignore their information and policy choice will be inefficient.
-Second, the analysis suggests that political rhetoric could be an important determinant of the quality of policy choice as well.
-In particular, if the consequences of severe climate change are portrayed as catastrophical, then the efficient equilibrium ceases to exist.
-Voters will ignore any information that climate change is severe, inducing office motivated politicians to ignore their information as well, and in the unique equilibrium politicians always choose policies for mild or no climate change.
-However, if the consequences of severe climate change are portrayed as moderate or mild, then the efficient equilibrium survives.
-Literature:
+Psychologists and economists alike have long been aware of such information processing biases, as exemplified by Kunda’s (1990) seminal paper or the recent survey by Amelio and Zimmermann, (2023), and multiple studies have provided evidence for their empirical relevance (for example, Lewandowsky et al.,, 2013, Thaler,, 2021, 2024, and Lois et al.,, 2024).
+Motivated reasoning alone, however, is unlikely to fully account for persistent political inaction on climate change.
+Recent empirical work has also emphasized the role of political trust in shaping climate policy.
+In particular, higher trust in political institutions has been associated with local politicians adopting more ambitious climate policies (Bose,, 2023).
+Figure 1 shows that a similar positive relationship between political trust and climate policy stringency emerges in a cross-country comparison.
+[^1] Data on climate policy performance come from the Climate Change Performance Index (CCPI), published jointly by Germanwatch, the NewClimate Institute, and the Climate Action Network International.
+The CCPI ranges from 0 to 100, with higher values indicating better performance in climate policy.
+It aggregates indicators on greenhouse gas emissions (40%), renewable energy (20%), energy use (20%), and other climate policies (20%).
+Data on trust are taken from the OECD’s Trust in Government Indicator (TGI), which measures the share of a country’s population reporting confidence in their national government and also ranges from 0 to 100.
+These patterns point to an interaction between belief distortions and political trust in determining policy outcomes, a mechanism that the analysis below formalizes.
+Figure 1: Scatterplot with fitted trendline and 95% confidence interval showing countries’ CCPI scores on the vertical axis and TGI scores on the horizontal axis.
+The aim of this paper is to construct a game-theoretic model that sheds light on how motivated reasoning affects politicians’ incentives to propose adequate climate change policies within an electoral competition framework.
+In the model, there are two possible states of the world: severe climate change and mild climate change.
+The state determines which policy is optimal and establishes a baseline level of welfare.
+If climate change is severe, baseline welfare is lower than in the mild state, regardless of the policy chosen.
+Voters and politicians receive signals about the state of the world, and society is best served when the implemented policy matches the true state.
+As in Brunnermeier and Parker, (2005), Bénabou and Tirole, (2016), or Spiegler, (2016), voters derive anticipatory utility and experience anxiety or stress when the future appears grim.
+To cope with these negative emotions, they may adopt motivated beliefs, interpreting information about the future in a non-Bayesian manner to thus increase anticipatory utility.
+How voters respond to information is therefore a key determinant of how politicians compete in elections.
+The analysis reveals that when the consequences of severe climate change are sufficiently large (“catastrophic”), voters maximize anticipatory utility by disregarding signals indicating that the problem is severe.
+This, in turn, induces politicians to campaign exclusively on policies designed for mild climate change.
+In contrast, when the welfare losses from severe climate change are moderate rather than catastrophic, multiple equilibria arise.
+In an efficient equilibrium, voters trust politicians to choose welfare-maximizing policies, thereby creating incentives for politicians to do so.
+In an inefficient equilibrium, voters expect politicians to ignore relevant information and to campaign on policies appropriate for mild climate change, which in turn gives politicians incentives to disregard severe climate risks.
+Thus, when the consequences of severe climate change are moderate, voter trust in welfare-maximizing policy choices becomes self-fulfilling, sustaining politicians’ incentives to propose such policies.
+Trust in government functions as a coordination device.
+When voters believe that politicians will implement policies that match the underlying state, signals are interpreted more faithfully, allowing society to converge to the efficient equilibrium.
+By contrast, when trust is low—interpreted as the belief that policy choices are largely independent of the underlying state—this link breaks down.
+In that case, pessimistic voters interpret information in ways that rationalize the policies they expect the government to pursue, giving rise to a self-confirming equilibrium in which inefficient policies persist.
+The model therefore identifies trust as an important determinant of whether political institutions can translate information about climate risks into effective action.
+The analysis shows that an efficient equilibrium is less likely to exist when the anticipated welfare losses from severe climate change are sufficiently large.
+How the severity of climate change is framed—whether it is presented as a “catastrophe” or as a more moderate “challenge”—may influence voters’ expectations about the magnitude of potential losses from climate change.
+An indirect implication of the model is, therefore, that climate change rhetoric can influence voters’ anticipatory utility and, through this channel, their belief formation, which in turn determines equilibrium outcomes.
+In this sense, highly alarming descriptions need not translate into greater support for policy action and may instead induce defensive belief formation, a pattern emphasized by Clayton et al., (2017) in the quote at the beginning of the paper.
+Related Literature:
 The paper contributes to several areas of literature.
 Firstly, it adds to the literature that examines the incentives of political candidates to select efficient policy platforms when the true state of the world is unknown.
-Important contributions to this literature include Heidhues and Lagerlof (2003), Laslier and van der Straeten (2004), and Kartik, Squintani, and Tinn (2015).
-Unlike Heidhues and Lagerlof (2003) and Kartik, Squintani, and Tinn (2015), I assume that voters also receive information, which creates the main tension of whether to ignore it or not.
-In Laslier and van der Straeten (2004), voters receive a single public signal as well, whereas in the current paper, each voter receives an independent signal.
-The two most closely related papers in this literature are Gratton (2014) and Crutzen, Sisak, and Swank (2024).
-Gratton (2014) assumes that candidates observe the state perfectly while voters receive imperfect but informative signals about the state, and the focus of the analysis is to identify conditions that lead equilibria such that candidates choose efficient policies.
-Crutzen, Sisak, and Swank (2024) study a game like Gratton (2014), but there are two groups of voters, elites and commoners, who stochastically differ in their policy preferences and who always differ in quality of the information they receive.
+Important contributions to this literature include Heidhues and Lagerlof, (2003), Laslier and van der Straeten, (2004), and Kartik et al., (2024).
+Unlike Heidhues and Lagerlof, (2003) and Kartik et al., (2024), we assume that voters also receive information, which creates the main tension of whether to ignore it or not.
+In Laslier and van der Straeten, (2004), voters receive a single public signal as well, whereas in the current paper, each voter receives an independent signal.
+The two most closely related papers in this literature are Gratton, (2014) and Crutzen et al., (2024).
+Gratton, (2014) assumes that candidates observe the state perfectly while voters receive imperfect but informative signals about the state, and the focus of the analysis is to identify conditions that lead to equilibria such that candidates choose efficient policies.
+Crutzen et al., (2024) study a game like Gratton, (2014), but there are two groups of voters, elites and commoners, who stochastically differ in their policy preferences and who always differ in quality of the information they receive.
 The analysis focusses on deriving conditions such that candidates cater to the elites or engage in populism.
-Unlike in those papers, the focus in the current paper is on misinterpreting information, and how this influences candidates’ incentives to provide efficient policies.
-A cornerstone of the model is that voters have anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare” (Bénabou and Tirole, 2016).
-As in Akerlof and Dickens (1982), Caplin and Leahy (2001), Bénabou and Tirole (2002), Spiegler (2016), Eliaz and Spiegler (2020), and other papers, this means that voters have preferences not only about states and policies but also about their own beliefs.
-A consequence of this is that voters may engage in motivated reasoning; see, for example, Kunda (1990), Bénabou and Tirole (2016), Zimmermann (2020), Oprea and Yuksel (2021), or Thaler (2024).
-The most closely related papers in this area are Levy (2014) and Le Yaouanq (2023), which also examine electoral contexts.
-Levy (2014) presents a model featuring a policymaker trying to signal quality/congruency to voters.
+Unlike in those papers, the focus of the current paper is on voters’ misinterpretation of information and how this influences candidates’ incentives to propose efficient policies.
+This is related to Millner et al., (2020), who study how confirmation bias affects electoral outcomes.
+In contrast, we consider a model in which the future affects voters’ well-being through anticipatory utility, giving them an incentive to hold motivated beliefs.
+As just foreshadowed, a cornerstone of the model is that voters derive anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare” (Bénabou and Tirole,, 2016).
+As in Akerlof and Dickens, (1982), Caplin and Leahy, (2001), Bénabou and Tirole, (2002), Brunnermeier and Parker, (2005), Spiegler, (2016), Eliaz and Spiegler, (2020), Little et al., (2022), and Engelmann et al., (2024), this means that voters have preferences not only about states and policies but also about their own beliefs.
+A consequence of this is that voters may engage in motivated reasoning; see, for example, Kunda, (1990), Bénabou and Tirole, (2016), Zimmermann, (2020), Oprea and Yuksel, (2021), or Thaler, (2024).
+The most closely related papers in this area are Levy, (2014) and Le Yaouanq, (2023), which also examine electoral contexts.
+Levy, (2014) presents a model featuring a policymaker trying to signal quality/congruency to voters.
 Voters have imperfect memory and can suppress certain news by purposefully conflating positive and negative signals.
@@
 This difference introduces a novel self-fulfilling aspect that becomes critical for equilibrium.
-Le Yaouanq (2023) suggests that voters may interpret signals in a contrary manner at a cost.
-In the current paper, I assume that voters can modulate the perceived informativeness of a signal, but they cannot alter its direction.
-Moreover, unlike in Le Yaouanq (2023), a key component of my model is that the state influences baseline welfare levels, which has important consequences for equilibrium.
-The paper also contributes to a recently revived literature on environmental policymaking.
-Delfgaauw and Swank (2024) show that there can be locked-in effects that prevent the adoption of environmentally friendly policies.
-Blumenthal (2024) shows that efficient environmental policies may not be adopted if voters’ preferences are expected to change over time.
-Van der Straten, Perotti, and Van der Ploeg (2024) study the relationship between climate change adoption policies and inequality in society.
-Besley and Persson (2023) examine the conditions for green transitions to occur, depending on voters’ ideological preferences and the extent to which the production sector remains reliant on brown energy sources.
-Gullberg (2008), Shapiro (2016), and Balles, Matter, and Stutzer (2024) show that special interest groups may prevent climate change policies from being adopted.
-Denter (2024) shows that the wrong kind of climate activism may decrease support for climate policies.
+Le Yaouanq, (2023) suggests that voters may interpret signals in a contrary manner at a cost.
+In the current paper, we assume that voters can modulate the perceived informativeness of a signal, but they cannot alter its direction.
+Moreover, unlike in Le Yaouanq, (2023), a key component of our model is that the state influences baseline welfare levels, which has important consequences for equilibrium.
+The paper also contributes to the literature on environmental policymaking.
+Delfgaauw and Swank, (2024) show that there can be locked-in effects that prevent the adoption of environmentally friendly policies.
+Blumenthal, (2024) shows that efficient environmental policies may not be adopted if voters’ preferences are expected to change over time.
+Van der Straten et al., (2024) study the relationship between climate change adoption policies and inequality in society.
+Besley and Persson, (2023) examine the conditions for green transitions to occur, depending on voters’ ideological preferences and the extent to which the production sector remains reliant on brown energy sources.
+Gullberg, (2008), Shapiro, (2016), and Balles et al., (2024) show that special interest groups may prevent climate change policies from being adopted.
 In contrast to these papers, the current paper provides an explanation based purely on the electoral incentives of office-motivated politicians in the presence of a common information processing bias among voters.
+Finally, the paper contributes to the literature studying trust as a determinant of policy making and welfare.
+Dasgupta, (2000) argues that higher levels of trust in society reduce transaction costs and thereby increase welfare.
+Relatedly, Zak and Knack, (2001) show that higher trust stimulates investment and economic growth.
+Aghion et al., (2010) demonstrate that lower trust in society causes voters to demand more regulation, and that excessive regulation, in turn, hinders the formation of trust.
+This mechanism is similar to the current model, where low trust in government leads voters to demand policies not targeted at climate change.
+However, their model predicts a negative relationship between trust and regulation, whereas in the present paper the relationship between trust and climate policy is positive.
+This is in line with Bose, (2023), who shows that higher trust induces politicians to provide more public goods, in particular, more climate change policies.
+Besley and Dray, (2024) show that state effectiveness is greater when citizens trust their government, because higher trust increases voluntary compliance and thereby reduces implementation costs.
+Ehrmann, (2025) reviews the literature on trust in central banks and conclude that a high level of trust is an important prerequisite for a successful monetary policy.
 The paper is organized as follows.
-In Section 2, I describe the baseline model, which is then solved in Section 3.
+In Section 2, we describe the baseline model, which is then solved in Section 3.
+In Section 4, we discuss the main implication of the model.
 Section 5 extends the baseline model along several dimensions.
-In Section 5, I derive testable predictions of the model and have a first quick look at the data.
 Section 6 concludes.
@@
 2 The Model
-In this section I introduce the main building blocks of the model.
+In this section we introduce the main building blocks of the model.
 There is an unknown state of the world, $\omega\in\{0,1\}$, indicating whether climate change is mild, $\omega=0$, or severe, $\omega=1$.
@@
 There are two different kinds of actors.
-On the one hand, there are two kinds of purely office motivated politicians, indexed by $i=1,2$.
+On the one hand, there are two purely office motivated politicians, indexed by $i=1,2$.
 On the other hand, there is a continuum of voters of mass 1.
@@
 Voters policy preferences depend on the state of the world $\omega$. Given a winning policy $p$ and given a state realization $\omega$, voters’ realized policy utility is
-$u(p,\omega)=\left\{\begin{array}[]{cl}-p&\text{if }\omega=0,\\ -\Delta-\beta\left|p-1\right|&\text{if }\omega=1.\end{array}\right.$ (1)
+$u(p,\omega)=\left\{\begin{array}[]{cl}-p&\text{if }\omega=0,\\ -\Delta-b\left|p-1\right|&\text{if }\omega=1.\end{array}\right$. (1)
 Hence, for each state realization $\omega$, the optimal policy is $p=\omega$.
@@
 In particular, if $\omega=1$, i.e., when climate change is severe, then baseline welfare decreases by $\Delta>0$, independent of the chosen policy, while it remains constant if climate change is mild, $\omega=0$.
-The parameter $\beta>0$ measures the importance of choosing the right policy when climate change is severe: the greater is $\beta$, the more important is taking the right action.
-Note that the worst outcome is choosing policy incorrectly in state 1, yielding $-\Delta-\beta$, while the best outcome is choosing correctly in state 0, yielding a utility of zero.
-I make the following assumption to capture the idea that at the outset, voters believe the optimal policy choice is $p=0$: [^1] This assumption is not without loss of generality for the model’s conclusions.
-Allowing for $q\geq\frac{1}{1+\beta}$ would imply that the set of possible equilibria is larger.
-I chose to assume $q<\frac{1}{1+\beta}$, because it seems realistic that voters need to be convinced to support anti-climate change policies.
+The parameter $b>0$ measures the importance of choosing the right policy when climate change is severe: the greater is $b$, the more important is taking the right action.
+We make the following assumption to capture the idea that at the outset, voters believe the optimal policy choice is p=0p=0: [^2] This assumption is not without loss of generality for the model’s conclusions.
+Allowing for $q\geq\frac{1}{1+b}$ would imply that the set of possible equilibria is larger.
+We chose to assume $q<\frac{1}{1+b}$, because it seems realistic that voters need to be convinced to support anti-climate change policies.
 Moreover, as a second benefit, Assumption 1 keeps the paper’s exposition more concise.
 Assumption 1.
-Throughout I assume that $q<\frac{1}{1+\beta}.$
+Throughout we assume that $q<\frac{1}{1+b}.$
 At the beginning of the game, both types of players receive a signal that is informative about $\omega$.
-In particular, as in Gratton (2014), before announcing policy platforms every politician $i$ receives a signal $s_{i}^{P}$ that perfectly reveals $\omega$ Hence, after receiving $s_{i}^{P}$, each politician perfectly knows wether climate change is severe or not.
+In particular, as in Gratton, (2014), before announcing policy platforms every politician $i$ receives a signal $s_{i}^{P}$ that perfectly reveals $\omega$.
+Hence, after receiving $s_{i}^{P}$, each politician perfectly knows whether climate change is severe or not.
 Voters receive two kinds of signals.
@@
 On the other hand, they receive information from news or from direct experience about $\omega$.
-I model this by assuming that each voter receives a signal $s\in\mathbb{R}$.
+We model this by assuming that each voter receives a signal $s\in\mathbb{R}$.
 When the state is $\omega$, the c.d.f.
-of $s$ is $\Phi\left(\frac{s+(2\omega-1)\mu}{\sigma}\right)$, where $\Phi$ is the c.d.f.
+of $s$ is $\Phi\left(\frac{s-(2\omega-1)\mu}{\sigma}\right)$, where $\Phi$ is the c.d.f.
 of the standard Gaussian distribution and $\mu>0$ and $\sigma>0$ are parameters.
 Therefore, the typical monotone likelihood ratio property (MLRP) is satisfied, meaning that any $s>0$ is evidence for $\omega=1$ and increases the belief compared to the prior, while any $s<0$ is evidence for $\omega=0$ and decreases the belief.
-Following Callander (2011) and others, I interpret $\mu$ as a measure of the complexity of the issue climate change.
+Following Callander, (2011) and others, we interpret $\mu$ as a measure of the complexity of the issue climate change.
 If $\mu$ is small, the issue is very complex, and thus voters tend to hold imprecise beliefs about $\omega$.
@@
 Assumption 2.
-Throughout I assume that
-$\mu\geq\mu^{\text{Bayes}}\equiv\frac{\sigma\sqrt{\ln\left(\frac{1-q}{\beta q}\right)}}{\sqrt{2}}.$
-I assume that voters have anticipatory utility.
+Throughout we assume that
+$\mu\geq\hat{\mu}\equiv\frac{\sigma\sqrt{\ln\left(\frac{1-q}{bq}\right)}}{\sqrt{2}}.$
+We assume that voters derive anticipatory utility.
 In particular, a voter has expectations about her future utility, which is derived from a belief bout the true state and a second belief about the enacted policy.
-This causes anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare(Bénabou and Tirole, 2016).
+This causes anticipatory utility, “meaning that the individual experiences pleasant or aversive emotions from thinking about future welfare(Bénabou and Tirole,, 2016).
 To increases anticipatory utility (to be defined precisely below), a voters may use motivated reasoning, which means that she may update beliefs using a distorted complexity parameter $\tilde{\mu}$.
 Clearly, processing information using such a distorted complexity parameter comes at a cost, because higher anticipatory utility may imply lower utility because of imperfect decision making.
-As Bénabou and Tirole (2016) write, “one can react to bad news objectively, which leads to better decisions but having to live with grim prospects for some time […], or adopt a more “defensive” cognitive response that makes life easier until the day of reckoning, when mistakes will have to be paid for.” I do not model these costs in the baseline model to be able to identify the pure effect of motivated reasoning, but in Section 4.4 two models of distortions are discussed.
-If a voter is indifferent between having $\tilde{\mu}=\mu$ and some other $\tilde{\mu}\neq\mu$, I assume without loss of generality that she chooses $\tilde{\mu}^{*}=\mu$.
+As Bénabou and Tirole, (2016) write, “one can react to bad news objectively, which leads to better decisions but having to live with grim prospects for some time […], or adopt a more “defensive” cognitive response that makes life easier until the day of reckoning, when mistakes will have to be paid for.” We do not model these costs in the baseline model to be able to identify the pure effect of motivated reasoning, but in Section 5.1 two models of distortions are discussed.
+If a voter is indifferent between having $\tilde{\mu}=\mu$ and some other $\tilde{\mu}\neq\mu$, we assume without loss of generality that she chooses $\tilde{\mu}^{*}=\mu$.
 Choosing $\tilde{\mu}>\mu$ implies the voter interprets the signal as more informative than it really is, and hence beliefs will change excessively, while choosing $\tilde{\mu}<\mu$ implies a more conservative stance and that beliefs move less than they should.
-Based on $s$ and $\mathbf{p}$, and after choosing $\tilde{\mu}$, each voter forms beliefs $\pi(s,\mathbf{p},\tilde{\mu})$ and $\kappa$.
-$\pi(s,\mathbf{p},\tilde{\mu})$ is the posterior belief about $\omega$, whereas $\kappa$ is the probability assigned to the event that policy 1 will be implemented after the election.
-Because candidates commit to policy platforms, clearly $\kappa=1$ if $p_{1}=p_{2}=1$, and $\kappa=0$ if $p_{1}=p_{2}=0$.
+[…]
 Equipped with these beliefs, we can now calculate a voter’s anticipatory utility:
-$AU(s,\mathbf{p},\tilde{\mu})=-\kappa\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot\Delta+(1-\pi(s,\mathbf{p},\tilde{\mu}))\right]-(1-\kappa)\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot(\Delta+\beta)\right]$ (2)
+$AU(s,\mathbf{p},\tilde{\mu})=-\kappa\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot\Delta+(1-\pi(s,\mathbf{p},\tilde{\mu}))\right]-(1-\kappa)\left[\pi(s,\mathbf{p},\tilde{\mu})\cdot(\Delta+b)\right]$ (2)
 Thus, anticipatory utility equals the utility the voter expects to receive once policies are determined.
 Note that a voter is never pivotal, because there is a continuum of voters.
-I thus assume that voters vote sincerely as defined by Austen-Smith and Banks (1996).
+We thus assume that voters vote sincerely.
 In words, each voter forms beliefs based on $s$, $\mathbf{p}$, and $\tilde{\mu}$ and then votes to maximize (1).
-We are now in a position to define equilibrium of our game.
-(i) each voter maximizes $AU(s,\mathbf{p},\tilde{\mu})$ given a belief about the probability that policy 1 will be enacted $\kappa$; (ii) each voter votes sincerely given $s$ and $\tilde{\mu}$; (iii) $\kappa$ is correct given $s$, $\tilde{\mu}$, and given the strategies of all other voters; (iv) politicians choose policy platforms that maximize the probability to be elected, given their expectations about voters’ behavior.
-As typical in signaling games, there are multiple equilibria.
-I focus attention on symmetric equilibria, that is, equilibria in which candidates, holding the same information and trying to achieve the same, choose identical strategies.
-Further, if on the equilibrium path both candidates choose the same pure strategy, then we must have $p_{1}=p_{2}$.
-If, off the equilibrium path, voters observe $p_{1}\neq p_{2}$, I assume that they do not learn anything from this about $\omega$.
-That is, if off the equilibrium path $p_{1}\neq p_{2}$, then observing $\mathbf{p}$ does not change voters’ belief about $\omega$ [^2] This could be justified, for example, because candidates have types as well.
-An ideological candidate always chooses $p_{i}=i-1$, $i\in\{0,1\}$.
-A strategic candidate chooses the policy that is expected to maximize the probability to win the election.
-A model in which each candidate $i$ is ideological with probability $\gamma\in(0,1)$ and non-ideological with probability $1-\gamma$ would yield the same results regarding the platform choices of office-motivated candidates as the assumption of non-informative deviations.
+[…]
 3 Equilibrium
-In this section, I solve the baseline game.
-I begin by analyzing equilibrium play in the voting subgame.
-In Section 3.2, I analyze candidates’ incentives and equilibrium behavior, and in Section 3.3 I study the equilibrium of the whole game.
+In this section, we solve the baseline game.
+We begin by analyzing equilibrium play in the voting subgame.
+In Section 3.2, we analyze candidates’ incentives and equilibrium behavior, and in Section 3.3 we study the equilibrium of the whole game.
 3.1 Voter Behavior
-First I turn to the optimal motivated beliefs of voters.
+First we turn to the optimal motivated beliefs of voters.
 The focus is on situations in which $p_{1}\neq p_{2}$, because otherwise voters choose a winner, but not a winning policy.
@@
 $\pi(s,\mathbf{p},\tilde{\mu})=\dfrac{q\rho_{1}(1-\rho_{1})}{q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}}.$ (3)
-To see how a voter optimally chooses her motivated belief $\tilde{\mu}$, take the derivative of $AU(s,\mathbf{p},\tilde{\mu})$ with respect to $\tilde{\mu}$:
-$\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}=-\frac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[\beta+\Delta-\kappa(1+\beta)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}$ (4)
-If $\kappa=\tilde{\kappa}\equiv\frac{\beta+\Delta}{\beta+1}$, then this is equal to zero, independent of $s$.
-If $\Delta>1$, then $\tilde{\kappa}>1$, and because $\kappa$ is a belief, it must be true that $\kappa<\tilde{\kappa}$, implying that the sign of (4) is independent of $\kappa$.
-In particular, if this is the case, then $AU(s,\mathbf{p},\tilde{\mu})$ decreases in $\tilde{\mu}$ if $s>0$, and it increases in $\tilde{\mu}$ if $s<0$.
-Otherwise, that is, if $\Delta\leq 1$, then whether or not $AU(s,\mathbf{p},\tilde{\mu})$ increases in $\tilde{\mu}$ depends on both $s$ and the belief $\kappa$.
-Note that for each voter $\kappa$ can be interpreted as a measure of trust in government.
-$\kappa$ is the distribution of beliefs about the policies that will be enacted.
-If a voter believes that the government will always enact the policy that matches the true state, then $\kappa=\pi(s,\mathbf{p},\tilde{\mu})$.
-If, however, there is little trust in government regarding policy choices, then $\kappa\neq\pi(s,\mathbf{p},\tilde{\mu})$, and if trust is very low, we have $\kappa\in\{0,1\}$.
-We can now derive the optimally distorted complexity of the issue climate change:
+Note that this belief is not properly defined when one of the mixed strategies becomes a pure strategy, i.e., when $\rho_{\omega}\in\{0,1\}$.
+[^4] The reason is that there cannot be a pure symmetric pure strategy equilibrium in which candidates choose $p_{1}\neq p_{2}$.
+In a symmetric pure strategy equilibrium, platforms are perfectly informative about $\omega$.
+[…]
+$\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}=-\frac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\mu s}{\sigma^{2}}}}{\sigma^{2}\left((1-q)(1-\rho_{0}){\rho_{0}}+q(1-\rho_{1})\rho_{1}e^{\frac{2\mu s}{\sigma^{2}}}\right)^{2}}$ (4)
+[…]
+[…]
+Lemma 1 states for each voter the optimal distortion $\tilde{\mu}^{*}$ as a function of $s$ and possibly $\kappa$:
 Lemma 1.
-Let $\tilde{\kappa}\equiv\frac{\beta+\Delta}{\beta+1}$.
+Let $\tilde{\kappa}\equiv\frac{b+\Delta}{b+1}$.
   1. (a)
-    If $s=0$ or $\kappa=\tilde{\kappa}$, then generally $\tilde{\mu}^{*}=\mu$.
+    If $s=0$ or $\kappa=\tilde{\kappa}$, then $\tilde{\mu}^{*}=\mu$ for all $\Delta$.
   2. (b)
-    If $\Delta>1$, then a voter’s optimal distortion of $\mu$ is (i) $\tilde{\mu}^{*}=\infty$ if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
+    If $\Delta>1$, then (i) $\tilde{\mu}^{*}$$\infty$ if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
   3. (c)
-    If $\Delta=1$, then a voter’s optimal distortion of $\mu$, if $\kappa<1$, is (i) $\tilde{\mu}^{*}=\infty$, if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
+    If $\Delta=1$ and $\kappa<1$, (i) $\tilde{\mu}^{*}$$\infty$, if $s<0$, and (ii) $\tilde{\mu}^{*}=0$, if $s>0$.
   4. (d)
-    If $\Delta<1$, then a voter’s optimal distortion of $\mu$ is (i) $\tilde{\mu}^{*}=\infty$, if either $s<0$ and $\kappa<\tilde{\kappa}$, or $s>0$ and $\kappa>\tilde{\kappa}$, and (ii) $\tilde{\mu}^{*}=0$, if either $s<0$ and $\kappa>\tilde{\kappa}$, or $s>0$ and $\kappa<\tilde{\kappa}$.
-The lemma provides us with an important intermediate result, the optimal distortions of $\mu$ for different voters.
+    If $\Delta<1$, then (i) $\tilde{\mu}^{*}$$\infty$, if either [$s<0$ and $\kappa<\tilde{\kappa}$] or [$s>0$ and $\kappa>\tilde{\kappa}$], and (ii) $\tilde{\mu}^{*}=0$, if either [$s<0$ and $\kappa>\tilde{\kappa}$] or [$s>0$ and $\kappa<\tilde{\kappa}$].
+The lemma provides us with an important intermediate result.
 We see that beliefs about policy only matter if severe climate change causes only moderate baseline welfare losses.
-Otherwise, if severe climate change leads to catastrophical welfare losses, then any signal indicating that $\omega=1$ will be interpreted as pure noise and hence completely ignored.
-To the contrary, any signal indicating $\omega=0$ will be accepted as a perfect indication that climate change is indeed mild, independent of the signal’s strength.
-Finally, a voter with a signal $s=0$ or a voter with $\kappa=\tilde{\kappa}$ has no incentive to distort her signal’s precision and hence updates her belief about $\omega$ like a perfect Bayesian.
-What does this imply for the beliefs about $\omega$ the voters hold?
-Clearly, if $\tilde{\mu}=0$, then beliefs do not change at all, and $\pi(s,\mathbf{p},0)=q$.
-The same is true for a voter receiving a signal $s=0$, which is not informative about $\omega$.
-Moreover, if $\tilde{\mu}=\infty$, then any signal that contains only the slightest bit of information will completely move beliefs to the extremes, and hence $\pi(s,\mathbf{p},\infty)\in\{0,1\}$.
-Only if $s\neq 0$ and $\kappa=\tilde{\kappa}$ will the belief be a non-constant continuous function of $s$ and it equals $\pi(s,\mathbf{p},\mu)$:
+Otherwise, if severe climate change leads to catastrophical welfare losses, then any signal indicating that $\omega=1$ will be interpreted as pure noise.
+To the contrary, any signal indicating $\omega=0$ will be accepted as a perfect indication that climate change is indeed mild, independent of the signal’s real strength.
+What does this imply for the beliefs about $\omega$ the voters hold? Clearly, if $\tilde{\mu}=0$, then
+$\pi(s,\mathbf{p},0)=\frac{q\rho_{1}(1-\rho_{1})}{q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})}.$
+This is the same belief a Bayesian voter observing $s=0$ holds.
+Moreover, if $\tilde{\mu}$$\infty$, then any signal that contains only the slightest bit of information will completely move beliefs to the extremes, and hence $\pi(s,\mathbf{p},\infty)\in\{0,1\}$.
+Only if $\kappa=\tilde{\kappa}$ will the belief be a non-constant continuous function of $s$ and it equals $\pi(s,\mathbf{p},\mu)$:
 Corollary 1.
@@
   2. (b)
-    If $\Delta\leq 1$,
-    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}0&\text{if }\left(s<0\wedge\kappa<\tilde{\kappa}\right)\\ \hat{\pi}&\text{if }\left(s>0\wedge\kappa<\tilde{\kappa}\right)\vee\left(s<0\wedge\kappa>\tilde{\kappa}\right)\\ \pi(s,\mathbf{p},\mu)&\text{if }\kappa=\tilde{\kappa}\\ 1&\text{if }\left(s>0\wedge\kappa>\tilde{\kappa}\right)\end{array}\right.$
+    If $\Delta=1$, then
+    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}\pi(s,\mathbf{p},\mu)&\text{if }\kappa=1\\ \hat{\pi}&\text{if }s\geq 0\wedge\kappa<1\\ 0&\text{if }s<0\wedge\kappa<1\end{array}\right$.
+  3. (c)
+    If $\Delta<1$,
+    $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\left\{\begin{array}[]{cl}\pi(s,\mathbf{p},\mu)&\text{if }\kappa=\tilde{\kappa}\lor s=0\\ 0&\text{if }s<0\wedge\kappa<\tilde{\kappa}\\ \hat{\pi}&\text{if }\left(s>0\wedge\kappa<\tilde{\kappa}\right)\vee\left(s<0\wedge\kappa>\tilde{\kappa}\right)\\ 1&\text{if }s>0\wedge\kappa>\tilde{\kappa}\end{array}\right$.
 We now know the beliefs of all voters as functions of their signals and of their policy belief $\kappa$.
 Corollary 1 shows that if $\Delta>1$, and hence severe climate change has catastrophic consequences, then the equilibrium has a simple structure, and $\kappa$ actually plays no roll.
-Voters have two different beliefs, $\pi(s,\mathbf{p},\tilde{\mu})=0$ and $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$.
+Voters have two different beliefs, $\pi=0$ and $\pi=\hat{\pi}$.
 What does this imply for voters’ decisions at the ballot?
 Recall that voters vote sincerely for the alternative that they believe maximizes (1).
-The expected utility from policy 1 is $u(p=1)=-\pi(s,\mathbf{p},\tilde{\mu})\Delta-(1-\pi(s,\mathbf{p},\tilde{\mu}))$, while from policy 0 she gets $u(p=0)=-\pi(s,\mathbf{p},\tilde{\mu})(\Delta+\beta)$.
+The expected utility from policy 1 is $u(p=1)=-\pi\Delta-(1-\pi)$, while from policy 0 she gets $u(p=0)=-\pi(\Delta+b)$.
 Hence, the voter cast her ballot for policy 1 iff
-$u(p=1)>u(p=0)\Leftrightarrow\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}\equiv\frac{1}{1+\beta}.$ (5)
-If the reverse is true, then policy 0 is strictly preferred, and if $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$, then a voter is indifferent.
-Clearly, $01>\tilde{\pi}>$ 0.
-Note that by Assumption 1 we have $\tilde{\pi}>q$.
-If $\Delta>1$, then voters hold two different beliefs, either $\pi(s,\mathbf{p},\tilde{\mu})=0$ or $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$.
-A voter with the former belief always votes for policy 0, while the decision at the ballot of the other voter depends on $q$, $\beta$, as well as on the politicians’ strategies $\rho_{0}$ and $\rho_{1}$.
+$u(p=1)>u(p=0)\Leftrightarrow\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}\equiv\frac{1}{1+b}.$ (5)
+If the reverse is true, then policy 0 is strictly preferred, and if $\pi=\tilde{\pi}$, then a voter is indifferent. Clearly, $1>\tilde{\pi}>0$. Note that by Assumption 1 we have $\tilde{\pi}>q$.
+If $\Delta>1$, then voters hold two different beliefs, either $\pi=0$ or $\pi=\hat{\pi}$.
+A voter with the former belief always votes for policy 0, while the decision at the ballot of the other voter depends on $q$, $b$, as well as on the politicians’ strategies $\rho_{0}$ and $\rho_{1}$.
 If $\hat{\pi}>\tilde{\pi}$, then these voters vote for policy 1, and as a consequence all voters vote informatively, i.e., they vote according to their signals.
@@
 If, however, $\hat{\pi}\leq\tilde{\pi}$, then a majority of voters always supports policy 0, implying it wins independent of the state.
-Things are slightly different when $\Delta\leq 1$.
-On the one hand, the equilibrium just described continues to exist also when severe climate change decreases baseline welfare only moderately.
-In particular, this is the case if $\hat{\pi}\leq\tilde{\pi}$ and $\kappa<\tilde{\kappa}$ for sufficiently many $s$.
-To the contrary, if $\hat{\pi}\geq\tilde{\pi}$ and $\kappa>\tilde{\kappa}$ for sufficiently many $s$, then policy 1 wins independent of the state.
-Hence, when severe climate change is not catastrophic, then there might exist an equilibrium in which politicians always campaign on the policy tailored for severe climate change, independent of the true state.
-It follows that we cannot yet exclude that, when $\Delta\leq 1$, a second inefficient equilibrium may exist.
-What about the efficient equilibrium, in which candidates choose the optimal policies given $\omega$?
+If $\Delta=1$, beliefs are the same except when $\kappa=1$, in which case a voter holds a Bayesian belief about $\omega$ and votes accordingly.
+Things are slightly different when $\Delta<1$, as now $\kappa$ starts to matter.
+As before, a situation in which all voters hold a belief $\kappa<\tilde{\kappa}$, leading to the above discussed situation of all voters voting for policy 0 if faced with a choice, continues to exist.
+On the other hand, if $\hat{\pi}\geq\tilde{\pi}$ and $\kappa>\tilde{\kappa}$ for sufficiently many $s$, then policy 1 wins independent of the state.
+What about an efficient outcome, in which voters collectively choose the policy that matches the state $\omega$?
 Assume each voters wants to believe her signal and thus has $\kappa<\tilde{\kappa}$ when $s<0$ and $\kappa>\tilde{\kappa}$ when $s>0$.
 Then every voter votes informatively, and therefore the policy matching the state wins for all $\hat{\pi}\in(0,1)$.
-Hence, if even severe climate change decreases baseline welfare only moderately, then there always may exist an equilibrium in which voters vote informatively, and hence the correct policy is chosen with probability 1.
+Hence, if even severe climate change decreases baseline welfare only moderately, then there may exist an equilibrium in which voters vote informatively, and hence the correct policy is chosen with probability 1.
 We summarize the results of this section in our next proposition:
 Proposition 1.
-Assume $\mathbf{p}\in\left\{\left(0,1\right),\left(1,0\right)\right\}$.
+Assume $p_{1}\neq p_{2}$.
   1. 1.
-    If $\Delta>1$, then there is a unique subgame equilibrium.
+    If $\Delta\geq 1$, then there is a unique equilibrium.
     1. (a)
-      If $\hat{\pi}\leq\tilde{\pi}$, then $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ for all $s$ and a majority of voters always votes for policy 1.
+      If $\hat{\pi}\leq\tilde{\pi}$, then $\kappa^{*}=0$ for all $s$ and a majority of voters votes for policy 0.
     2. (b)
-      If $\hat{\pi}>\tilde{\pi}$, then $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$ and $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=1$ if $s\geq 0$, and a majority always votes for the policy $p_{i}$ that matches the true state $\omega$.
+      If $\hat{\pi}>\tilde{\pi}$, then $\kappa^{*}=0$ if $s<0$ and $\kappa^{*}=\hat{\pi}$ if $s\geq 0$, and a majority votes for $p=\omega$.
   2. 2.
-    If $\Delta\leq 1$, then there exist multiple subgame equilibria.
+    If $\Delta<1$, then there exist multiple equilibria.
     1. (a)
-      For any $\hat{\pi}\in(0,1)$, there exists a subgame equilibrium with $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$ and $\kappa^{*}(s,\mathbf{p},\tilde{\mu})=1$ if $s\geq 0$, and a majority of voters always votes for the policy $p_{i}$ that matches the true state $\omega$.
+      For any $\hat{\pi}\in(0,1)$, there exists a subgame equilibrium with $\kappa^{*}=0$ if $s<0$ and $\kappa^{*}=1$ if $s>0$, and a majority of voters always votes for policy $p=\omega$.
     2. (b)
-      For $\hat{\pi}\geq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=1$ and policy 1 always wins.
+      For $\hat{\pi}\geq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa^{*}=1$ and policy 1 always wins.
     3. (c)
-      For $\hat{\pi}\leq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=0$ and policy 0 always wins.
+      For $\hat{\pi}\leq\tilde{\pi}$, there exists a subgame equilibrium with $\kappa^{*}=0$ and policy 0 always wins.
     4. (d)
-      There is no subgame equilibrium in which the policy $p_{i}$, that does not match the state $\omega$, always wins.
-On the positive side, there always exists a subgame equilibrium in which information aggregates, if severe climate change decreases baseline welfare not too much, i.e., if $\Delta$ is small.
-However, this equilibrium depends on the expectations about the efficiency of political process.
+      There is no subgame equilibrium in which policy $p\neq\omega$ always wins.
+There always exists a subgame equilibrium in which information aggregates, if severe climate change decreases baseline welfare not too much, i.e., if $\Delta$ is small.
+However, this equilibrium depends on the expectations about the efficiency of the political process.
 If voter believe that the policy suggested by their signal is chosen, then the election leads to efficient results.
-However, if voters are pessimistic, and believe policy 0 will be chosen no matter what is the true state $\omega$, then this becomes a self-fulfilling prophecy if $\hat{\pi}$ is low.
+But if voters are pessimistic, and believe policy 0 will be chosen no matter what is the true state $\omega$, then this becomes a self-fulfilling prophecy if $\hat{\pi}$ is low.
 When $\Delta$ becomes larger, and hence severe climate change leads to more dire welfare losses, then the efficient equilibrium disappears when $\hat{\pi}$ is small.
@@
 We now study the platform choices of the candidates as a function of their signals.
-Of course, when they decide which policies to offer they think ahead which policies are likely to lead to electoral success, and therefore voters’ reactions, in particular the one’s formalized in Proposition 1, matter for candidates’ incentives.
+Of course, when they decide which policies to offer, they reason forward thinking about which policy is more likely to lead to electoral success.
+Therefore voters’ reactions to policy choices, in particular the one’s formalized in Proposition 1, matter for candidates’ incentives.
 From the perspective of an office motivated candidate, the optimal platform is the one that maximizes the chance to win the election.
@@
 Focussing on symmetric equilibria, the probability that each candidate choose policy 1 in state $\omega$ is $\rho_{\omega}$ as before.
-Note that if $\rho_{1}\neq\rho_{2}$, voters learn from platform choices even if $p_{1}\neq p_{2}$.
+If $\rho_{1}\neq\rho_{2}$, voters learn from platform choices even if $p_{1}\neq p_{2}$.
 Choosing such a mixed strategy can only be an equilibrium if it leads to a chance of winning of 50%, because candidates need to be indifferent.
@@
 Hence, we need to have $\kappa=\frac{1}{2}$ for all $s$.
-If $\tilde{\kappa}<\frac{1}{2}$, any voter with $s>0$ will hold belief $\pi(s,\mathbf{p},\tilde{\mu})=1$ and thus they all vote for 1.
+If $\tilde{\kappa}<\frac{1}{2}$, any voter with $s>0$ will hold belief $\pi=1$ and thus they all vote for 1.
 But this means that in state $\omega=1$, the candidate offering $p=1$ wins with certainty, and thus this cannot be equilibrium.
-If $\tilde{\kappa}>\frac{1}{2}$, any voter with $s<0$ will hold belief $\pi(s,\mathbf{p},\tilde{\mu})=0$ and thus they all vote for policy 0.
+If $\tilde{\kappa}>\frac{1}{2}$, any voter with $s<0$ will hold belief $\pi=0$ and thus they all vote for policy 0.
 But this means that in state $\omega=0$ the candidate offering $p=0$ wins with certainty, and thus this cannot be equilibrium, either.
 Finally, if $\tilde{\kappa}=\frac{1}{2}$, any voter will choose $\tilde{\mu}^{*}(s)=\mu$.
-But then information aggregates by Assumption 2, and thus this cannot be an equilibrium belief, either.
+But then information aggregates by Assumption 2, and thus also this cannot be an equilibrium belief.
 Hence, we can conclude that no mixed strategy equilibrium can exist.
-What about pure strategy equilibria?
-In a symmetric equilibrium, we must have $p_{1}=p_{2}$, because candidates hold the same information.
-Suppose voters observe off-equilibrium that $p_{1}\neq p_{2}$ and both candidates expect voters to vote informatively.
-Hence, they vote for policy 1 if $s>0$, for policy 0 if $s<0$, and they randomize when $s=0$.
-Then, in each state $\omega$, the candidate offering the policy that matches the state wins the election.
-Hence, when choosing which policy platform to offer, candidate $i$ knows that $p_{i}=\omega$ wins the election if $p_{-i}\neq\omega$ and otherwise both win with a probability of 50%.
-To the contrary, choosing $p_{i}\neq\omega$ wins the election with 50% if also $p_{-i}\neq\omega$ and otherwise loses the election for sure.
-Hence, $p_{i}=s_{i}$ is the unique optimal strategy.
-It follows that if sufficiently many voters vote informative, candidates will choose informative policy platforms.
-To the contrary, if candidates expect a sufficiently large number of voters to not vote informatively when $p_{1}\neq p_{2}$, then there exists a policy $\hat{p}\in\{0,1\}$ that wins the election with certainty, independent of the state $\omega$.
-Candidates then know they win the election by choosing $p_{i}=\hat{p}$ if $p_{-i}\neq\hat{p}$, and they win with a probability of 50% if $p_{-i}=\hat{p}$.
-Choosing $p_{i}\neq\hat{p}$ will lose the election for sure if $p_{-i}=\hat{p}$, and hence both candidates are best served by offering $\hat{p}$.
-This leads to the next formal result:
+[…]
+This leads to the next result:
 Proposition 2.
-There is a unique symmetric equilibrium in the platform choice stage, which is in pure strategies.
+For given $\tilde{\mu}^{*}(s)$ for all $s$, there is a unique symmetric equilibrium in the platform choice stage, which is in pure strategies.
 Candidates choose $p_{i}^{*}=\omega$, if they expect a majority of voters to vote for the policy that matches the state, when $p_{1}\neq p_{2}$.
@@
 We can now determine equilibrium platform choices as functions of the parameters of the game.
-If severe climate change leads to catastrophic baseline welfare level losses, $\Delta>1$, then voters ignore any signal indicating that $\omega=1$.
-Our analysis so far reveals that then a majority always votes for policy 0, whenever $p_{1}\neq p_{2}$.
+If severe climate change leads to catastrophic baseline welfare level losses, $\Delta\geq 1$, then voters ignore any $s>0$.
+Our analysis reveals that then a majority always votes for policy 0, whenever $p_{1}\neq p_{2}$.
 Proposition 2 tells us that in this case candidates never campaign on policy $p=1$, independent of the information they hold.
 Hence, if severe climate change is catastrophic, then candidates will ignore any information and always choose the optimal policy for mild climate change, $\mathbf{p}^{*}=(0,0)$.
-When $\Delta\leq 1$, the set of equilibria becomes larger.
+When $\Delta<1$, the set of equilibria becomes larger.
 The reason is that now voters’ beliefs about the enacted policy become self-fulfilling.
-If voters trust the government in the sense that they believe the policy that is optimal given $\omega$ wins, then voters always vote informatively, implying that $\mathbf{p}^{*}=(\omega,\omega)$ can indeed be equilibrium.
-However, there always coexists another equilibrium, in which candidates again ignore their information.
-If voters are convinced that $p=0$ is chosen, then any voter with $s<0$ will choose $\tilde{\mu}=\infty$, while a voter with $s>0$ chooses $\tilde{\mu}=0$.
-Therefore, as in the case of $\Delta>1$, a majority always votes for policy 0, and candidates choose $\mathbf{p}^{*}=(0,0)$.
-The next proposition is the paper’s main result and formalizes the above intuitions:
+If voters trust the government in the sense that they believe the policy that is optimal given $\omega$ wins, then voters always vote informatively, implying that $\mathbf{p}^{*}=(\omega,\omega)$ is indeed an equilibrium.
+However, there always coexists the equilibrium in which which candidates ignore their information and choose $\mathbf{p}^{*}=(0,0)$.
+The next proposition states the paper’s main result and formalizes the above intuitions:
 Proposition 3.
-If $\Delta>1$, then there exists a unique pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
-If instead $\Delta\leq 1$, then there are multiple equilibria:
-  1. 1.
-    If $\Delta\leq 1$ and voters have trust in the government, then there exists a pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=\pi(s,\mathbf{p},\tilde{\mu}^{*})$ and candidates choose $\mathbf{p}^{*}=(\omega,\omega)$.
-  2. 2.
-    If $\Delta\leq 1$ and voters have no trust in the government, then there exists a pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
+If $\Delta\geq 1$, then there exists a unique pure strategy equilibrium in which $\kappa^{*}=0$ for all $s$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
+If instead $\Delta<1$, then there exist multiple equilibria.
+In particular, there exists one pure strategy equilibrium in which, for each $s$, $\kappa^{*}=\pi(s,\mathbf{p},\infty)$ and candidates choose $\mathbf{p}^{*}=(\omega,\omega)$, and there exists another pure strategy equilibrium in which $\kappa^{*}=0$ for all $s$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
+The analysis shows that when motivated beliefs are present, a self-fulfilling efficient equilibrium exists as long as the worst consequences of climate change are not too severe, i.e., $\Delta<1$.
+In this equilibrium, each voter takes her signal at face value and therefore votes informatively.
+As a result, candidates have the correct incentives and campaign on the policy that is optimal given $\omega$.
+Consequently, the probability that the winning policy matches $\omega$ is 1.
+However, there also always exists a bury-your-head-in-the-sand equilibrium, independent of $\Delta$.
+In this equilibrium, voters choose to ignore any signal $s>0$, while treating any signal $s<0$ as conclusive evidence that $\omega=0$.
+The ex ante probability that the correct policy is chosen in this equilibrium is 1q1-q.
+4 Discussion and Implications
+In this section, we discuss two important implications of the model: that the efficiency of policy-making in the face of climate change depends on both rhetoric and trust in government.
+4.1 Trust in Government
+When anticipated welfare losses from severe climate change are sufficiently large, the model admits a unique, inefficient equilibrium.
+By contrast, when these losses are moderate, multiple equilibria arise.
+In this case, equilibrium selection depends critically on voters’ expectations about how policy choices relate to the underlying state of the world, which we interpret as trust in government.
+Following Cabral, (2005) and Besley and Dray, (2024), we define trust as voters’ expectations that political institutions implement policies that are responsive to the true state.
+In the model, voters hold two distinct beliefs: a belief $\pi$ about the state of the world and a belief $\kappa$ about the probability that policy 1 will be implemented.
+In the efficient equilibrium with moderate welfare losses, voters’ beliefs about the state take extreme values.
+As shown in Section 3, motivated belief formation implies that $\pi\in\{0,1\}$, with $\pi=1$ when the voter’s signal satisfies $s>0$ and $\pi=0$ when $s<0$.
+In this environment, trusting the political process means that voters expect policy choices to mirror the realized state.
+Since policy 1 is optimal if and only if climate change is severe, a voter who believes $\pi=1$ expects policy 1 to be implemented with probability one, whereas a voter who believes $\pi=0$ expects policy 1 to be implemented with probability zero.
+It therefore follows directly in this equilibrium that the perceived probability that policy 1 is implemented coincides with the belief that the state is severe, so that $\kappa=\pi$.
+When welfare losses from severe climate change are moderate and voters have trust, they optimally take their signals at face value and vote informatively.
+Anticipating this behavior, politicians have incentives to follow their information and propose welfare-maximizing policies, thereby sustaining the efficient equilibrium.
+Trust thus operates as a coordination device: by aligning voters’ expectations about policy implementation with their beliefs about the state, it ensures that information is processed and aggregated efficiently.
+Conversely, when trust is low, voters expect policy choices to be weakly related or unrelated to the underlying state, so that $\kappa$ becomes insensitive to information contained in signals.
+Such pessimistic expectations weaken incentives to process information accurately and can become self-confirming.
+Anticipating that voters will not respond informatively, politicians optimally ignore their own information and campaign on policies designed for mild climate change, even when climate change is severe.
+In this way, low trust sustains an inefficient equilibrium.
+The model therefore implies a positive relationship between trust in government and the likelihood that electoral competition leads to effective climate policy, provided that welfare losses from severe climate change are not too large.
+This implication is consistent with Figure 1 and recent empirical evidence showing that higher trust in political institutions is associated with greater provision of public goods, including climate policy (Bose,, 2023).
+Importantly, this relationship arises only through the interaction between trust and motivated belief formation; absent motivated reasoning, voters’ beliefs about the state would be independent of expectations about policy implementation, and trust would not affect equilibrium policy choices.
+4.2 The Importance of Rhetoric
+A first implication concerns political rhetoric.
+In the model, an efficient equilibrium is less likely to exist when the anticipated welfare losses from severe climate change are sufficiently large.
+The reason is that larger anticipated losses increase anticipatory disutility and thereby strengthen incentives for voters to engage in defensive belief formation, which undermines coordination on the efficient outcome.
+An indirect corollary is that the way climate risks are described may matter for equilibrium selection.
+To the extent that rhetoric affects voters’ expectations about the magnitude of potential welfare losses from climate change, it can shift the strength of anticipatory disutility and, through this channel, the extent of belief distortion.
+In this sense, highly alarming descriptions need not translate into greater support for policy action and may instead trigger defensive responses, echoing the concern raised by Clayton et al., (2017).
+This argument does not require taking a stand on any particular model of communication.
+One interpretation is that rhetoric operates through framing, in the sense that equivalent information can induce systematically different perceptions depending on how outcomes are described (Tversky and Kahneman,, 1981, 1989; Ellingsen et al.,, 2012; Spiegler,, 2014).
+Another interpretation is that rhetoric is informative, in line with cheap-talk and persuasion frameworks in which language conveys signals about states or consequences (Crawford and Sobel,, 1982; Kartik and Van Weelden,, 2019; Kamenica and Gentzkow,, 2011; Alonso and Câmara,, 2016).
+The present model highlights a simple channel common to both interpretations: by shaping perceived severity, rhetoric can affect anticipatory utility and thus the mapping from information to policy outcomes.
 5 Extensions
-In this section I informally discuss further extensions that seem relevant.
-For the sake of brevity and clarity, throughout this section I assume that $\Delta>1$.
-In Section 4.1 I discuss how conclusions change if candidates are not only office motivated, but care also about welfare.
-In Section 4.2 I discuss the implications of differences in candidate valence.
-In Section 4.3, I discuss the implication of misinformation campaigns.
-Finally, in Section 4.4 I study the implication of costly deviations from Bayesian rationality.
+In this section we informally discuss further extensions that seem relevant.
+For the sake of brevity and clarity, throughout this section we assume that $\Delta>1$.
+In Section 5.1 we study the implication of costly deviations from Bayesian information processing.
+In Section 5.2 we discuss how conclusions change if candidates are not only office motivated, but care also about welfare.
+5.1 Costly Distortions
+In this section, we extend the previous analysis by accounting for the costs that a voter may incur when distorting the issue’s complexity.
+Introducing costs allows us to scrutinize the above predictions about how beliefs may be distorted, and it enables us to assess the robustness of the earlier results on policy choices.
+We distinguish between two types of costs.
+First, we analyze utility costs from making bad decisions (Bénabou and Tirole,, 2002, 2016; Brunnermeier and Parker,, 2005).
+We then study purely cognitive costs from choosing distorted beliefs (Levy, (2014); Le Yaouanq, (2023); Engelmann et al., (2024)).
+For proofs of the statements, see Appendix B.
+5.1.1 Cost from Imperfect Decision Making
+Assume each voter places a weight $\lambda\in[0,1]$ on anticipatory utility and $1-\lambda$ on the expected utility loss relative to the Bayesian-optimal action.
+In particular, voters aim at maximizing $W^{UC}=\lambda AU+(1-\lambda)C^{U}$, where $C^{U}\leq 0$ is the utility cost of deviating from Bayesian information processing.
+[…]
+If the distortion does change behavior, it has an expected cost.
+This cost is the difference between expected utility under Bayesian beliefs with the optimal action and the expected utility under Bayesian belief with the suboptimal action.
+The only voters for whom distortion may change the action are those with sufficiently high signals, $s>\tilde{s}$.
+When deciding if and how to distort beliefs, a voter faces two potentially optimal choices: (i) $\tilde{\mu}=0$, and thus interpreting any signal as irrelevant; and (ii) choosing the smallest possible $\tilde{\mu}$ such that the choice at the ballot is not altered but anticipatory utility increases.
+[^5] This option strictly dominates $\tilde{\mu}=\mu$.
+The voter’s welfare from (i) is
+$W^{UC}(\tilde{\mu}=0)=-q(b+\Delta)\lambda-\left(b-\frac{(1-q)(1+b)}{1-\big(1-e^{\frac{2s\mu}{\sigma^{2}}}\big)q}\right)(1-\lambda).$
+To determine the utility from (ii), note that the belief about $\omega$ given this must equal $1/(1+b)$. Thus,
+$W^{UC}(\tilde{\mu}=\mu_{c})=-\frac{(b+\Delta)\lambda}{1+b}.$
+When comparing these two welfare levels, note that when $s$ is such that a Bayesian voter is indifferent between voting for policy 0 or policy $1$, the utility cost from distortion is zero, and thus $W^{UC}(\tilde{\mu}=0)>W^{UC}(\tilde{\mu}=\mu_{c})$.
+Moreover, $W^{UC}(\tilde{\mu}=0)$ is strictly decreasing in $s$.
+The reason is that if a voter has a very strong signal, the Bayesian belief is very strong, and the utility cost from acting against this belief is high.
+Therefore, if $W^{UC}(\tilde{\mu}=0)>W^{UC}(\tilde{\mu}=0)$ even when $s\to\infty$, then any voter with $s>0$ will choose $\tilde{\mu}=0$.
+This is true if $\lambda$ is sufficiently large.
+In particular, if
+$\lambda>\frac{b(1+b)}{b\big(2-q+(1-q)b\big)+\big(1-q-qb\big)\Delta}\in(0,1).$
+Otherwise, for lower levels of $\lambda$, a voter with a very strong signal will choose the non-consequential distortion $\mu_{c}$ and vote according to the signal.
+Any other voter with $s>0$, however, chooses $\tilde{\mu}=0$, thereby ignoring the signal and voting for policy 0.
+Hence, the extended model of motivated beliefs provides a more nuanced prediction of voter behavior.
+If $\lambda$ is large, introducing costs changes nothing.
+However, if $\lambda$ is sufficiently small, then a strong signal will be interpreted as evidence of severe climate change, and voters receiving such a signal will vote for policy 1.
+Denote the threshold for a sufficiently strong signal by $\hat{s}$ and note that $\hat{s}>\tilde{s}$.
+To shift the equilibrium toward one in which politicians offer policies for severe climate change, this is not yet sufficient.
+The reason is that if $\mu$ is small—i.e., if climate change is a highly complex issue—then only few voters receive $s>\tilde{s}$.
+Thus, in this case, there exists a unique equilibrium in which no politician ever offers policy 1.
+However, for large enough $\mu$, more than 50% of voters receive $s>\tilde{s}$ in state 1, and thus a majority votes for policy 1.
+This creates incentives for politicians to follow their information and offer policies targeting severe climate change.
+5.1.2 Cognitive Costs
+Next consider pure cognitive costs from distorting information processing.
+Distorted information processing can lead to cognitive costs because, for example, sustaining self-serving or overly optimistic beliefs may conflict with one’s deeper desire for accuracy and coherence, creating dissonance and reducing overall well-being.
+Hence, in the following we assume that there is a cost of $C^{C}(\tilde{\mu},\mu)$ associated with choosing $\tilde{\mu}$.
+[…]
+$C^{C}(\tilde{\mu},\mu)=\left\{\begin{array}[]{ll}\frac{c}{2}\left(\tilde{\mu}-\mu\right)^{2}&\text{if }\tilde{\mu}\geq\mu\\ \frac{c}{2}\left(\frac{\mu}{\tilde{\mu}}\left(\tilde{\mu}-\mu\right)\right)^{2}&\text{if }\tilde{\mu}<\mu\end{array}\right$.
+(6)
+With this cost function we have $\left.C^{C}(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\delta\cdot\mu}=\left.C^{C}(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\delta}=\frac{c}{2}\mu^{2}\left(\delta-1\right)^{2}$ for all $\delta>1$.
+In the left panel of Figure 2, the cost function is plotted for $\mu=1$ and $c=2$.
+One property of it is that $\lim_{\tilde{\mu}\rightarrow\infty}C^{C}(\tilde{\mu},\mu)=\lim_{\tilde{\mu}\rightarrow 0}C^{C}(\tilde{\mu},\mu)=\infty$, and therefore we have $\tilde{\mu}\in(0,\infty)$ for any $s$.
+When deciding how to optimally distort $\mu$, each voter aims to maximize
+Figure 2: Left panel: Cost function when $\mu=1$ and $c=2$. Right panel: $\tilde{\mu}$ as a function of $s$ for $\mu=1$ and different values of $c$. Higher $c$ moves $\tilde{\mu}$ closer to $\mu$.
+$W^{CC}=AU(s,\mathbf{p},\tilde{\mu})-C^{C}(\tilde{\mu},\mu).$
+[…]
+If $s<0$, voters choose optimally at the ballot despite $\tilde{\mu}\neq\mu$.
+For $s>0$, this is not necessarily the case.
+These voters interpret signals as less informative than they truly are and may therefore cast a ballot for policy 0 despite $s>0$.
+Intuitively, as $c$ increases, voters choose to distort $\mu$ less.
+Consequently, $\lim_{c\to\infty}\tilde{\mu}^{*}(s)=\mu$.
+In this case, a majority will always cast a ballot for $p=\omega$, giving candidates an incentive to heed their information.
+However, if $c$ is small, the outcome reverts to the result in Proposition 3.
+By continuity, there exists a $\hat{c}>0$ such that if $c\in[0,\hat{c}$], motivated beliefs lead to voting behavior that prevents candidates from following their information.
+Conversely, if $c>\hat{c}$, then despite choosing $\tilde{\mu}^{*}\neq\mu$, a majority of voters will vote for policy 1 in state 1, thereby creating incentives for candidates to follow their information and campaign on policy 1.
 5.2 Policy Motivated Politicians
@@
 Candidate 1 has a strict incentive to also propose $p_{1}=1$ if
-$\frac{V}{2}-\Delta\geq V-\Delta-\beta\Leftrightarrow\beta\geq\frac{V}{2}.$
-If, however, $\beta<\frac{V}{2}$, there remains a unique pure strategy equilibrium with $\mathbf{p}^{*}=(0,0)$.
-Thus, if policy motivation is sufficiently strong, the previously identified problems may not occur.
-However, there is always another equilibrium in pure strategies where candidates continue to ignore their information despite policy motivation.
+$\frac{V}{2}-\Delta\geq V-\Delta-b\Leftrightarrow b\geq\frac{V}{2}.$
+If, however, $b<\frac{V}{2}$, there remains a unique pure strategy equilibrium with $\mathbf{p}^{*}=(0,0)$.
+Moreover, increasing the spoils of office $V$ decreases the chance for an efficient equilibrium to exist.
+However, even if the spoils of office are sufficiently low, the inefficient equilibrium continues to exist.
 Assume candidate 1 expects candidate 2 to choose $p_{2}=0$.
-In this case, the unique best response is to also choose $p_{1}=0$, because policy 0 now wins with certainty.
-Although candidate 1 cannot influence the policy, she can increase her chances of winning.
+In this case, the unique best response is to also choose $p_{1}=0$.
+The reason is that now policy 0 wins with certainty.
+Candidate 1 cannot influence the policy, but she can increase her chances of winning.
 Therefore, despite policy motivations, and even in the absence of any office motivation, there always exists an equilibrium with $\mathbf{p}^{*}=(0,0)$.
-4.2 Valence Differences
-Next, consider valence differences as in Ansolabehere and Snyder (2000), Groseclose (2001), Aragones and Palfrey (2002), and more recently Denter (2021).
-Specifically, assume that voters evaluate candidates not only by the policies they offer but by $v_{i}+u(p_{i},\omega)$, where $v_{i}\in\mathbb{R}$ is the valence utility the voter gets from candidate $i$ in the office and $u(p,\omega)$ is policy utility and defined as before in (1).
-Let $v_{1}>v_{2}$, implying that candidate 1 has a valence advantage.
-How does introducing valence alter the model’s conclusions?
-On the one hand, it has no effect on $\tilde{\mu}^{*}(s)$, which remains as described in Lemma 1, because valence is independent of the state $\omega$.
-However, valence does affect a voter’s choice at the ballot box for a given belief $\pi(s,\mathbf{p},\tilde{\mu})$.
-If both candidates choose the same policy, candidate 1 always wins because $v_{1}>v_{2}$.
-Equilibrium in the extended game resembles equilibrium in Ansolabehere and Snyder (2000).
-If the valence advantage is not too large, candidate 1 will always choose $p_{1}=0$, while candidate 2 may choose any $p_{2}\in\{0,1\}$ without having a chance to win.
-If the valence advantage is large enough, candidate 1 may choose any $p_{1}\in\{0,1\}$ and still win with certainty.
-Thus, it is possible that a candidate with a large valence advantage might use this advantage to propose an optimal policy.
-In fact, if we combine a large valence advantage with policy motivation as described in Section 4.1, then it is a dominant strategy for candidate 1 to choose $p_{1}=\omega$, and this policy indeed wins the election.
-4.3 Disinformation Campaigns
-The analysis so far has revealed a robust mechanism through which the electoral incentives of candidates are sufficient to prevent the implementation of adequate policies against climate change.
-However, climate change is also one of the topics where organized misinformation is most prevalent (see, for example, Lewandowsky, 2021).
-Perhaps the most famous example of such campaigns is that of ExxonMobil, which, despite having very precise knowledge about the causes and consequences of climate change (see Supran, Rahmstorf, and Oreskes, 2023), chose to downplay the phenomenon in their public communications and editorial-style advertisements in The New York Times (see Supran and Oreskes, 2017).
-I follow Papanastasiou (2020), Sisak and Denter (2024) and Denter and Ginzburg (2024) to incorporate misinformation into the model.
-In particular, I assume that each voter receives a fake signal with a probability of $\alpha$, while with the complementary probability of $1-\alpha$ the signal comes from a proper source and is informative.
-An informative signal is drawn from the distribution $F_{\omega}$ with corresponding density $f_{\omega}$.
-I assume that $f_{1}(s)/f_{0}(s)$ is an strictly increasing function of $s$, which is the standard MLRP assumption.
-Note that in the model described in Section 2 with normally distributed signals this is automatically satisfied.
-Moreover, without loss of generality let $f_{1}(0)=f_{0}(0)$, implying that $s=0$ is an uninformative signal as before.
-Unlike proper signals, a fake signal is not informative.
-In particular, I assume that the distribution of fake signals is $G(s)$ which is independent of $\omega$.
-For simplicity, further assume that $F_{\theta}$ and $G$ have full support on $\mathbb{R}$.
-Because the literature has shown that in the presence of fake news motivated reasoning is often regarding the reliability of an information source, voters cannot change the precision of a signal anymore, but have motivated beliefs about the probability that a given signal is fake or proper.
-In particular, they can choose $\tilde{\alpha}\in[0,1]$ instead of $\tilde{\mu}\in[0,\infty]$.
-For a given signal $s$ and a given choice $\tilde{\alpha}$, the belief about the state $\omega$ is
-$\pi(s,\mathbf{p},\tilde{\alpha})=\frac{q\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right]}{q\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right]+(1-q)\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right]}.$
-Choosing $\hat{\alpha}=0$ means ignoring the chance that the signal could be fake, and hence the belief equals the perfect Bayesian belief when no fake signals exist.
-To the contrary, setting $\hat{\alpha}=1$ means ignoring the signal completely because it originates from a fake source.
-Lemma 2.
-The optimal motivated belief about $\alpha$ is
-$\hat{\alpha}^{*}(s)=\left\{\begin{array}[]{cl}1&\text{if }s>0\\ \alpha&\text{if }s=0\\ 0&\text{if }s<0\end{array}\right.$
-Hence, any bad news about climate change will be interpreted as fake, while any good news will be interpreted as proper and hence credible.
-This is in line with recent empirical evidence (Thaler, 2024).
-What are the consequences for political decision making?
-The political equilibrium is the same as in Section 3.
-The reason is that any signal indicating climate change is severe will be ignored, and hence a majority of the electorate will prefer policy 00 in each state.
-Proposition 4.
-Suppose $\Delta>1$.
-There exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\alpha}^{*}(s))=0$, where $\tilde{\alpha}^{*}(s)$ follows from Lemma 2, and in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
-4.4 Costly Distortions
-In this section, I extend the previous analysis by accounting for the costs that a voter may incur when distorting the issue’s complexity.
-The goal is twofold: first, introducing costs allows us to analyze how beliefs may be distorted, which is valuable in itself.
-Second, it enables us to assess the robustness of the earlier results.
-I begin by considering a fixed cost of distortion, following the approach of other recent papers, such as Levy (2014) and Le Yaouanq (2023).
-I then explore a continuous cost function, where the magnitude of the distortion affects the costs, in Section 4.4.2.
-I will concentrate on conditions under which equilibria with inactive candidates continue to exist.
-4.4.1 Fixed Costs
-First consider that distorting the issue’s complexity implies a fixed cost of $\gamma>0$. That is, each voter now maximizes
-$W(s,\mathbf{p},\tilde{\mu})=AU(s,\mathbf{p},\tilde{\mu})-\gamma\cdot\mathbbm{1}_{\tilde{\mu}\neq\mu}.$
-As before, $W(s,\mathbf{p},\tilde{\mu})$ strictly decreases in $s$, implying the greatest incentive to distort $\mu$ when $s>0$ has a voter receiving $s\rightarrow\infty$, and when $s<0$ a voter receiving $s\approx 0$.
-Of course, if $\gamma$ is too large, no voter has an incentive to choose $\tilde{\mu}\neq\mu$.
-Lemma 3.
-If $\gamma<\gamma^{+}\equiv(1-q)(\beta(1-\kappa)+\Delta-\kappa)$, then there exists $s^{+}>0$ such that
-$\tilde{\mu}^{*}=\left\{\begin{array}[]{cl}0&\text{if }s\geq s^{+}\\ \mu&\text{if }s\in[0,s^{+})\end{array}\right.$
-However, if $\gamma\geq\gamma^{+}$, then $\tilde{\mu}^{*}=\mu$ for all $s\geq 0$. Moreover, if $\gamma<\gamma^{-}\equiv q(\beta+\Delta-(\beta+1)\kappa)$, then there exists $s^{-}<0$ such that
-$\tilde{\mu}^{*}=\left\{\begin{array}[]{cl}\infty&\text{if }s\in[s^{-},0)\\ \mu&\text{if }s<s^{-}\end{array}\right.$
-However, if $\gamma\geq\gamma^{-}$, then $\tilde{\mu}^{*}=\mu$ for all $s<0$.
-Having established voters’ optimal deviations from Bayesian rationality, we next seek to determine under which conditions inactivity against climate change, despite better information, remains an equilibrium.
-First, note that if $\gamma\geq\max\left\{\gamma^{-},\gamma^{+}\right\}$, then no voter distorts their beliefs, and all voters interpret information as perfect Bayesians.
-It follows from Assumption 2 that in this case, information aggregates correctly in an election with $p_{1}\neq p_{2}$, and therefore the conditions for efficient policy platforms in equilibrium are satisfied.
-If $\gamma<\min\left\{\gamma^{-},\gamma^{+}\right\}$, there will be both voters who received good news ($s<0$) and those who received bad news ($s>0$) who choose $\tilde{\mu}^{*}\neq\mu$, while others will interpret signals correctly.
-Proposition 5.
-Suppose $\Delta>1$.
-There exists $\hat{\gamma}>0$ such that iff $\gamma<\hat{\gamma}$, then there exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, where $\tilde{\mu}^{*}$ follows from Lemma 3, and in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
-The proposition shows that our earlier analysis is robust to the introduction of a fixed cost that needs to be paid if a voter chooses $\tilde{\mu}\neq\mu$.
-The result is of course not very surprising given Proposition 3.
-However, one thing to note is that the incentives to distort $\mu$ of voters with $s<0$ do not matter for the result.
-In fact, those voters vote more often informatively with motivated beliefs than without, which is beneficial for choosing the optimal policy.
-4.4.2 Quadratic Costs
-Next consider the case of a smooth cost of distortion function.
-We demand a number of properties our cost function should satisfy.
-It should be convex increasing when we move away from $\mu$ and zero when there is no distortion.
-Moreover, it should not be symmetric in the sense that $\tilde{\mu}=\mu+\epsilon$ and $\tilde{\mu}=\mu-\epsilon$ should not lead to the same cost.
-To see why assume $\mu=1$.
-A symmetric cost function would imply that $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=0}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=2}$.
-But note that $\tilde{\mu}=0$ implies that the issue is perceived as infinitely incomplex, and any signal $s\neq 0$ will lead to updating to either 0 or 1.
-At the same time, $\tilde{\mu}=2$ represents only a gradual increase in complexity.
-If both have the same cost, then downward distortions are generally “cheaper” than upward distortions.
-We therefore demand that the cost of equal relative distortions upwards and downwards are the same.
-This implies the cost function is compressed for downwards distortions.
-More formally, $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\lambda\cdot\mu}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\lambda}$ for all $\lambda>0$ and all $\mu>0$.
-In particular, we assume the following cost function:
-$C(\tilde{\mu},\mu)=\left\{\begin{array}[]{ll}\frac{c}{2}\left(\tilde{\mu}-\mu\right)^{2}&\text{if }\tilde{\mu}\geq\mu\\ \frac{c}{2}\left(\frac{\mu}{\tilde{\mu}}\left(\tilde{\mu}-\mu\right)\right)^{2}&\text{if }\tilde{\mu}<\mu\end{array}\right.$ (6)
-With this cost function we have $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\lambda\cdot\mu}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\lambda}=\frac{c}{2}\mu^{2}\left(\lambda-1\right)^{2}$ for all $\lambda>1$ as desired.
-In the left panel of Figure 1, the cost function is plotted for $\mu=1$ and $c=2$.
-One property of it is that $\lim_{\tilde{\mu}\rightarrow\infty}C(\tilde{\mu},\mu)=\lim_{\tilde{\mu}\rightarrow 0}C(\tilde{\mu},\mu)=\infty$, and therefore we have $\tilde{\mu}\in(0,\infty)$ for any $s$.
-When deciding how to optimally distort $\mu$, each voter aims to maximize
-Figure 1: Left panel: Cost function when $\mu=1$ and $c=2$. Right panel: $\tilde{\mu}$ as a function of $s$ for $\mu=1$ and different values of $c$. Higher $c$ moves $\tilde{\mu}$ closer to $\mu$.
-$W(s,\mathbf{p},\tilde{\mu})=AU(s,\mathbf{p},\tilde{\mu})-C(\tilde{\mu},\mu).$
-As before, focus on $p_{1}\neq p_{2}$.
-When the signal is uninformative, $s=0$, any distortion only causes cost, but does not change $AU(s,\mathbf{p},\tilde{\mu})$.
-Hence, in this case, $\tilde{\mu}^{*}=\mu$.
-But what happens when signals are very informative?
-Suppose $s\rightarrow\infty$.
-Then any distortion except $\tilde{\mu}=0$ will move the belief about $\omega$ to $\pi=1$.
-But the cost of choosing $\tilde{\mu}=0$ is infinite, and hence this cannot be optimal.
-Therefore, it is optimal to not distort at all.
-The same is true for $s\rightarrow-\infty$.
-For any other signal realization, $s\in\left(-\infty,0\right)\cup\left(0,\infty\right)$, the direction of the optimal distortion is the same as in Lemma 1.
-Moreover, $W(s,\mathbf{p},\tilde{\mu})$ is a continuous and smooth function of both $s$ and $\tilde{\mu}$, and hence also $\tilde{\mu}$ is a continuous and smooth function of $s$.
-To get a better intuition of how $\tilde{\mu}^{*}$ changes with $s$, I plotted it for $\mu=1$ and different values of $c$ in the right panel of Figure 1.
-The next lemma states the above intuitions formally:
-Lemma 4.
-Suppose $\Delta>1$.
-If $s\in\{-\infty,0,\infty\}$, then $\tilde{\mu}(s)=\mu$.
-Otherwise, and if $s>0$, then $\tilde{\mu}(s)<\mu$, and there exists $\bar{s}>0$ such that $\tilde{\mu}(s)$ decreases in $s$ if $s\in[0,s^{+})$, and it increases in $s$ if $s>s^{+}$.
-If $s<0$, then $\tilde{\mu}(s)>\mu$, and there exists $s^{-}<0$ such that $\tilde{\mu}(s)$ decreases in $s$ if $s\in(s^{-},0]$, and it increases in $s$ if $s<s^{-}$.
-Moreover, for every $s$, the absolute distortion $\left|\mu-\tilde{\mu}\right|$ decreases in $c$
-Now let us discuss the implications of the lemma for voters’ decisions at the ballot, and by extension, for candidates’ incentives to offer the optimal policy.
-First, just as in the case of fixed costs, if $s<0$, voters tend to amplify the signals, leading more voters to vote informatively compared to the absence of motivated beliefs.
-This improves the ability to select the right policy through an election.
-Therefore, to understand the conditions under which candidates ignore information and choose policy 0 regardless of $\omega$, we need to focus on voters receiving $s>0$.
-These voters interpret signals as less informative than they truly are, resulting in fewer voters casting informative votes compared to the absence of motivated beliefs.
-Because $q<\tilde{\pi}$, if either the distortion of $\mu$ is significant or the signal is too weak, such a voter may cast a ballot for policy 0 despite $s>0$.
-Intuitively, as $c$ increases, voters choose to distort $\mu$ less and less.
-Consequently, $\lim_{c\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$.
-In this case, by Assumption 2, a majority will always cast a ballot for $p=\omega$, giving candidates an incentive to heed their information.
-However, if $c$ is small, we revert to the result from Proposition 3.
-By continuity, there exists a $\hat{c}>0$ such that if $c\in[0,\hat{c}]$, motivated beliefs will lead to distortions and voting behavior that prevent candidates from following their information.
-Proposition 6.
-Suppose $\Delta>1$.
-There exists $\hat{c}>0$ such that iff $c<\hat{c}$, then there exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, where $\tilde{\mu}^{*}$ follows from Lemma 4, in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
-4.4.3 Comparing Fixed and Continuous Costs
-The analysis in this section has given us more nuanced results for the optimal distortion of beliefs.
-Moreover, in both extensions discussed so far, the inactive equilibrium prevails as the unique equilibrium if the cost of distorting informativeness of signals is low enough.
-However, the two models discussed differ significantly in how voters choose to distort information and hence beliefs, see Figure 2, allowing to derive testable predictions.
-Figure 2: Equilibrium belief distortion (solid, red curve) with fixed costs (left panel) and with a strictly convex cost function (right panel). The blue and dashed curve is the Bayesian belief.
-With fixed costs, only the worst signals are distorted, both in the case of good news ($s<0$) and bad news ($s>0$).
-When news are bad, the worst signals are the ones suggesting $\omega=1$ the strongest.
-This leads to low anticipatory utility and thus the signal is interpreted as not informative.
-However, if bad news is weak, then the belief does not change a lot anyway, and hence choosing $\tilde{\mu}\neq\mu$ at a cost is not attractive.
-When news is good, the worst signals are the ones suggesting $\omega=0$ the least.
-Weak good news leads to little updating of beliefs, and overreacting to information is better for anticipatory utility, even if it is somewhat costly.
-However, if good news is strong enough, then overreacting to the signal does not change beliefs that much anymore, and hence distorting beliefs is not in a voters interest.
-If, instead, we consider a continuous and convex cost function, the results are to some degree reversed.
-In fact, when $s$ is very large, implying a voter receives strong bad news, then interpreting the signal as not informative is very costly, and hence a voters will be better served simply interpreting the signal in an (almost) Bayesian fashion.
-The same is true when $s$ is close to zero.
-However, bad news of intermediate strength will be downplayed significantly.
-When a voter receives good news we find similar results.
-If $s$ is very low, the voter’s belief about $\omega$ is close to zero even if $\tilde{\mu}=\mu$, and hence distorting $\mu$ is not very attractive.
-The same is true if good news is very weak.
-However, if good news is of intermediate strength, then distorting $\mu$ upwards is attractive.
-5 Discussion and Some Evidence
-The formal analysis highlights at least two important determinants of the efficiency of policy-making in the face of climate change, rhetoric and trust in government.
-The Importance of Rhetoric:
-When $\Delta>1$, there is no equilibrium in which office-motivated candidates campaign with the policies needed for severe climate change, even if they know for sure that $\omega=1$.
-Of course, in real life, we do not discuss the issue of climate change in terms of baseline welfare losses, but rather using verbal descriptions.
-In recent years, an increasing number of people have stopped using the relatively neutral terms “global warming” or “climate change” and have instead adopted more extreme and loaded language such as “climate crisis” or “climate emergency.” Indeed, a comparison of Google search trends in the U.S.
-from January 2014 to July 2024 shows that, starting around 2019, the relative frequency with which these terms are used has been increasing, see Figure 3.
-One reason for this change is that many activist groups started to use a more aggressive rhetoric in an effort to emphasize the importance of swift climate policy changes.
-For example, the activist group Extinction Rebellion writes on their homepage: [^3] See https://rebellion.global/ (accessed on 24/09/2024).
-(a) Climate Change
-(b) Climate Emergency
-(c) Climate Crisis
-Figure 3: Five months moving averages of Google search trends in the U.S. for “Climate Change”, “Climate Emergency”, and “Climate Crisis” from January 2014 to July 2024.
-This is an Emergency.
-Life on Earth is in crisis.
-Our climate is changing faster than scientists predicted and the stakes are high.
-Biodiversity loss.
-Crop failure.
-Social and ecological collapse.
-Mass extinction.
-We are running out of time, and our governments have failed to act.
-The model suggests that such a shift in language may have undesired consequences, because it may lead voters to ignore information suggesting that climate change is a problem, whereas information stating the opposite is trusted more than it should.
-And in fact, the fraction of U.S.
-Americans who stated in PEW polls that they “personally care a great deal about the issue of climate change” increased between 2016 and 2018 from 36% to 44%, but then dropped to 37% in 2023.
-At the same time, the fraction of people who stated to either not care too much or to not care at all about the issue first decreased from 26% in 2016 to 22% in 2018, but then increased again to 27% in 2023.444See https://www.pewresearch.org/science/2023/10/25/how-americans-view-future-harms-from-climate-change-in-their-community-and-around-the-u-s/.
-These changes can be explained by the model, and they create stronger incentives for politicians to ignore their information about the true state of the world when designing policies.
-That framing and language may have an important effect on wether people learn or not is well established in psychology.
-For example, Smith and Mackie (2007) mention the importance of not creating too much anxiety in receivers when the intent is to change their attitudes:
-To be successful, fear appeals must arouse just the right amount of anxiety by showing the negative consequences […] that will follow if behavior doesn’t change.
-The provision of an explicit avenue of action […] eliminates the anxiety by showing how the negative consequences can be avoided.
-Smith and Mackie (2007), page 257
-Excessively negative rhetoric may create too much anxiety and may hence prevent attitude changes and learning, see also McDonald, Chai, and Newell (2015).
-The risk to trigger such excessive emotional responses is particularly large when people’s general livelihood seems threatened, which is true in the case of climate change.
-Writing explicitly about the issue climate change, Clayton, Manning, Krygsman, and Speiser (2017) state that
-[t]he ability to process information and make decisions without being disabled by extreme emotional responses is threatened by climate change.
-Some emotional response is normal, and even negative emotions are a necessary part of a fulfilling life.
-In the extreme case, however, they can interfere with our ability to think rationally, plan our behavior, and consider alternative actions.
-Clayton, Manning, Krygsman, and Speiser (2017), page 16
-Other papers that highlight the importance of framing for the efficacy to learn about climate change include Spence and Pidgeon (2010), Jones and Peterson (2017), and Ngo, Poortvliet, and Klerkx (2022).
-Hence, language or rhetoric may indeed be important factors in explaining voters’ attitudes towards climate change, and therefore may as well be determinants of political inactivity in the face of climate change.
-The Importance of Trust in Government:
-On the other hand, Trust in the Government to choose appropriate policies for the problems at hand matters.
-In the model this is captured by $\kappa(s,\mathbf{p},\tilde{\mu})$.
-If voters belief the correct policy will be chosen, then $\kappa(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\tilde{\mu})$, and in this case an efficient equilibrium exists if $\Delta<1$.
-However, if voters are pessimistic about politician’s ability or incentives to implement the right policies—$\kappa(s,\mathbf{p},\tilde{\mu})$ is then independent of the signal $s$—then this becomes self-confirming, leading politicians to ignore their information, and the optimal policy may not be chosen.
-Hence, the model suggests that there is a positive relationship between efforts against climate change and trust in government.
-Note that in a model with no motivated beliefs, such a prediction could not emerge, because the belief about climate change is independent of trust in government, and hence so is the implemented policy.
-It is the interaction of trust in government $\kappa$ and motivated beliefs $\tilde{\mu}$ that causes this result.
-Of course, trust in government is not the only determinant of progress against climate change.
-Different countries face different costs of adopting new policies.
-In the model, this is captured by $\beta$.
-When $\beta$ is large, inaction may be very costly, while if $\beta$ is small, inaction has little consequences.
-Empirically, $\beta$ could be interpreted as economic cost of (not) dealing with climate change in an appropriate way.
-Often, anti-climate change policies aim at reducing CO2 emissions.
-This may be quite costly for countries rich in oil and gas, while it is less costly for other countries.
-Hence, countries rich in fossil energy sources are less likely to make progress against climate change than countries with little or no such resources.
-Statistic N Mean St. Dev. Min Max
-CCPI 19 53.66 13.59 26.47 79.61
-TGI 19 51.33 13.07 25.63 77.54
-Oil 19 9.95 38.82 0.00 170.00
-Risk 19 2.20 1.15 1.00 5.40
-Table 1: Summary statistics of included variables.
-We can test the hypotheses using cross-sectional data.
-As the dependant variable, I use the Climate Change Performance Index (CCPI) published jointly by Germanwatch, the NewClimate Institute, and the Climate Action Network International.
-[^5] See https://ccpi.org/.
-The CCPI takes values between 0 and 100, where higher values indicate that a country performs better in enacting appropriate policies against climate change.
-The considered policies include reductions in greenhouse gas emissions (40%), expansion of renewable energy sources (20%), general energy use (20%), and other climate policies (20%).
-As controls for trust in government I use the OECD’s Trust in Government Indicator (TGI).
-[^6] See https://www.oecd.org/en/data/indicators/trust-in-government.html The TGI also takes values between 0 and 100 and equals the share of people reporting to have confidence in the national government.
-As additional control variables, I use countries’ confirmed oil reserves in billion barrels (Oil), which can be interpreted as a measure of the cost of climate change policies, and the Global Climate Risk Index (Risk), published by Germanwatch, which can be interpreted as a measure of the benefits of climate change policies.
-[^7] See https://www.germanwatch.org/en/cri In Table 1 I summarize the included variables.
-Figure 4: Scatterplot of Climate Change Performance Index (CCPI) in 2023 and the OECD Trust in Government Indicator (TGI) from 2022 with OLS trendline.
-Unfortunately, the number of countries for which all of these variables are available is small ($N=19$), and hence power is limited.
-Nevertheless, let us now take a first look at the data.
-Figure 4 shows a scatter plot of CCPI and TGI for the different countries including a trend line.
-The plot indicates a positive relation between TGI and CCPI, as predicted by the model.
-The Scandinavian countries have the highest trust in government and also show some of the highest CCPI scores.
-The US, Poland, or Colombia have relatively low trust in government, and they also have low CCPI scores.
-Canada has a low CCPI score despite having a medium-high TGI score.
-[^8] Note that the theoretical analysis is valid only for democracies.
-However, because all countries in the sample are classified either as Democracy or Full Democracy in the Polity IV data series, and because the sample size is very small, I do not control for polity scores in the regressions.
-However, Canada has significant confirmed Oil reserves.
-To shed further light on the empirical relation between trust in government and climate change policies, I regressed CCPI on TGI and a subset of the other control variables.
-In particular, I estimate the following regression equation:
-$CCPI_{i}=\alpha+\beta\,TGI_{i}+\gamma\,\mathbf{X}_{i}+\epsilon_{i}$
-The estimate of $\beta$ captures the relationship we are interested in.
-Table 2 below shows the regression results.
-Despite the low power, $\hat{\beta}$ is positive and statistically significant in all specifications.
-Moreover, a one standard deviation increase of TGI leads to an increase of CCPI between 0.463 and 0.495 standard deviations, depending on the chosen specification.
-Therefore, the relation appears to be of relevance.
-Checking for potential heteroscedasticity, the $p$-values of the Breusch-Pagan test are larger than 0.4 for all specifications, suggesting that heteroscedasticity is no problem.
-Inspecting the Q-Q plots of the different specifications (see Figure 5 in Appendix B), further supports this conclusion.
-Hence, there indeed seems to be a significant and positive correlation between CCPI and TGI, as suggested by the model.
-Dependent variable:
-CCPI
-(1) (2) (3) (4)
-TGI $0.484^{**}$ $0.481^{**}$ $0.498^{*}$ $0.515^{**}$
-$(0.223)$ $(0.193)$ $(0.250)$ $(0.216)$
-Oil $-0.168^{**}$ $-0.170^{**}$
-$(0.065)$ $(0.067)$
-Risk2022 $0.397$ $0.982$
-$(2.832)$ $(2.458)$
-Constant $28.825^{**}$ $30.629^{***}$ $27.249$ $26.754^{*}$
-$(11.803)$ $(10.251)$ $(16.566)$ $(14.316)$
-Observations 19 19 19 19
-R2 0.217 0.446 0.218 0.452
-Adjusted R2 0.170 0.377 0.120 0.343
-F Statistic 4.699∗∗ (df = 1; 17) 6.451∗∗∗ (df = 2; 16) 2.224 (df = 2; 16) 4.128∗∗ (df = 3; 15)
-Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
-Table 2: Regression Results.
+This logic is closely related to the one in Calvert, (1985), who shows that the median voter theorem may continue to hold even in the presence of policy motivations.
 6 Conclusion
@@
 The analysis reveals that—in the presence of motivated reasoning—trust in the functioning of government plays a critical role for the efficiency of policy-making.
-A first look at the data confirms this result.
 Moreover, the paper suggests that political rhetoric plays an important role as well: if politicians, scientists, journalists, and activists aim to contribute to better climate change policy, they must navigate a delicate balance between conveying information truthfully and avoiding the risk of pushing voters into denial.
@@
 The proof follows from (4):
-$\begin{array}[]{rcl}\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}&>&0\\ \Leftrightarrow-\dfrac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}&>&0\\ \Leftrightarrow s\left[\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)\right]&<&0\end{array}$
-Here we have $s$ and the expression in brackets that jointly determine the derivative’s sign. Consider next the term in parentheses:
-$\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)>0\Leftrightarrow\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}\equiv\frac{\Delta+\beta}{1+\beta}$
-If $\Delta>1$, then $\tilde{\kappa}>1$, and hence we cannot have $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$.
-It follows that in this case $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$.
-Hence, $\tilde{\mu}^{*}=\infty$ if $s>0$ and $\tilde{\mu}^{*}=0$ if $s<0$.
+$\begin{array}[]{rcl}\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}>0\Leftrightarrow-\dfrac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[b+\Delta-\kappa(1+b)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}&>&0\\ \Leftrightarrow s\left[b+\Delta-\kappa(1+b)\right]&<&0\end{array}$
+The realization of $s$ and the expression in brackets jointly determine the derivative’s sign. Consider next the term in parentheses:
+$b+\Delta-\kappa(1+b)>0\Leftrightarrow\kappa<\tilde{\kappa}\equiv\frac{\Delta+b}{1+b}$
+If $\Delta>1$, then $\tilde{\kappa}>1$, and hence we must have $\kappa<\tilde{\kappa}$.
+It follows that $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$.
+Hence, $\tilde{\mu}^{*}$$\infty$ if $s<0$ and $\tilde{\mu}^{*}=0$ if $s>0$.
 If $s=0$, any $\tilde{\mu}$ yields the same $AU$ and hence $\tilde{\mu}^{*}=\sigma$.
-If $\Delta<1$, $\tilde{\kappa}<1$ as well.
-Hence, if $\kappa(s,\mathbf{p},\tilde{\mu})<\tilde{\kappa}$, then again $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$.
-It follows that $\tilde{\mu}^{*}=\infty$ if $s>0$, $\tilde{\mu}^{*}=0$ if $s<0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$.
-To the contrary, if $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$.
-Hence, $\tilde{\mu}^{*}=0$ if $s>0$, $\tilde{\mu}^{*}=\infty$ if $s<0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$.
-If $\kappa(s,\mathbf{p},\tilde{\mu})=\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=0$, and any $\tilde{\mu}$ yields the same anticipatory utility.
-Thus, $\tilde{\mu}^{*}=\sigma$.
-Finally, if $\Delta=1$, then we cannot have $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$.
-Thus, $AU(s,\mathbf{p},\tilde{\mu})$ is either flat in $\tilde{\mu}$ if $\kappa(s,\mathbf{p},\tilde{\mu})=1$, implying any $\tilde{\mu}$ yields the same anticipatory utility, and we have $\tilde{\mu}^{*}=\sigma$, or the situation resembles the case with $\Delta>1$ discussed above.∎
+If $\Delta<1$, then $\tilde{\kappa}<1$ as well.
+Hence, if $\kappa<\tilde{\kappa}$, then again $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$.
+Thus, $\tilde{\mu}^{*}$$\infty$ if $s<0$, $\tilde{\mu}^{*}=0$ if $s>0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$.
+To the contrary, if $\kappa>\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$.
+Hence, $\tilde{\mu}^{*}=0$ if $s<0$, $\tilde{\mu}^{*}$$\infty$ if $s>0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$.
+If $\kappa=\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=0$, and any $\tilde{\mu}$ yields the same anticipatory utility.
+Thus, $\tilde{\mu}^{*}=\mu$.
+[…]
 A.2 Proof of Corollary 1
@@
 $\Delta>1$:
-We know from Corollary 1 that all voters with $s>0$ have belief $\pi(s)=\hat{\pi}$, while those with $s<0$ have belief $\pi(s)=0$.
+We know from Corollary 1 that all voters with $s>0$ have belief $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$, while those with $s<0$ have belief $\pi(s,\mathbf{p},\tilde{\mu})=0$.
 Hence, the latter always vote for policy 0.
-Wether a voter with $s>0$ votes for policy 1 or policy 0 depends on the comparison of $\tilde{\pi}$ and $\hat{\pi}$.
-In particular, we know from (5) that a voter votes for policy 1 if $\pi(s)>\tilde{\pi}$, for policy 0 if $\pi(s)<\tilde{\pi}$, and they choose each with a probability of 50% if $\pi(s)=\tilde{\pi}$.
+whether a voter with $s>0$ votes for policy 1 or policy 0 depends on the comparison of $\tilde{\pi}$ and $\hat{\pi}$.
+In particular, we know from (5) that a voter votes for policy 1 if $\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}$, for policy 0 if $\pi(s,\mathbf{p},\tilde{\mu})<\tilde{\pi}$, and they choose each with a probability of 50% if $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$.
 Hence, if $\hat{\pi}<\tilde{\pi}$, all voters vote for policy 0.
@@
 As a consequence, policy 0 wins all the time if $\hat{\pi}\leq\tilde{\pi}$.
-This implies that we must have $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ for all $s$.
+This implies that we must have $\kappa^{*}=0$ for all $s$.
 However, if $\hat{\pi}>\tilde{\pi}$, all voters vote informatively, and thus the optimal policy always wins.
@@
 Hence, voters know that the policy that matches the state wins all the time.
-The belief about the state is $\pi(s)=\hat{\pi}$ if $s\geq 0$, implying $\left.\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s\geq 0}=\hat{\pi}$, and $\pi(s)=0$ if $s<0$, and hence $\left.\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s<0}=0$.
-Could there be other pure strategy equilibria?
-$\tilde{\mu}^{*}$ is uniquely determined, and hence also $\pi(s,\mathbf{p},\tilde{\mu}^{*})$ is uniquely determined.
-But this implies that equilibrium vote shares are unique, leaving no room for beliefs $\kappa(s,\mathbf{p},\tilde{\mu})$ that are consistent with these vote shares and that differ from the ones established before.
-Hence, if $\Delta>1$, no other equilibrium in the voting subgame can exist.
-$\Delta\leq 1$:
-First assume that $\kappa(s,\mathbf{p},\tilde{\mu})$ is increasing in $s$ in the sense that $\text{Sign}\left[\kappa(s,\mathbf{p},\tilde{\mu})-\tilde{\kappa}\right]=\text{Sign}\left[s\right]$.
-Then any voter with $s<0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0.
-Moreover, any voter with $s>0$ chooses also $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1.
-This implies that indeed the policy that matches the state always wins and we need to have $\left.\kappa(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s<0}=0$ as well as $\left.\kappa(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s>0}=1$, proving part 2 (a) of the proposition.
+The belief about the state is $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$ if $s\geq 0$, implying $\left.\kappa^{*}\right|_{s\geq 0}=\hat{\pi}$, and $\pi(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$, and hence $\left.\kappa^{*}\right|_{s<0}=0$.
+[…]
+$\Delta=1$:
+If $\kappa=1$, then $\pi(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\mu)$.
+Hence, any voter with $s<0$ will always vote for policy 0.
+This means policy 0 wins at least if $\omega=0$, contradicting $\kappa=1$.
+Hence, if $\Delta=1$, there cannot be an equilibrium with $\kappa=1$.
+If $\kappa<1=\tilde{\kappa}$, then $\tilde{\mu}^{*}=0$ if $s>0$ and $\tilde{\mu}^{*}$$\infty$ if $s<0$. It follows that equilibrium is the same as in the case of $\Delta>1$.
+$\Delta<1$:
+First assume that $\kappa$ positively correlates with $s$ in the sense that $\text{Sign}\left[\kappa-\tilde{\kappa}\right]=\text{Sign}\left[s\right]$.
+Any voter with $s<0$ chooses $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0.
+Moreover, any voter with $s>0$ chooses also $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1.
+This implies that indeed the policy that matches the state always wins and we need to have $\left.\kappa\right|_{s<0}=0$ as well as $\left.\kappa\right|_{s>0}=1$, proving part 2 (a) of the proposition.
 Next, assume voters generally believe that policy 0 will be implemented, independent of the true state $\omega$.
-Then $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=0$ for all $s$.
-Then any voter with $s<0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0.
-Moreover, any voter with $s>0$ chooses $\tilde{\mu}^{*}=0$.
-If and only if $\hat{\pi}\leq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 0, and hence when $\hat{\pi}\leq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 0 is always chosen.
+Then $\kappa=0$ for all $s$.
+It follows that any voter with $s<0$ chooses $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0.
+Moreover, any voter with $s>0$ chooses $\tilde{\mu}^{*}=0$, implying $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\hat{\pi}$.
+Iff $\hat{\pi}\leq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 0, and hence when $\hat{\pi}\leq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 0 is always chosen.
 This proves part 2 (b).
 Now assume voters generally believe that policy 1 will be implemented, independent of the true state $\omega$.
-Then $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=1$ for all $s$, and any voter with $s>0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1.
-Moreover, any voter with $s<0$ chooses $\tilde{\mu}^{*}=0$.
+Then $\kappa=1$ for all $s$, and any voter with $s>0$ chooses $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1.
+Moreover, any voter with $s<0$ chooses $\tilde{\mu}^{*}=0$, inducing $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\hat{\pi}$.
 If and only if $\hat{\pi}\geq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 1, and hence when $\hat{\pi}\geq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 1 is always chosen.
 This proves part 2 (c).
-Consider a situation in which beliefs are just as in the equilibrium discussed above for the case of $\Delta>1$.
-That is, $\kappa(s,\mathbf{p},\tilde{\mu})=0$ for all $s$.
-Then, as before, $\tilde{\mu}^{*}(s)=\infty$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ if $s<0$, and $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=q$ else.
-It follows that if $q\leq\frac{1}{1+\beta}$, a majority of voters always votes for policy 0.
-Hence, this is indeed an equilibrium.
-Consider next a situation in which beliefs are just opposite to what we have seen above, namely $\kappa(s,\mathbf{p},\tilde{\mu})=1$ for all $s$.
-Then, $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ if $s>0$, and $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=q$ else.
-It follows that if $q\geq\frac{1}{1+\beta}$, a majority of voters always votes for policy 1.
-Hence, this is also an equilibrium.
 Finally, consider part 2 (d).
 An equilibrium in which the optimal policy never wins would imply policy 1 wins if $\omega=0$ and policy 0 wins if $\omega=1$.
-Hence, we would need to have $\kappa(s,\mathbf{p},\tilde{\mu})$ weakly decreasing in $s$.
+Hence, we would need to have $\kappa$ weakly decreasing in $s$.
 Consider a voter with signal $s>0$.
@@
 Hence, it cannot be true that $\kappa\geq\tilde{\kappa}$.
-If $\kappa(s,\mathbf{p},\tilde{\mu})<\tilde{\kappa}$, these voters all hold belief $\hat{\pi}$.
+If $\kappa<\tilde{\kappa}$, these voters all hold belief $\hat{\pi}$.
 Next consider voters with $s<0$.
-It must be true that most of these voters vote for policy 1, implying $\kappa(s,\mathbf{p},\tilde{\mu})\leq\tilde{\kappa}$ is not possible.
-Thus, because $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$, these voters all also hold belief $\hat{\pi}$.
+It must be true that most of these voters vote for policy 1, implying $\kappa\leq\tilde{\kappa}$ is not possible.
+Thus, because $\kappa>\tilde{\kappa}$, these voters all also hold belief $\hat{\pi}$.
 If $\hat{\pi}>\tilde{\pi}$, policy 1 always wins, contradicting that the optimal policy never wins.
@@
 Finally, if $\hat{\pi}=\tilde{\pi}$, in each state, each policy wins with 50%, and thus also the optimal policy wins with a chance of 50%.
-This implies that the actual state is irrelevant for the probability of each policy winning, and thus we must have $\kappa(s,\mathbf{p},\tilde{\mu})=\frac{1}{2}$ for all $s$.
-But then it is not possible that $\kappa(s,\mathbf{p},\tilde{\mu})$ differs for voters with $s>0$ and voters with $s<0$.
+This implies that the actual state is irrelevant for the probability of each policy winning, and thus we must have $\kappa=\frac{1}{2}$ for all $s$.
+But then it is not possible that $\kappa$ differs for voters with $s>0$ and voters with $s<0$.
 Hence, such an equilibrium of the voting subgame cannot exist.
-This complete the proof of the proposition.∎
+This completes the proof of the proposition.∎
 A.4 Proof of Proposition 2
@@
 A.5 Proof of Proposition 3
-We first consider the case of severe climate change being catastrophic, $\Delta>1$.
-The other case of severe climate change having only mild baseline welfare consequences will be considered thereafter.
-Note that $\hat{\pi}=q\leq\tilde{\pi}$ in a symmetric pure strategy equilibrium.
-Suppose that off the equilibrium path voters observe $\tilde{\mathbf{p}}\in\{\{0,1\},\{1,0\}\}$.
-$\Delta>1$:
-We know from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=\hat{\pi}$ if $s\geq 0$.
-Hence, $\hat{\pi}\leq\tilde{\pi}$, and thus no voter ever votes for policy 1, whereas policy 0 always wins at the ballot.
-It follows that $\kappa^{*}(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$, with $\tilde{\mu}^{*}(s)$ as described in Lemma 1.
-Thus, the unique equilibrium is candidates choosing $\mathbf{p}^{*}=(0,0)$.
-This is supported by $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$ and $\tilde{\mu}^{*}(s)$ as described in Lemma 1.
-$\Delta\leq 1$:
-Assume first $\kappa(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ for all $s$.
-Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=\hat{\pi}=q$ if $s>0$.
-Because $q\leq\frac{1}{1+\beta}$, a majority of voters always votes for policy 0.
-It follows from Proposition 2 that no candidate has an incentive to campaign on $p=1$.
-Consequently, $\mathbf{p}^{*}=(0,0)$, $\kappa(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, and $\tilde{\mu}^{*}$ follows from Lemma 1.
-Suppose next that $\Delta<1$ and that $\kappa>\tilde{\kappa}$ if $s>0$ and $\kappa<\tilde{\kappa}$ if $s<0$.
-Note that this means that we must have $\Delta<1$.
-Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=1$ if $s>0$.
-Hence, all voters all the time vote informatively, and thus the policy that matches the state wins, implying $\kappa=\pi(s,\tilde{\mathbf{p}},\tilde{\mu})$.
-It follows from Proposition 2 that, if candidates expect voters to behave this way, they have an incentive to choose the socially optimal policies.
-Hence, $\mathbf{p}^{*}=(\omega,\omega)$, $\kappa(s,\mathbf{p}^{*},\tilde{\mu}^{*})=\pi(s,\mathbf{p}^{*},\tilde{\mu}^{*})$, and $\tilde{\mu}^{*}$ follows from Lemma 1.
-Now suppose that $\Delta=1$.
-Moreover, suppose an efficient equilibrium exists also in this case, with $\kappa(s,\mathbf{p},\tilde{\mu})=1=\tilde{\kappa}$ if $s>0$, and $\kappa=0<\tilde{\kappa}$ if $s<0$.
-Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$, and thus any voter with $s<0$ votes for policy 0.
-For voters with $s>0$, $\tilde{\mu}^{*}=\mu$, and thus these voters hold perfect Bayesian beliefs.
-This implies, by Assumption 2, that voters vote informatively.
-However, then the right policy is chosen in any state, and thus we need to have $\kappa(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\tilde{\mu})$.
-But $\pi(s,\mathbf{p},\tilde{\mu})<1$, and therefore voters with $s>0$ would choose $\tilde{\mu}=0$ instead of $\tilde{\mu}=\mu$.
-Hence, when $\Delta=1$, no efficient equilibrium exists.
+In a symmetric equilibrium, each candidate wins with a 50% probability.
+We next show that no lucrative deviations exist.
+Hence, suppose that off the equilibrium path voters observe $\tilde{\mathbf{p}}\in\{(0,1),(1,0)\}$.
+$\Delta\geq 1$:
+We know from Corollary 1 that after $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=\hat{\pi}$ if $s\geq 0$.
+Hence, $\hat{\pi}\leq\tilde{\pi}$, and thus no voter ever votes for policy 1.
+It follows that policy 1 never wins.
+Consequently, a deviation from $\mathbf{p}=(0,0)$ is never profitable, and thus $\mathbf{p}^{*}=(0,0)$ is indeed an equilibrium.
+Moreover, $(1,1)$ cannot be an equilibrium, because a deviation is always profitable, because when spotting such a deviation, voters will ignore information implying that climate change is severe.
+Thus, the unique symmetric equilibrium is candidates choosing $\mathbf{p}^{*}=(0,0)$.
+This is supported by $\kappa^{*}=0$ and $\tilde{\mu}^{*}(s)$ as described in Lemma 1.
+$\Delta<1$:
+It is easy to show that the inefficient equilibrium with $\mathbf{p}^{*}=(0,0)$ still exists. Hence, here we focus on the existence of the efficient equilibrium.
+Suppose candidates choose $\mathbf{p}=(1,1)$.
+Hence, as before, each wins with a 50% probability.
+If a candidate deviates, they face conflicting policy platforms, and hence do not learn from $\mathbf{p}$ itself.
+The only difference between voters is the signals they receive.
+Hence, if they hold different beliefs, these need to be based on signals.
+[…]
 It follows from Proposition 1 part 2 (d) that no equilibrium can exist in which always the wrong policy is chosen.
 Hence, we now only need to show that no equilibrium, in which policy 1 always wins, can exist.
-Suppose to the contrary that voters believe that policy 1 wins with certainty, $\kappa(s,\mathbf{p},\tilde{\mu})=1\geq\tilde{\kappa}$ for all $s$.
-Then it follows from Corollary 1 that, if $\Delta<1$, $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=q\leq\tilde{\pi}$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=1$ if $s>0$.
-Therefore, all voters with $s>0$ always vote for policy 1.
-If sufficiently many voters with $s<0$ also vote for policy 1, policy 1 wins all the time.
-This is the case iff $q\geq\frac{1}{1+\beta}$.
-However, this contradicts Assumption 1.
-A similar argument shows that $\mathbf{p}=(1,1)$ cannot be equilibrium, either, if $\Delta=1$.
+Suppose to the contrary that voters expect policy 1 to be implemented with certainty, $\kappa=1\geq\tilde{\kappa}$ for all $s$.
+Then it follows from Corollary 1 that $\pi(s,\mathbf{p},\tilde{\mu})=q<\tilde{\pi}$ if $s<0$ and $\pi(s,\mathbf{p},\tilde{\mu})=1$ if $s>0$.
+Therefore, all voters vote informatively, implying $p=\omega$ wins if $p_{1}\neq p_{2}$, thus contradicting $\kappa=1$.
 Hence, it follows that no equilibrium, in which policy 1 is always chosen, exists.
-This proves the proposition.
-∎
-A.6 Proof of Lemma 2
-Take the belief about $\omega$ of a voter who receives $s$ and chooses $\tilde{\alpha}$:
-$\pi(s,\tilde{\alpha})=\frac{q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)}{q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)+(1-q)\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right)}.$
-Anticipatory utility is defined as before in (2). Hence, it’s derivative with respect to $\tilde{\alpha}$ is
-$\dfrac{\partial AU}{\partial\tilde{\alpha}}=\frac{(1-q)qg(s)\left[\beta+\Delta-(\beta+1)\kappa\right](f_{1}(s)-f_{0}(s))}{\left[q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)+(1-q)\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right)\right]^{2}}$
-The derivative is positive if $f_{1}(s)-f_{0}(s)>0\Leftrightarrow s>0$, negative if $f_{1}(s)-f_{0}(s)<0\Leftrightarrow s<0$, and zero else. This proves the result. ∎
-A.7 Proof of Proposition 4
-It follows from Lemma 2 that independent of the true state $\omega$, a majority of voters prefers to vote for policy 0 when offered the choice. The proof then follows from Proposition 2. ∎
-A.8 Proof of Lemma 3
-We focus on pure strategy equilibria.
-It then follows from Lemma 1 that when $s>0$, a voter would either like to choose $\tilde{\mu}=\infty$, or $\tilde{\mu}=\mu$.
-Moreover, when $s<0$, a voter would either like to choose $\tilde{\mu}=0$, or $\tilde{\mu}=\mu$.
-Next have a look how a voter’s objective function changes with $s$:
-$\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial s}=-\frac{2\mu(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\mu s}{\tilde{\mu}^{2}}}}{\left(q\tilde{\mu}\left(e^{\frac{2\mu s}{\tilde{\mu}^{2}}}-1\right)+\tilde{\mu}\right)^{2}}<0,$ (7)
-and hence $W(s,\mathbf{p},\tilde{\mu})$ strictly decreases in $s$ for all $\tilde{\mu}$.
-Note that when $\tilde{\mu}=0$, then the realization of $s$ is irrelevant as the belief is $\left.\pi(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}=q$.
-Hence, the voter with the greatest incentive to distort $\mu$ has $s\rightarrow\infty$.
-For this voter, the payoff of not distorting and instead choosing $\tilde{\mu}=\mu$ is
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu\wedge s\rightarrow\infty}=-\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))-\Delta.$
-If, to the contrary, any such voter chooses $\tilde{\mu}=0$, the payoff is
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}=-\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta)-\gamma.$
-If $\gamma\geq\gamma^{+}\equiv(1-q)(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta-\kappa(s,\mathbf{p},\tilde{\mu}))$, then no voter with $s>0$ has an incentive to distort the issue’s complexity.
-Otherwise, that is, if $\gamma<\gamma^{+}$, then there exists a unique signal realization $s=s^{+}>0$ such that a voter receiving this signal is indifferent between choosing $\tilde{\mu}=0$ and $\tilde{\mu}=\mu$.
-The payoff of a voter choosing to not distort a signal $s$ is
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=-\frac{\kappa(s,\mathbf{p},\tilde{\mu})(1-q)+q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta)e^{\frac{2\mu s}{\sigma^{2}}}}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}.$
-To find the indifferent voter, we need to solve
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}\Leftrightarrow s=s^{+}\equiv\frac{\sigma^{2}\ln\left(1+\frac{\gamma}{q(\beta-\gamma+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta+\Delta))}\right)}{2\mu}.$
-Next consider a voter with $s<0$.
-It follows from Lemma 1 that such a voter would either like to choose $\tilde{\mu}=\infty$ or $\tilde{\mu}^{*}=\mu$.
-Moreover, by (7), when $\tilde{\mu}=\mu$, utility is the lowest when $s=0$.
-Hence, if a voter receiving $s=0$ has no incentive to distort $\mu$, then no voter with $s<0$ has an incentive to do so.
-The payoff of not distorting is
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu\wedge s=0}=-\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta).$
-If, however, a voter with $s<0$ chooses to distort beliefs, he receives
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\infty\wedge s<0}=-\kappa(s,\mathbf{p},\tilde{\mu})-\gamma.$
-Hence, if $\gamma\geq\gamma^{-}\equiv q(\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu}))$, then no voter with $s<0$ chooses $\tilde{\mu}\neq\mu$.
-However, if $\gamma<\gamma^{-}$, then there exists a voter with $s=s^{-}<0$ who is indifferent between distorting and not distorting.
-To find the indifferent voter, we need to solve
-$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\infty\wedge s<0}\Leftrightarrow s=s^{-}\equiv\frac{\sigma^{2}\ln\left(\frac{\gamma(q-1)}{q(\beta(\kappa(s,\mathbf{p},\tilde{\mu})-1)+\gamma-\Delta+\kappa(s,\mathbf{p},\tilde{\mu}))}\right)}{2\mu}.$
-Any voter with $s<s^{-}$ chooses $\tilde{\mu}=\mu$, while any voter with $s\in(s^{-},0)$ chooses $\tilde{\mu}=0$.∎
-A.9 Proof of Proposition 5
-In an inactivity equilibrium, we have $\kappa(s,\mathbf{p},\tilde{\mu})=0$.
-Then, $\gamma_{0}^{+}\equiv\left.\gamma^{+}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=(1-q)(\beta+\Delta)$ and $\gamma_{0}^{-}\equiv\left.\gamma^{-}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=q(\beta+\Delta)$.
-Note that when $q>\frac{1}{2}$, then $\gamma_{0}^{-}>\gamma_{0}^{+}$.
-Hence, if severe climate change is more likely a priori, then there are more distortions of good news than of bad news, and vice versa.
-Moreover, $\left.s^{+}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=\frac{\sigma^{2}\ln\left(1+\frac{\gamma}{q((\Delta+\beta)(1-q)-\gamma)}\right)}{2\mu}$ and $\left.s^{-}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=\frac{\sigma^{2}\ln\left(\frac{\gamma(1-q)}{q(\Delta+\beta-\gamma)}\right)}{2\mu}.$ Hence, if $\gamma<\gamma_{0}^{+}$, then beliefs after $s>0$ are $q$ if $s\geq s^{+}_{0}$ and $\pi(s,\mathbf{p},\tilde{\mu})\in(q,1)$ else.
-Moreover, if $s<0$ and $\gamma<\gamma_{0}^{-}$, then beliefs are 00 if $s\in[s^{-}_{0},0)$ and $\pi(s)\in(0,q)$ else.
-Note that all voters with $s<0$ hold beliefs weakly below Bayesian beliefs.
-Because of motivated beliefs, they vote to a greater extent informatively than they should, which is positive for information aggregation.
-Because $q\leq\tilde{\pi}$, they all vote informatively.
-If $\omega=0$, therefore a majority votes for policy 0.
-Hence, what matters for candidate incentives is the share of $s>0$ voters voting informatively.
-Because $q<\tilde{\pi}$, those with a belief of $q$ vote for policy 0.
-That is, all voters with $s>s_{0}^{+}$.
-Those with $s\in(0,s^{+})$ may vote either way.
-The belief of a voter who chooses $\tilde{\mu}=\mu$ and receives signal $s=s^{+}_{0}$ is $\pi(s_{0}^{+},\mathbf{p},0)=\frac{\gamma+\Delta q+q}{\Delta+1}$. This is smaller than $\tilde{\pi}$ if
-$\frac{\gamma+\Delta q+q}{\Delta+1}<\frac{1}{1+\beta}\Leftrightarrow\gamma<\tilde{\gamma}\equiv\frac{(\Delta+1)(1-q(1+\beta))}{\beta+1}>0.$
-The belief of a voter who chooses $\tilde{\mu}=\mu$ and receives signal $s>0$ is
-$\pi(s,\mathbf{p},\tilde{\mu})=\frac{q}{q+(1-q)e^{-\frac{2s\mu}{\sigma^{2}}}}.$
-The signal at which such a voter starts to vote for policy 1 is the signal such that $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$.
-In particular, this signal realization is $s^{*}$, where $s^{*}\equiv\frac{\sigma^{2}\ln\left(\frac{1-q}{\beta q}\right)}{2\mu}$.
-If $\Phi\left(\frac{s-\mu}{\sigma}\right)$ is the CDF of the normal distribution with mean $\mu$ and standard deviation $\sigma$, the share of voters voting for policy 1 if $\omega=1$ is
-$\mathcal{V}_{0}=\Phi\left(\frac{s^{+}_{0}-\mu}{\sigma}\right)-\Phi\left(\frac{s^{*}-\mu}{\sigma}\right).$
-Recall that $s_{0}^{+}$ is increasing in $\gamma$, and hence $\mathcal{V}_{0}$ is monotonically increasing in $\gamma$.
-When $\gamma=0$, we know from our earlier analysis that voters all vote for policy 0 despite $s>0$.
-If $\gamma\rightarrow\infty$, no voters distorts $\mu$ and, by Assumption 2, a majority votes for policy 1 iff $\omega=1$.
-By continuity, there is a unique $\hat{\gamma}$ such that a majority votes for policy 0 if and only if $\gamma<\hat{\gamma}$.
-∎
-A.10 Proof of Lemma 4
-To see incentives, first consider how $W(s,\mathbf{p},\tilde{\mu})$ changes with $\tilde{\mu}$, when $\tilde{\mu}\geq\mu$.
-It is easy to show that $W(s,\mathbf{p},\tilde{\mu})$ is strictly concave in $\tilde{\mu}$ and thus an interior equilibrium exists for all $s$.
+This proves the proposition. ∎
+Appendix B Costly Distortions
+B.1 Cost from Imperfect Decision Making
+Recall that voters aim at maximizing $W^{UC}=\lambda AU+(1-\lambda)C^{U}$.
+Each voter chooses $\tilde{\mu}$ to maximize $W$, taking into account that $\tilde{\mu}\neq\mu$ may negatively impact the quality of the decision at the ballot.
+When $\tilde{\mu}$ is chosen in a way that is inconsequential for the decision at the ballot, $C^{U}=0$. Otherwise, $C^{U}$ equals the expected cost from deciding incorrectly.
+If a voters chooses to vote for policy 1 at the ballot, her expected utility is
+$u_{1}=\frac{-\Delta qe^{\frac{2\mu s}{\sigma^{2}}}+q-1}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}.$
+When voting for policy 0, expected utility is
+$u_{0}=-\frac{q(b+\Delta)}{q-(q-1)e^{-\frac{2\mu s}{\sigma^{2}}}}$
+Hence, a voter votes for policy 1 iff
+$u_{1}\geq u_{0}\Leftrightarrow s>\tilde{s}:=\frac{\sigma^{2}\log\left(\frac{1-q}{bq}\right)}{2\mu}>0.$
+Any voter with $s<\tilde{s}$ will vote for policy 0 with Bayesian information processing, and this does not change with motivated beliefs.
+Thus, If $s<\tilde{s}$, then $C^{U}=0$ because the choice at the ballot is identical to the Bayesian optimal choice.
+If $s>\tilde{s}$, this may change, as the Bayesian optimal choice would be to vote for policy 1, whereas with distorted information processing policy 0 might be chosen.
+If the choice at the ballot is indeed altered by choosing $\tilde{\mu}\neq\mu$, then the voter incurs a cost of
+$C^{U}=-b+\frac{(b+1)(1-q)}{1-q+qe^{\frac{2\mu s}{\sigma^{2}}}}.$
+This equals the expected utility loss from taking an incorrect decision.
+As explained in the main text, when deciding if and how to distort beliefs, a voter has only two potentially optimal choices: $\tilde{\mu}=0$ and $\tilde{\mu}=\mu_{c}$, where $\mu_{c}$ is the distortion that induces a posterior belief of $1/(1+b)$.
+These lead to the following welfare levels:
+$\begin{array}[]{rcl}W^{UC}(\tilde{\mu}=0)&=&-(1-\lambda)\left(b+\frac{(b+1)(q-1)}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}\right)-\lambda q(b+\Delta)\\ W^{UC}(\tilde{\mu}=\mu_{c})&=&-\frac{\lambda(b+\Delta)}{b+1}\end{array}$
+Note that
+$\left.W^{UC}(\tilde{\mu}=0)\right|_{s=\tilde{s}}-W(\tilde{\mu}=\mu_{c})=\frac{\lambda(b+\Delta)(1-bq-q)}{b+1}>0$
+for all $\lambda>0$. Moreover,
+$\frac{\partial W^{UC}(\tilde{\mu}=0)}{\partial s}=-\frac{2(b+1)(1-\lambda)\mu(1-q)qe^{\frac{2\mu s}{\sigma^{2}}}}{\sigma^{2}\left[q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1\right]^{2}}<0.$
+Therefore, if
+$\begin{array}[]{c}\lim_{s\to\infty}W^{UC}(\tilde{\mu}=0)>W(\tilde{\mu}=\mu_{c})\Leftrightarrow b(\lambda+\lambda(-q)-1)-\Delta\lambda q>-\frac{\lambda(b+\Delta)}{b+1}\\ \Leftrightarrow\lambda>\tilde{\lambda}:=-\dfrac{b(b+1)}{\Delta(bq+q-1)+b(b(q-1)+q-2)}\in(0,1),\end{array}$
+then any voter with $s>0$ will choose $\tilde{\mu}^{*}=0$. Conversely, if $\lambda<\tilde{\lambda}$, there exists $\hat{s}>\tilde{s}$, where
+$\hat{s}=\frac{\sigma^{2}}{2\mu}\log\!\left(\frac{(q-1)\!\left[-\Delta\lambda+\lambda+b^{2}\lambda q+b\big((\Delta+1)\lambda q-1\big)+\Delta\lambda q-1\right]}{q\!\left[b+b^{2}(\lambda(q-1)+1)+b\lambda(\Delta q+q-2)+\Delta\lambda(q-1)\right]}\right),$
+such that $\tilde{\mu}^{*}=0$ if $s\in[0,\hat{s}$] and $\tilde{\mu}^{*}=\mu_{c}$ if $s>\hat{s}$.
+If $\lambda>\tilde{\lambda}$, then all results are as in the baseline version of the model.
+Otherwise, whether or not candidates will ignore their information depends on the complexity of climate change.
+For candidates to have incentives to offer policy 1, at least 50% of voters must receive $s>\hat{s}$ if $\omega=1$.
+If $\mu$ is sufficiently large, this condition is satisfied.
+However, if complexity is high, this will not hold, and the result from the baseline model continues to apply.
+B.2 Cognitive Costs
+Following the discussion in the main text, we now establish a useful lemma:
+Lemma 2.
+[…]
+Proof.
+To see incentives, first consider how $W^{CC}$ changes with $\tilde{\mu}$, when $\tilde{\mu}\geq\mu$.
+It is easy to show that $W^{CC}$ is strictly concave in $\tilde{\mu}$ and thus an interior equilibrium exists for all $s$.
 The FOC for an interior optimum is
-$\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c(\tilde{\mu}-\mu).$
+$\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c(\tilde{\mu}-\mu).$
 if $\tilde{\mu}>\mu$ and
-$\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu).$
+$\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu).$
 if $\tilde{\mu}<\mu$.
@@
 If $s=0$, the first term is zero, and hence we need to have $\tilde{\mu}^{*}(s)=\mu$.
-Moreover, $\lim_{s\rightarrow\infty}\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu)$, and hence $\lim_{s\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$.
-Similarly, $\lim_{s\rightarrow-\infty}\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-c(\tilde{\mu}-\mu)$, and hence also $\lim_{s\rightarrow-\infty}\tilde{\mu}^{*}(s)=\mu$.
+Moreover, $\lim_{s\rightarrow\infty}\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu)$, and hence $\lim_{s\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$.
+Similarly, $\lim_{s\rightarrow-\infty}\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-c(\tilde{\mu}-\mu)$, and hence also $\lim_{s\rightarrow-\infty}\tilde{\mu}^{*}(s)=\mu$.
 Now consider $s<0$, for which we have $\tilde{\mu}>\mu$. Rearranging the FOC yields
-$-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}+\mu=\tilde{\mu}$
+$\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}+\mu=\tilde{\mu}$
 That is, we can express $\tilde{\mu}$ as a deviation from $\mu$, and
-$\Gamma=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}$
+$\Gamma=-\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}$
 defines the deviation. $\Gamma$ increases in $s$ iff
-$\frac{\partial\Gamma}{\partial s}=\frac{-2(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left((1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)\right)}{c\sigma^{4}\left(q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+1\right)^{3}}>0,$
+$\frac{\partial\Gamma}{\partial s}=\frac{-2(1-q)q\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left((1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)\right)}{c\sigma^{4}\left(q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+1\right)^{3}}>0,$
 which is the case iff $\Omega=(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)<0$.
@@
 Next consider $s>0$, implying $\tilde{\mu}<\mu$. Rearranging the FOC yields
-$-\frac{2\tilde{\mu}^{3}(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}+\mu=\tilde{\mu}$
+$\frac{2\tilde{\mu}^{3}(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}+\mu=\tilde{\mu}$
 That is, we can again express $\tilde{\mu}$ again as a deviation from $\mu$, and
-$\Gamma^{\prime}=-\frac{2\tilde{\mu}^{3}(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}$
+$\Gamma^{\prime}=-\frac{2\tilde{\mu}^{3}(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}$
 defines the deviation. $\Gamma^{\prime}$ increases in $s$ iff
-$\frac{\partial\Gamma^{\prime}}{\partial s}=\frac{2\tilde{\mu}^{3}(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)\right)}{c\sigma^{4}\left(\mu+\mu q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)\right)^{3}}>0$
+$\frac{\partial\Gamma^{\prime}}{\partial s}=\frac{2\tilde{\mu}^{3}(1-q)q\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)\right)}{c\sigma^{4}\left(\mu+\mu q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)\right)^{3}}>0$
 This is the case iff $\Omega^{\prime}=qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)>0$.
@@
 ∎
-A.11 Proof of Proposition 6
-The proof basically follows along the lines of the discussion in the text.
-We know from Lemma 4 that $\left|\tilde{\mu}(s)-\mu\right|$ decreases in $c$.
-Moreover, $\lim_{c\rightarrow\infty}\tilde{\mu}(s)=\mu$, because then both $\Gamma=0$ and $\Gamma^{\prime}=0$ (see the proof of Lemma 4).
-Moreover, as $c\searrow 0$, we are back in the scenario of Proposition 3.
-Finally note that both $\tilde{\mu}(s)$ and $\pi(s,\mathbf{p},\tilde{\mu})$ are continuous functions of both $s$ and $c$, and thus so is the share of voters voting for policy 1 in state $\omega=1$ if $p_{1}\neq p_{2}$.
-Hence, there must exist $\hat{c}$ such that the share of voters $\mathcal{V}_{0}$ voting for policy 1 in state 1 is just $\frac{1}{2}$.
-Moreover, $\text{Sign}\left[c-\hat{c}\right]=\text{Sign}\left[\mathcal{V}_{0}-\frac{1}{2}\right]$.
-This proves the proposition.
-∎
-Appendix B Additional Tables and Figures
-Figure 5: Q-Q plots for the different regressions reported in Table 2.
+The remaining discussion from the main text follows from the lemma.